### v1

# On RoCK blocks of double covers of symmetric and alternating groups and the refined Broué conjecture

Abstract
Recently, Kleshchev and Livesey [21, 22] proved the existence of RoCK $p$-blocks for double covers of symmetric and alternating groups over large enough coefficient rings. They proved that RoCK blocks of double covers are Morita equivalent to standard “local” blocks via bimodules with endopermutation source. Based on this, Kleshchev and Livesey proved that these RoCK blocks are splendidly Rickard equivalent to their Brauer correspondents. The analogous result for blocks of symmetric groups, a theorem of Chuang and Kessar, was an important step in Chuang and Rouquier ultimately proving Broué’s abelian defect group conjecture for symmetric groups. In this paper we show that in most cases the Morita equivalences and splendid Rickard equivalences constructed by Kleshchev and Livesey descend to the ring $\mathbb{Z}_{p}$ of $p$-adic integers, hence prove Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture for most of these RoCK blocks.
keywords:
splendid Rickard equivalence , descent , abelian defect group conjecture , symmetric groups , alternating groups , double covers , RoCK blocks
1 Introduction
In [19], Kessar and Linckelmann proposed a refined version of Broué’s abelian defect group conjecture by considering non splitting coefficient rings:
Conjecture 1.1 (The refined Broué conjecture [19]).
For an arbitrary complete discrete valuation ring $\mathcal{O}$ and a block idempotent $b$ of a finite group $G$ over $\mathcal{O}$ with an abelian defect group, there is a splendid Rickard equivalence between $\mathcal{O}Gb$ and its Brauer correspondent.
Comparing to the refined version, the usual Broué conjecture is with the assumption that the complete discrete valuation rings is large enough. Conjecture 1.1 has been proved for blocks with cyclic defect groups in [19], blocks with a Klein four defect group in [12], blocks of alternating groups in [13], blocks of ${\rm SL}_{n}(q)$ and ${\rm GL}_{n}(q)$ in defining characteristic (see [15]), and unipotent blocks ${\rm GL}_{n}(q)$ in non-defining characteristic (see [16]). Conjecture 1.1 has also been verified for certain blocks of $p$-nilpotent and $p$-solvable groups in [3], [27] and [28]. Since any field is a splitting field for symmetric groups, Conjecture 1.1 has been proved for blocks of symmetric groups by Chuang and Rouquier [6, Theorem 7.6].
For a non-negative integer $n$, denote by $S_{n}$ the symmetric group on $n$ letters and by $\widetilde{S}_{n}$ one of the two double covers $\widetilde{S}_{n}^{+}$ and $\widetilde{S}_{n}^{-}$ of $S_{n}$. Denote by $A_{n}$ the corresponding alternating group and by $\widetilde{A}_{n}$ its double cover. In this paper we provide additional evidence for Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture by considering blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$.
Let $\mathcal{O}$ be a complete discrete valuation ring. Denote by $z$ the central element in $\widetilde{S}_{n}$ of order $2$. Since $\mathcal{O}\widetilde{S}_{n}(1+z)/2$ is isomorphic to $\mathcal{O}S_{n}$ and $\mathcal{O}\widetilde{A}_{n}(1+z)/2$ is isomorphic to $\mathcal{O}A_{n}$, Conjecture 1.1 is true for blocks of $\mathcal{O}\widetilde{S}_{n}(1+z)/2$ and of $\mathcal{O}\widetilde{A}_{n}(1+z)/2$. So we restrict our attention in this paper to blocks of $\mathcal{O}\widetilde{S}_{n}(1-z)/2$ and of $\mathcal{O}\widetilde{A}_{n}(1-z)/2$, which are called the spin blocks of the symmetric and alternating groups, respectively. For block with trivial defect group, Conjecture 1.1 is trivially true, hence we only need to consider Conjecture 1.1 for block with nontrivial defect groups. Let $K$ denote the field of fractions of $\mathcal{O}$ and let $k$ denote the residue field of $\mathcal{O}$, i.e. $k=\mathcal{O}/J(\mathcal{O})$. Let $p:={\rm char}(k)$, a prime number. For the rest of this introduction we assume that $K$ has characteristic 0, and when dealing with a finite group $G$, we will assume that $K$ (and consequently $\mathcal{O}$) contains a primitive $|G|$-th root of unity; we also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$). Let $\mathbb{Z}_{p}$ denote the ring of $p$-adic integers, $\mathbb{Q}_{p}$ the field of $p$-adic numbers and $\mathbb{F}_{p}$ the field with $p$ elements. Since $z$ is a central element of order $2$, if $p=2$, then $\widetilde{S}_{n}$ (resp. $\widetilde{A}_{n}$) satisfies Conjecture 1.1 if and only if $S_{n}\cong\widetilde{S}_{n}/\langle z\rangle$ (resp. $A_{n}\cong\widetilde{A}_{n}/\langle z\rangle$) satisfies Conjecture 1.1; see [31, Lemma 10.2.11] (we note that the blanket assumption in [31] that the coefficient ring is large enough is not used in the proof of [31, Lemma 10.2.11]). Hence Conjecture 1.1 holds for $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ if $p=2$. So from now on we assume that $p\geq 3$.
Let $R\in\{\mathcal{O},k\}$, $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$ and let $e_{z}:=(1-z)/2\in R\widetilde{S}_{n}$. From [21, 22] we see that it is important to consider $R\tilde{S}_{n}$ as a superalgebra via
$(R\widetilde{S}_{n})_{\bar{0}}:=R\widetilde{A}_{n}~~~{\rm and}~~~(R\widetilde{S}_{n})_{\bar{1}}:=R(\widetilde{S}_{n}\backslash\widetilde{A}_{n}).$
Recall from [22] that the spin superblock algebras (i.e. the indecomposable superideals) of $R\tilde{S}_{n}e_{z}$ are labeled by pairs $(\rho,d)$, where $\rho$ is a $\bar{p}$-core ($p$-bar core) and $d\in\mathbb{N}$ (called the weight of the block) such that $n=|\rho|+d\rho$. Denote such a superblock algebra by $B^{\rho,d}_{R}$ and denote by $e_{\rho,d}$ the central idempotent of $R\widetilde{S}_{n}$ such that $B^{\rho,d}_{R}=R\widetilde{S}_{n}e_{\rho,d}$. We prove that the idempotent $e_{\rho,d}$ is contained in the subring $R_{p}\widetilde{S}_{n}$ (see Proposition 5.9 below). Hence we can define $B^{\rho,d}_{R_{p}}:=R_{p}\widetilde{S}_{n}e_{\rho,d}$. Note that the case $d=0$ corresponds to the defect zero situations, so we assume that $d>0$ in the rest of this section. In this case $B^{\rho,d}_{R}$ is in fact a spin block algebra (and not just a superblock algebra) of $R\widetilde{S}_{n}$.
We denote by $D^{\rho,d}$ a defect group of $B^{\rho,d}_{R}$. The defect group $D^{\rho,d}$ of $B^{\rho,d}_{R}$ is abelian if and only if $d<p$. Moreover, if $d<p$, then $D^{\rho,d}\cong(C_{p})^{\times d}$. By [22, Remark 4.4], the idempotent $e_{\rho,d}$ is contained in the subalgebra $R\widetilde{A}_{n}e_{z}$, and hence (by Proposition 5.9) in $R_{p}\widetilde{A}_{n}e_{z}$. By [22, Theorem 4.5] (or [18, Propsition 3.16]), as an element of $R\widetilde{A}_{n}$, $e_{\rho,d}$ is a single block unless $d=0$ and $\rho$ is even. So $B^{\rho,d}_{R}$ can be regarded as an $R$-superalgebra with even part $(B^{\rm\rho,d}_{R})_{\bar{0}}:=R\widetilde{A}_{n}e_{\rho,d}$. We denote $R_{p}\widetilde{A}_{n}e_{\rho,d}$ by $(B^{\rm\rho,d}_{R_{p}})_{\bar{0}}$. It is known that $D^{\rho,d}$ is also a defect group of $B^{\rho,d}_{\bar{0}}$ (see [18, Proposition 3.16] or Theorem 5.8 below). Let $b^{\rho,d}_{R}$ (resp. $(b^{\rho,d}_{R})_{\bar{0}}$) be the Brauer correspondent of $B^{\rho,d}_{R}$ (resp. $(B^{\rho,d}_{R})_{\bar{0}}$) in $RN_{\widetilde{S}_{n}}(D^{\rho,d})$ (resp. $RN_{\widetilde{A}_{n}}(D^{\rho,d})$). Similarly, we let $b^{\rho,d}_{R_{p}}$ (resp. $(b^{\rho,d}_{R_{p}})_{\bar{0}}$) be the Brauer correspondent of $B^{\rho,d}_{R_{p}}$ (resp. $(B^{\rho,d}_{R_{p}})_{\bar{0}}$) in $R_{p}N_{\widetilde{S}_{n}}(D^{\rho,d})$ (resp. $R_{p}N_{\widetilde{A}_{n}}(D^{\rho,d})$). Note that $(b^{\rho,d}_{R})_{\bar{0}}$ (resp. $(b^{\rho,d}_{R_{p}})_{\bar{0}}$) is indeed the even part of the superalgebra $B^{\rho,d}_{R}$ (resp. $B^{\rho,d}_{R_{p}}$). Let $\mathcal{T}_{n}^{R}:=R\widetilde{S}_{n}e_{z}$. For a superalgebra $A$, a twisted wreath superproduct $A\wr_{s}\mathcal{T}_{d}^{R}$ is defined in [22, §3.7] (see also 2.7 below).
In [21], Kleshchev and Livesey proved the following structure theorem of $b^{\rho,d}_{k}$:
Theorem 1.2 ([21, Theorem A]).
Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Then
  1. 1.
    $b^{\rho,d}_{k}\cong B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})$ as $k$-superalgebras.
  2. 2.
    $(b^{\rho,d}_{k})_{\bar{0}}\cong(B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}))_{\bar{0}}$ as $k$-algebras.
Based on the above theorem we prove a similar structure theorem of $b^{\rho,d}_{\mathbb{F}_{p}}$:
Theorem 1.3 (Theorem 6.2).
Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$. Then
  1. 1.
    $b^{\rho,d}_{\mathbb{F}_{p}}\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})$ as $\mathbb{F}_{p}$-superalgebras.
  2. 2.
    $(b^{\rho,d}_{\mathbb{F}_{p}})_{\bar{0}}\cong(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}))_{\bar{0}}$ as $\mathbb{F}_{p}$-algebras.
In [21], Kleshchev and Livesey proved the following analogue of “the Rickard and Marcus step for symmetric groups” for double covers of symmetric and alternating groups:
Theorem 1.4 ([21, Theorem C]).
Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Then
  1. 1.
    $B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})$ is derived equivalent to $B^{\rho,0}_{k}\otimes_{k}(B^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})$.
  2. 2.
    $(B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}))_{\bar{0}}$ is derived equivalent to $(B^{\rho,0}_{k}\otimes_{k}(B^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}))_{\bar{0}}.$
We prove a similar result over $\mathbb{F}_{p}$:
Theorem 1.5 (Theorem 6.6).
Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$. Then
  1. 1.
    $B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})$ is splendidly Rickard equivalent to $B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})$.
  2. 2.
    $(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}))_{\bar{0}}$ is splendidly Rickard equivalent to $(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}))_{\bar{0}}$.
See Definitions 4.9 and 4.11 below for splendid Rickard equivalences between general interior $P$-algebras. In [21], Kleshchev and Livesey defined the notion of a $d$-Rouquier $\bar{p}$-core for $d\in\mathbb{N}$ (see 5.16 below) and proved the following result:
Theorem 1.6 ([21, Theorem D], [22, Theorem A]).
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Then $B^{\rho,d}_{R}$ is Morita superequivalent to $B^{\rho,0}_{R}\otimes_{R}(B^{\emptyset,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})$ via a bimodule with endopermutation source.
For the definition of a Morita superequivalence, see [22, §3.3] or 2.4 below. Following [22] we call the block $e_{\rho,d}$ of $R\widetilde{S}_{n}$ in Theorem 1.6 an RoCK block of $\widetilde{S}_{n}$. Since $e_{\rho,d}$ is also a block idempotent of $R\widetilde{A}_{n}$, following [22] we also called the the block $e_{\rho,d}$ of $R\widetilde{A}_{n}$ an RoCK block of $\widetilde{A}_{n}$. Using Theorem 1.6, Kleshchev and Livesey proved Broué’s abelian defect group conjecture for RoCK blocks of $R\widetilde{S}_{n}$ and $R\widetilde{A}_{n}$:
Theorem 1.7 ([22, Theorem B]).
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Then $B^{\rho,d}_{R}$ and $(B^{\rho,d}_{R})_{\bar{0}}$ are splendidly Rickard equivalent to $b^{\rho,d}_{R}$ and $(b^{\rho,d}_{R})_{\bar{0}}$ respectively.
In this paper we prove that in most cases Theorems 1.6 and 1.7 descend to $R_{p}$:
Theorem 1.8 (Theorem 7.4).
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$. Then $B^{\rho,d}_{R_{p}}$ is Morita superequivalent to $B^{\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\emptyset,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})$.
Corollary 7.5 (a corollary of Theorem 1.8) answers a question of Kleshchev and Livesey proposed in [22, Remark 8.15] in the highest generality of their question.
Theorem 1.9 (Theorem 7.10).
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$. Then $B^{\rho,d}_{R_{p}}$ and $(B^{\rho,d}_{R_{p}})_{\bar{0}}$ are splendidly Rickard equivalent to $b^{\rho,d}_{R_{p}}$ and $(b^{\rho,d}_{R_{p}})_{\bar{0}}$ respectively. In particular, Conjecture 1.1 holds for RoCK blocks of $\widetilde{S}_{n}^{+}$ and $\widetilde{A}_{n}$, and if $p\equiv 1~{\rm mod}~4$, Conjecture 1.1 also holds for RoCK blocks of $\widetilde{S}_{n}^{-}$.
Besides Theorems 1.3, 1.5 and 1.8, our proof of Theorem 1.9 makes use of a recent theorem of Miller [28, Theorem 1.3 (b)], which gives a descent criterion of an endosplit $p$-permutation resolution of a module.
Remark 1.10.
For the group $\widetilde{S}_{n}^{-}$, if $p\equiv 3~{\rm mod}~4$, Theorems 1.2, 1.4, 1.6 and 1.7 do not descend to $R_{p}$, but descend to the subring of $R$ generated by $R_{p}$ and a primitive $4$-th root of unity $\sqrt{-1}$. The reason is that by (5.2) below we have an isomorphism of superalgebras
$R\widetilde{S}_{n}^{-}e_{z}\cong\mathcal{T}_{n}^{R},~t_{i}e_{z}\mapsto(-1)^{i}\sqrt{-1}\tau_{i},\ ze_{z}\mapsto-1,$
but the element $\sqrt{-1}\in R$ is not in the subring $R_{p}$.
Using [1, Theorem 1.4] and [14, Theorem 1], we obtain a corollary of Theorem 1.9:
Corollary 1.11.
Refinements of the Alperin–McKay conjecture by Isaacs–Navarro, Navarro, Turull and Boltje (see [1, (1.5), (1.6), (1.7), (1.8)]) and the blockwise Navarro/Galois Alperin weight conjecture (see [14, Conjecture 2]) hold for RoCK blocks of $\widetilde{S}_{n}^{+}$ and $\widetilde{A}_{n}$, and if $p\equiv 1~{\rm mod}~4$, these conjectures also hold for RoCK blocks of $\widetilde{S}_{n}^{-}$.
The blockwise Navarro Alperin weight conjecture has been proved for all blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ in [8] using different method.
Let $\mathbb{P}$ be the set of prime numbers, and $\mathbb{P}_{1}:=\{p\in\mathbb{P}\mid p\equiv 1~{\rm mod}~4\}$. Now consider the following three families of group-prime pairs: $\Omega_{1}:=\{(\widetilde{S}_{n}^{+},p)\mid n\in\mathbb{N},~p\in\mathbb{P}\}$, $\Omega_{2}:=\{(\widetilde{A}_{n},p)\mid n\in\mathbb{N},~p\in\mathbb{P}\}$, $\Omega_{3}:=\{(\widetilde{S}_{n}^{-},p)\mid n\in\mathbb{N},~p\in\mathbb{P}_{1}\}$. For a partition $\lambda$ we denote by $h(\lambda)$ the number of nonzero parts of $\lambda$. Now by the results (Theorems 1.3, 1.5, 1.8) in this paper, to prove Conjecture 1.1 for all spin blocks of these three families, it suffices to prove the following refined version of the Kessar–Schaps conjecture [20]:
Conjecture 1.12.
Arbitrary spin blocks $B^{\rho,d}_{R_{p}}$ and $B^{\rho^{\prime},d^{\prime}}_{R_{p}}$ of symmetric groups (or spin blocks $(B^{\rho,d}_{R_{p}})_{\bar{0}}$ and $(B^{\rho^{\prime},d^{\prime}}_{R_{p}})_{\bar{0}}$ of alternating groups) are splendidly Rickard equivalent if and only if $d=d^{\prime}$ and $|\rho|-h(\rho)\equiv|\rho^{\prime}|-h(\rho^{\prime})~({\rm mod}~2)$.
If one replaces “splendidly Rickard equivalent” with “derived equivalent” and replaces $R_{p}$ with $k$, the above conjecture is the usual Kessar–Schaps conjecture. Recently, proofs of the usual Kessar–Schaps conjecture were announced in [9] and [4].
This paper is organised as follows: In Section 2 we review some relevant results on superalgebras and supermodules which are necessary for later use. In Section 3, we prove general descent results for supermodules and superequivalences. In Section 4, we formulate convenient reference for results on graded Rickard equivalences and splendid complexes. In Section 5 we review some block theory of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$. Base on [21], we prove in Section 6 the structure results for Brauer correspondents of spin blocks over $\mathbb{F}_{p}$. In Section 7, we analyse the Morita superequivalence bimodule for an RoCK block constructed by Kleshchev and Livesey [22], and then realise this bisupermodule over $\mathbb{Z}_{p}$; we prove the refined Broué conjecture for RoCK spin blocks in the end of Section 7.
2 Preliminaries
2.1.
Generalities. Throughout the rest of this paper $p$ is an odd prime, $\mathcal{O}$ is a complete discrete valuation ring with residue field $k$ of characteristic $p$ and quotient field $K$ of characteristic 0. Such a triple $(K,\mathcal{O},k)$ is called a $p$-modular system. (Note that we didn’t assume $(K,\mathcal{O},k)$ to be large enough.) An extension of the $p$-modular system $(K,\mathcal{O},k)$ is a $p$-modular system $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ such that $\mathcal{O}$ is a subring of $\mathcal{O}^{\prime}$ with $J(\mathcal{O})\subseteq J(\mathcal{O}^{\prime})$. For an algebra $A$, we denote by $A^{\rm op}$ the opposite algebra of $A$, by $A^{\times}$ the set of units in $A$, by $A$-mod the category of finitely generated $A$-modules, and by $K^{b}(A)$ the homotopy category of bound complexes over $A$-mod (see e.g. [23, page 117]). Unless otherwise stated, modules in this paper are left modules. By a block of $\mathcal{O}G$ (resp. $kG$) we usually mean an indecomposable direct factor $B$ of $\mathcal{O}G$ (resp. $kG$). Some times we also mean the correspondent primitive central idempotent.
Let $R\in\{\mathcal{O},k\}$ and $A$ a finitely generated $R$-algebra. A projective cover $(P,\varphi)$ of a finitely generated $A$-module $M$ is a projective $A$-module $P$ and a surjective $A$-module homomorphism $\varphi:P\to M$ such that no proper direct summand of $P$ surjects onto $M$. The assumption on $A$ ensure that projective covers of $M$ always exist and are unique up to isomorphism. The $A$-module $\Omega_{A}(M):={\rm ker}(\varphi)$ is called the Heller translate of $M$. We inductively define powers of the Heller translate $\Omega_{A}^{n}$, for a positive integer $n$. Finally, $\Omega_{A}^{0}(M)$ is defined as the direct sum of all the non-projective indecomposable summands of $M$.
[…]
$2.$
Superalgebras and supermodules. We review some definitions and notation from [22, §3]. Let $R$ be a commutative ring. An $R$-superspace is a finitely generated $R$-module $V$ with decomposition $V=V_{\bar{0}}\oplus V_{\bar{1}}$. For $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$, we refer to the elements of $V_{\varepsilon}$ as the homogeneous elements of parity $\varepsilon$, and write $|v|=\varepsilon$ for $v\in V_{\varepsilon}$. The subspace $V_{\bar{0}}$ (resp. $V_{\bar{1}}$) is called the even (resp. odd ) part of $V$. An $R$-subsuperspace $W$ of $V$ is an $R$-submodule of $V$ such that $W=(W\cap V_{\bar{0}})\oplus(W\cap V_{\bar{1}})$. Of course we can treat $W$ as an $R$-subsuperspace in its own right. Let $V$ and $W$ be $R$-superspaces. The direct sum $V\oplus W$ is considered as an $R$-superspace in an obvious way. The tensor product $V\otimes_{R}W$ is considered as an $R$-superspace via $|v\otimes w|:=|v|+|w|$. A homomorphism $f:V\to W$ of $R$-superspaces is an $R$-linear map satisfying $f(V_{\varepsilon})\subseteq W_{\varepsilon}$, for $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$.
An $R$-superalgebra is an $R$-superspace AA that is also an $R$-algebra such that $A_{\varepsilon}A_{\delta}\subseteq A_{\varepsilon+\delta}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$. A homomorphism $f:A\to B$ of $R$-superalgebras is an algebra homomorphism that is also a homomorphism of $R$-superspaces. Let $A$ be an $R$-superalgebra. If $A_{\bar{1}}\cap A^{\times}\neq\emptyset$, then $A$ is called a superalgebra with superunit, and any $u\in A_{\bar{1}}\cap A^{\times}$ is called a superunit.
An $A$-supermodule $M$ is an $A$-module which is also an $R$-superspace such that $A_{\varepsilon}M_{\delta}\subseteq M_{\varepsilon+\delta}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$. If $B$ is another $R$-superalgebra, an AA-$B$-bisupermodule module $M$ is an AA-$B$-bimodule which is also an $R$-superspace such that $A_{\varepsilon}M_{\delta}\subseteq M_{\varepsilon+\delta}$ and $M_{\delta}B_{\varepsilon}\subseteq M_{\varepsilon+\delta}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$. A homomorphism $f:M\to N$ of $A$-supermodules is an $A$-module homomorphism that is also a homomorphism of $R$-superspaces; a homomorphism of bisupermodules is defined similarly.
Assume that $1\neq 1$ in $R$. We defined the $R$-superalgebra homomorphism
$\sigma=\sigma_{A}:A\to A,~a\mapsto(-1)^{|a|}a.$ (2.1)
Note that the superstructure (i.e. the $\mathbb{Z}/2\mathbb{Z}$-grading on $A$) is completely determined by $\sigma_{A}$. The tensor product $A\otimes_{R}B$ of $R$-superalgebras AA and $B$ is considered to be an $R$-superalgebra via
$(a\otimes b)(a^{\prime}\otimes b^{\prime})=(-1)^{|b||a^{\prime}|}aa^{\prime}\otimes bb^{\prime}.$
Following [22], we define the superalgebra $A^{\rm sop}$ (denoted $A^{\rm op}$ in [21]) to be equal to $A$ as an $R$-superspace but with multiplication given by $a\cdot b=(-1)^{|a||b|}ba$, for all $a,b\in A$. We can view an $(A,B)$-bisupermodule $M$ as an $(A\otimes_{R}B^{\rm sop})$-supermodule via
$(a\otimes b)\cdot m=(-1)^{|b||m|}amb,$
for all $a\in A$, $b\in B$ and $m\in M$. In this way, we identify the notions of an $(A,B)$-bisupermodule and an $(A\otimes_{R}B^{\rm sop})$-supermodule.
If $V$ is an $R$-superspace we denote by $|V|$ the $R$-space which is $V$ with grading forgotten. In particular, if $A$ is an $R$-superalgebra and $M$ is an $A$-supermodule, we have an $R$-algebra $|A|$ and an $|A|$-module $|M|$. Similarly, if $B$ is another superalgebra and $M$ is an $(A,B)$-bisupermodule, we have an $(|A|,|B|)$-bisupermodule $|M|$. An $A$-supermodule $M$ is called module-indecomposable if $|M|$ is indecomposable as an $|A|$-module. An $(A,B)$-bisupermodule $M$ is called module-indecomposable if $|M|$ indecomposable as an $(|A|,|B|)$-bimodule.
Let $A,B,C$ and $D$ be $R$-superalgebras, $M$ an $(A,C)$-bisupermodule and $N$ a $(B,D)$-bisupermodule. We define the $(A\otimes_{R}B,C\otimes_{R}D)$-bisupermodule $M\boxtimes_{R}N$ to be the $R$-superspace $M\otimes_{R}N$ with the action
$(a\otimes b)(m\otimes n)=(-1)^{|b||m|}(am\otimes bn),\qquad(m\otimes n)(c\otimes d)=(-1)^{|c||n|}(mc\otimes nd),$
for all $a\in A$, $b\in B$, $c\in C$, $d\in D$, $m\in M$ and $n\in N$. In particular, given an $A$-supermodule $M$ and a $B$-supermodule NN, we have the $(A\otimes_{R}B)$-supermodule $M\boxtimes_{R}N$ with the action $(a\otimes b)(m\otimes n)=(-1)^{|b||m|}(am\otimes bn).$ If $X$ is a complex of $(A,C)$-bisupermodules with differential $\delta=(\delta_{n}:X_{n}\to X_{n-1})_{n\in\mathbb{Z}}$ and $Y$ is a complex of $(B,D)$-bisupermodules with differential $\epsilon=(\epsilon_{n}:Y_{n}\to Y_{n-1})_{n\in\mathbb{Z}}$, then we can define the complex of $(A\otimes_{R}B,C\otimes_{R}D)$-bisupermodules $X\boxtimes_{R}Y$ with $(X\boxtimes_{R}Y)_{n}=\oplus_{i+j=n}(X_{i}\boxtimes_{R}Y_{j})$ and with differential given by taking the direct sum of the maps $(-1)^{i}{\rm id}_{X_{i}}\otimes\epsilon_{j}:X_{i}\boxtimes_{R}Y_{j}\to X_{i}\boxtimes_{R}Y_{j-1}$ and $\delta_{i}\otimes{\rm id}_{Y_{j}}:X_{i}\boxtimes_{R}Y_{j}\to X_{i-1}\boxtimes_{R}Y_{j}$.
Remark 2.3.
Note that any finite generated algebra (module, bimodule) $A$ can be regarded as an superalgebra (a supermodule, a bisupermodule) with $A$ as the even part and 0 as the odd part. So the above tensor products, as well as most other results in this paper, for superalgebras, supermodules and bisupermodules apply to usual algebras, modules and bimodules.
2.4.
Morita superequivalences. Let $R$ be a commutative ring and let $A$, $B$ be $R$-superalgebras. Following [22], a Morita superequivalence between $A$ and $B$ is a Morita equivalence between $A$ and $B$ induced by an $(A,B)$-bisupermodule $M$ and a $(B,A)$-bisupermodule NN, i.e. $M\otimes_{B}N\cong A$ as $(A,A)$-bisupermodules and $N\otimes_{A}M\cong B$ as $(B,B)$-bisupermodules.
2.5.
Dual supermodules. Let $R$ be a commutative ring with $1\neq 1$ in $R$ and let $A$ be an $R$-superalgebra and $M$ a finitely generated $A$-supermodule. Following [22] we define the dual of $M$ to be $M^{*}:={\rm Hom}_{R}(M,R)$ with the structure of an $R$-superspace through
${\rm Hom}_{R}(M,R)_{\varepsilon}:=\{f\in{\rm Hom}_{R}(M,R)\mid f(M_{\varepsilon+\bar{1}})=0\}\qquad(\varepsilon\in\mathbb{Z}/2\mathbb{Z}).$
We treat $M^{*}$ as a right $A$-supermodule via $(f\cdot a)(m):=f(am)$, for all $a\in A$, $f\in M^{*}$ and $m\in M$, hence also as an $A^{\rm sop}$-module via $(a\cdot f)(m):=(-1)^{|a||f|}f(am)$. For more properties of dual supermodules and relevant results on Morita superequivalences, we refer to [22, §3.7].
2.6.
Crossed superproducts. Let $R$ be a commutative ring with $1\neq 1$ in $R$ and $G$ a finite group. A $G$-graded crossed $R$-superproduct is an $R$-superalgebra AA with a decomposition $\bigoplus_{g\in G}A_{g}$ into subsuperspaces such that $A_{g}A_{h}\subseteq A_{gh}$, for all $g,h\in G$, and such that, for all $g\in G$, we have $A_{g}\cap A^{\times}\neq\varnothing$. If $A$ is a $G$-graded crossed $R$-superproduct, then so is $A^{\rm sop}$ by defining $(A^{\rm sop})_{g}=A_{g^{-1}}$, for all $g\in G$. Note that $A_{1_{G}}$ is always a subsuperalgebra of $A$, and $(A_{1_{G}})^{\rm sop}=(A^{\rm sop})_{1_{G}}$.
If $A$ and $B$ are $G$-graded crossed $R$-superproducts, following [22, (3.5)] we define
$(A,B)_{G}:=\sum_{g\in G}A_{g}\otimes_{R}B_{g^{-1}}=\sum_{g\in G}A_{g}\otimes_{R}(B^{\rm sop})_{g}\subseteq A\otimes_{R}B^{\rm sop}.$ (2.2)
The definition of a $G$-graded crossed superproduct ensures that
$A_{1_{G}}\otimes_{R}B_{1_{G}}^{\rm sop}\subseteq(A,B)_{G}\subseteq A\otimes_{R}B^{\rm sop}$
are $R$-subsuperalgebras.
2.7.
Let $R\in\{\mathcal{O},k\}$ and let $n$ be a positive integer. The twisted group superalgebra $\mathcal{T}_{n}^{R}$ over $R$ of the symmetric group $S_{n}$ is the $R$-superalgebra given by odd generators $\tau_{1},\cdots,\tau_{n-1}$ subject to the relations
$\tau_{r}^{2}=1,\quad\tau_{r}\tau_{s}=-\tau_{s}\tau_{r}\text{ if }|r-s|>1,\quad(\tau_{r}\tau_{r+1})^{3}=1.$
For each $w\in S_{n}$, choosing a reduced expression $w=s_{r_{1}}\cdots s_{r_{k}}$ in terms of simple transpositions, we define $\tau_{w}:=\tau_{r_{1}}\cdots\tau_{r_{k}}$ (which depends up to a sign on the choice of a reduced expression). It is well-known that $\{\tau_{w}\mid w\in S_{n}\}$ is an $R$-basis of $\mathcal{T}_{n}^{R}$ (one way to see that is to use the first isomorphism in (5.2) below).
Given an $R$-superalgebra AA, following [21, Section 5.1.a], we define the twisted wreath superproduct $A\wr_{s}\mathcal{T}_{n}^{R}$ as the free product $A^{\otimes n}\star\mathcal{T}_{n}^{R}$ of superalgebras subject to the relations
$\tau_{r}\,(a_{1}\otimes\dots\otimes a_{n})=(-1)^{\sum_{u\neq r,r+1}|a_{u}|}\left({}^{s_{r}}(a_{1}\otimes\dots\otimes a_{n})\,\tau_{r}\right)$
for all $a_{1},\cdots,a_{n}\in A$ and $1\leq r\leq n-1$, where $s_{r}$ is the transposition $(r,r+1)\in S_{n}$.
Proposition 2.8.
Let $R\in\{\mathcal{O},k\}$, $n$ a positive integer and $A$ an $R$-superalgebra. Then
$A\wr_{s}\mathcal{T}_{n}^{R}=\bigoplus_{w\in S_{n}}A^{\otimes n}\tau_{w},$ (2.3)
where $A^{\otimes n}$ denotes $\underbrace{A\otimes_{R}\cdots\otimes_{R}A}_{n}$. Hence $A\wr_{s}\mathcal{T}_{n}^{R}$ is an $S_{n}$-graded crossed $R$-superproduct.
Proof. Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$ such that $\sqrt{-2}$ (a root of the polynomial $x^{2}+2\in\mathcal{O}[x]$) is contained in $\mathcal{O}^{\prime}$. Let $R^{\prime}\in\{\mathcal{O}^{\prime},k^{\prime}\}$ and let $A^{\prime}=R^{\prime}\otimes_{R}A$. Then by [22, (3.8)], $A^{\prime}\wr_{s}\mathcal{T}_{n}^{R^{\prime}}$ is an $S_{n}$-graded crossed $R^{\prime}$-superproduct via
$A^{\prime}\wr_{s}\mathcal{T}_{n}^{R^{\prime}}=\bigoplus_{w\in S_{n}}(A^{\prime})^{\otimes n}\tau_{w}.$ (2.4)
[…]
Remark 2.9.
To explain the equality (2.4), [22, (3.8)] uses the element $\sqrt{-2}$; see [22, Proposition 3.22]. But after the explanation, $\sqrt{-2}$ can be throw away - because $\sqrt{-2}$ would not involved in this equality.
2.10.
Keep the notation of Proposition 2.8. Given the $S_{n}$-grading in (2.3), recall the definition of $(A\wr_{s}\mathcal{T}_{n}^{R},A\wr_{s}\mathcal{T}_{n}^{R})_{S_{n}}$ from (2.2). Let $M$ be an $(A,A)$-bisupermodule. Following [22, (3.9)], we denote by $M^{\boxtimes n}_{S_{n}}$ the $(A\wr_{s}\mathcal{T}_{d}^{R},A\wr_{s}\mathcal{T}_{n}^{R})_{S_{n}}$-supermodule given by extending the $(A^{\otimes n},A^{\otimes n})$-bisupermodule $M^{\boxtimes n}:=\underbrace{M\boxtimes_{R}\cdots\boxtimes_{R}M}_{n}$, via
[…]
for all $m_{1},\cdots,m_{n}\in M$ and $1\leq r\leq n-1$.
Following [22, (3.10)] we now define
$M\wr_{s}\mathcal{T}_{n}^{R}:={\rm Ind}_{(A\wr_{s}\mathcal{T}_{n}^{R},A\wr_{s}\mathcal{T}_{n}^{R})_{S_{n}}}^{(A\wr_{s}\mathcal{T}_{n}^{R})\otimes_{R}(A\wr_{s}\mathcal{T}_{n}^{R})^{\rm sop}}(M^{\boxtimes n}_{S_{n}}).$ (2.6)
2.11.
Let $R\in\{\mathcal{O},k\}$, $n$ a positive integer and $A$ an $R$-superalgebra with a superunit $u$. If $C_{2}$ is a cyclic group with generator $g$, we can give $A$ the structure of a $C_{2}$-graded crossed $R$-superproduct with graded components $A_{g^{\varepsilon}}=A_{\bar{0}}u^{\varepsilon}=A_{\varepsilon}$ for $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$. For any $n\in\mathbb{Z}_{>0}$, following [21, 5.1.c], we consider the wreath product
$C_{2}\wr S_{n}=\{g_{1}^{\varepsilon_{1}}\cdots g_{n}^{\varepsilon_{n}}w\mid\varepsilon_{1},\dots,\varepsilon_{n}\in\mathbb{Z}/2\mathbb{Z},\,w\in S_{n}\}$
where $g_{r}$ is the generator of the $r$-th $C_{2}$ factor in the base group $C_{2}^{\times n}$. Then we can give $A\wr_{s}\mathcal{T}_{n}^{R}$ the structure of a $(C_{2}\wr S_{n})$-graded crossed $R$-superproduct with graded components
$(A\wr_{s}\mathcal{T}_{n})_{g_{1}^{\varepsilon_{1}}\cdots g_{n}^{\varepsilon_{n}}w}=A_{\bar{0}}^{\otimes n}u_{1}^{\varepsilon_{1}}\dots u_{n}^{\varepsilon_{n}}\tau_{w}$
where each $u_{r}$ is $1^{\otimes(r-1)}\otimes u\otimes 1^{\otimes(n-r)}$.
Let $A$ now be a superalgebra with a superunit $u_{A}$ and $B$ be a superalgebra with a superunit $u_{B}$. Recalling (2.2), we have the subsuperalgebras
$(A,B)_{C_{2}}=(A_{\bar{0}}\otimes_{R}B_{\bar{0}}^{\rm sop})\oplus(A_{\bar{1}}\otimes_{R}B_{\bar{1}}^{\rm sop})\subseteq A\otimes_{R}B^{\rm sop},$
$(A\wr_{s}\mathcal{T}_{n}^{R},B\wr_{s}\mathcal{T}_{n}^{R})_{C_{2}\wr S_{n}}\subseteq(A\wr_{s}\mathcal{T}_{n}^{R})\otimes_{R}(B\wr_{s}\mathcal{T}_{n}^{R})^{\rm sop}.$
Setting $U_{r}:=(u_{A})_{r}\otimes(u_{B}^{-1})_{r}$, for $1\leq r\leq n$, we have that $A_{\bar{0}}^{\otimes n}\otimes_{R}(B_{\bar{0}}^{\otimes n})^{\rm sop}$ together with the $U_{r}$’s generate a subsuperalgebra of $(A\wr_{s}\mathcal{T}_{n}^{R},B\wr_{s}\mathcal{T}_{n}^{R})_{C_{2}\wr_{s}S_{n}}$ isomorphic to $((A,B)_{C_{2}})^{\otimes n}$. We identify the two superalgebras. Then $(A\wr_{s}\mathcal{T}_{n}^{R},B\wr_{s}\mathcal{T}_{n}^{R})_{C_{2}\wr_{s}S_{n}}$ is generated by $((A,B)_{C_{2}})^{\otimes n}$ and $T_{r}:=\tau_{r}\otimes\tau_{r}^{-1}$, for $1\leq r\leq n-1$ subject to the relations listed in [21, (5.1.13)].
Proposition 2.12 (a slight generalisation of [21, Proposition 5.1.14]).
Let $R\in\{\mathcal{O},k\}$, let $n$ be a positive integer, and let $A$ and $B$ be $R$-superalgebras with superunit. If $X$ is a complex of $A_{\bar{0}}$-$B_{\bar{0}}$-bimodules that extends to a complex of $(A,B)_{C_{2}}$-modules, then $X^{\otimes n}$ extends to a complex of $(A\wr_{s}\mathcal{T}_{n},B\wr_{s}\mathcal{T}_{n})_{C_{2}\wr S_{n}}$-modules.
[…]
2.13.
The groups $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$. The double covers of $S_{n}$ (where $n$ is a non-negative integer) are given by
$\widetilde{S}_{n}^{+}:=\langle z,t_{1},\cdots,t_{n-1}~|~z^{2}=1,~t_{i}z=zt_{i},~t_{i}^{2}=1,~t_{i}t_{j}=zt_{j}t_{i}~{\rm if}~|i-j|>1,~(t_{i}t_{i+1})^{3}=1\rangle,$
$\widetilde{S}_{n}^{-}:=\langle z,t_{1},\cdots,t_{n-1}~|~z^{2}=1,~t_{i}z=zt_{i},~t_{i}^{2}=z,~t_{i}t_{j}=zt_{j}t_{i}~{\rm if}~|i-j|>1,~(t_{i}t_{i+1})^{3}=1\rangle.$
We will use $\widetilde{S}_{n}$ to simultaneously represent both double covers. Denote by $\pi_{n}$ the surjective group homomorphism $\widetilde{S}_{n}\to S_{n}$, $z\mapsto 1$, $t_{i}\mapsto s_{i}$, where $s_{i}$ denotes the transposition $(i,i+1)\in S_{n}$. Then the double cover of $A_{n}$ is given by $\widetilde{A}_{n}:=\pi_{n}^{-1}(A_{n})$.
3 Descent and supermodules
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$.
Proposition 3.1.
Let $R\subseteq R^{\prime}$ be two commutative rings. Let $A$ be an $R$-superalgebra, $M$ an $A$-supermodule. Then $A^{\prime}:=R^{\prime}\otimes_{R}A$ is a superalgebra with $A^{\prime}_{\bar{0}}:=R^{\prime}\otimes_{R}A_{\bar{0}}$ and $A^{\prime}_{\bar{1}}:=R^{\prime}\otimes_{R}A_{\bar{1}}$, and $M^{\prime}:=R^{\prime}\otimes_{R}M$ is an $A^{\prime}$-supermodule with $M^{\prime}_{\bar{0}}:=R^{\prime}\otimes_{R}M_{\bar{0}}$ and $M^{\prime}_{\bar{1}}:=R^{\prime}\otimes_{R}M_{\bar{1}}$. A similar result also holds for bisupermodules instead of supermodules.
[…]
3.2.
Let $R\subseteq R^{\prime}$ be two commutative rings. Let $A$ be an $R$-superalgebra and let $A^{\prime}:=R^{\prime}\otimes_{R}A$. Let $U^{\prime}$ be an $R^{\prime}$-supermodule. We said that $U^{\prime}$ is defined over $R$ or $U^{\prime}$ descends to $R$, if there is an $A$-supermodule $U$ such that $U^{\prime}\cong R^{\prime}\otimes_{R}U$ as $A^{\prime}$-supermodules. In this case, by the definition of an isomorphism of $R^{\prime}$-supermodules, we have $U^{\prime}_{\bar{0}}\cong R^{\prime}\otimes_{R}U_{\bar{0}}$ and $U^{\prime}_{\bar{1}}\cong R^{\prime}\otimes_{R}U_{\bar{1}}$ as $A^{\prime}_{\bar{0}}$-modules (recall from Proposition 3.1 that $A^{\prime}_{\bar{0}}=R^{\prime}\otimes_{R}A_{\bar{0}}$). The converse holds if $A$ has a super unit:
Proposition 3.3.
Let $R\subseteq R^{\prime}$ be two commutative rings. Let $A$ be an $R$-superalgebra with a superunit, and let $A^{\prime}:=R^{\prime}\otimes_{R}A$. Let $U^{\prime}$ be an $R^{\prime}$-supermodule. If the $A^{\prime}_{\bar{0}}$-module $U^{\prime}_{\bar{0}}$ descends to $R$, then $U^{\prime}$ descends to $R$.
Proof. By assumption there is an $A_{\bar{0}}$-module $U_{0}$ such that $U^{\prime}_{\bar{0}}\cong R^{\prime}\otimes_{R}U_{\bar{0}}$ as $A^{\prime}_{\bar{0}}$-module. Let $U:=A\otimes_{A_{\bar{0}}}U_{\bar{0}}$. It is easy to check that $U$ is an $A$-supermodule. By [22, Lemma 3.7 (i)], we have $U^{\prime}\cong A^{\prime}\otimes_{A^{\prime}_{\bar{0}}}U^{\prime}_{\bar{0}}$. Hence we have
$U^{\prime}\cong(R^{\prime}\otimes_{R}A)\otimes_{A^{\prime}_{\bar{0}}}(R^{\prime}\otimes_{R}U_{\bar{0}})\cong R^{\prime}\otimes_{R}(A\otimes_{A_{\bar{0}}}U_{\bar{0}})=R^{\prime}\otimes_{R}U$
[…]
Proposition 3.4 (Descent of dual supermodules).
Let $R\subseteq R^{\prime}$ be two commutative rings. Let $A$ be an $R$-superalgebra and let $A^{\prime}:=R^{\prime}\otimes_{R}A$. Let $U$ be an $A$-supermodule which is projective as $R$-module. Then ${\rm Hom}_{R^{\prime}}(R^{\prime}\otimes_{R}U,R^{\prime})\cong R^{\prime}\otimes_{R}{\rm Hom}_{R}(U,R)$ as right $A^{\prime}$-supermodules.
Proof. By [23, Corollary 1.12.12], we have an $R^{\prime}$-module isomorphism
$R^{\prime}\otimes_{R}{\rm Hom}_{R}(U,R)\cong{\rm Hom}_{R^{\prime}}(R^{\prime}\otimes_{R}U,R^{\prime})$
[…]
Proposition 3.5 (Descent of Morita superequivalences).
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$. Let $A$ and $B$ be $\mathcal{O}$-superalgebras, and let $A^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A$, $B^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}B$. Let $M$ be an $(A,B)$-bisupermodule, NN a $(B,A)$-bisupermodule, $M^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}M$ and $N^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}N$. Then $M^{\prime}$ and $N^{\prime}$ induce a Morita superequivalence between $A^{\prime}$ and $B^{\prime}$ if and only if $M$ and $N$ induce a Morita superequivalence between $A$ and $B$. (Note that in this paper, all superspaces are defined to be finitely generated.) A similar statement holds with $\mathcal{O}^{\prime}$ and $\mathcal{O}$ replaced by $k^{\prime}$ and $k$ respectively.
Proof. One direction of the implications is trivial. Now assume that
$M^{\prime}\otimes_{B^{\prime}}N^{\prime}\cong A^{\prime}~~{\rm and}~~N^{\prime}\otimes_{A^{\prime}}M^{\prime}\cong B^{\prime}$ (3.1)
as $(A^{\prime},A^{\prime})$-bisupermodules and $(B^{\prime},B^{\prime})$-bisupermodules, respectively. Then by [19, Proposition 4.5],
$M\otimes_{B}N\cong A~~{\rm and}~~N\otimes_{A}M\cong B$ (3.2)
[…]
Remark 3.6.
Proposition 3.5 can be easily extend to the case for cross products (see 4.1) instead of superalgebras, and graded Morita equivalences (see [26, Definition 3.3]) or graded Rickard equivalences (see 4.1) instead of Morita superequivalences. The proof is almost the same.
3.7.
Projective supermodules and Heller translates. Let $R\in\{\mathcal{O},k\}$ and let $A$ be a superalgebra with a superunit. Following [22, §3.4], we call an $A$-supermodule $P$ a projective supermodule if $P$ is isomorphic to a direct summand of $A^{\oplus n}$ for some $n$. A projective cover of an $A$-supermodule is defined exactly like for modules. Our conditions on the superalgebra $A$ ensure that a projective cover of a finitely generated $A$-supermodule $M$ exists and is unique up to isomorphism. So we can define Heller translate $\Omega_{A}(M)$ of $M$ just like for modules.
Proposition 3.8 (Descent of Heller translates).
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$. Let $A$ be a finite generated $\mathcal{O}$-algebra (resp. $\mathcal{O}$-superalgebra), $A^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A$, and $M^{\prime}$ a finitely generated $A^{\prime}$-module (resp. $A^{\prime}$-supermodule). If $M^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}M$ for some $A$-module (resp. $A$-supermodule) $M$, then $\Omega_{A^{\prime}}(M^{\prime})\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}\Omega_{A}(M)$ as $A^{\prime}$-modules (resp. $A^{\prime}$-supermodules). A similar statement holds with $\mathcal{O}^{\prime}$ and $\mathcal{O}$ replaced by $k^{\prime}$ and $k$ respectively.
[…]
We refer to [23, Theorem 5.2.1] for the definition of Green correspondents. The following result is a descent criterion for Green correspondents.
Proposition 3.9.
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$. Let $G$ be a finite group, $P$ a subgroup of $G$ and $H$ a subgroup of $G$ such that $N_{G}(P)\subseteq H$. Let $U^{\prime}$ be an $\mathcal{O}^{\prime}$-free indecomposable $\mathcal{O}^{\prime}G$-module having $P$ as a vertex, and let $V^{\prime}$ be an $\mathcal{O}^{\prime}H$-module which is the Green correspondent of $U^{\prime}$ with respect to $P$. Assume that $k$ is finite. Then $U^{\prime}$ descends to $\mathcal{O}$ if an only if $V^{\prime}$ descends to $\mathcal{O}$. A similar statement holds for $k^{\prime}$ and $k$ instead of $\mathcal{O}^{\prime}$ and $\mathcal{O}$ respectively.
Proof. Assume that $U^{\prime}$ descends to $\mathcal{O}$, that is, there exists an $\mathcal{O}G$-module $U$ such that $U^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}U$. By definition, $V^{\prime}$ is the unique (up to isomorphism) direct summand of ${\rm Res}_{H}^{G}(U^{\prime})$ with vertex $P$. By [19, Lemma 6.1], we deduce that there is an indecomposable direct summand of $V$ of ${\rm Res}_{P}^{G}(U)$ such that $V^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}V$.
[…]
3.10.
Super group algebra. Throughout the rest of this section $G$ will denote a finite group with an index $2$ subgroup $G_{\bar{0}}$. Let $R$ be a commutative ring. $RG$ has the structure of an $R$-superalgebra with superunit via
$(RG)_{\bar{0}}:=RG_{\bar{0}}~~~{\rm and}~~~(RG)_{\bar{1}}:=\Big\{\sum_{g\in G\backslash G_{\bar{0}}}\alpha_{g}g\mid\alpha_{g}\in R\Big\}.$
For example $G$ and $G_{\bar{0}}$ can be chosen to be $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ ($n\geq 2$) respectively. For any subgroup $H$ of $G$ we will use $H_{\bar{0}}$ to denote $H\cap G_{\bar{0}}$ and $RH$ will inherit its superalgebra structure from that of $RG$.
Let $R\in\{\mathcal{O},k\}$. Following [22], for a module-indecomposable $RG$-supermodule $M$, a vertex of $M$ means a vertex of $|M|$. The same applies to relative projectivity with respect to a subgroup and other standard notions of finite group representation theory.
In [22, Theorem 3.39], Kleshchev and Livesey introduced a super version of the usual Green correspondence:
Theorem 3.11 ([22, Theorem 3.39]).
Let $P$ be a $p$-subgroup of $G$ and $H$ a subgroup of $G$ such that $N_{G}(P)\subseteq H$ and $H\nsubseteq G_{\bar{0}}$. Then there exists a correspondence between the module-indecomposable $RG$-supermodules with a vertex $P$ and the module-indecomposable $RH$-supermodules with a vertex $P$. Suppose that $U$ is a module-indecomposable $RG$-supermodule with a vertex $P$ corresponding to a module-indecomposable $RH$-supermodule $V$ with a vertex $P$. Then
${\rm Ind}_{H}^{G}(V)\cong U\oplus M_{G}~~~{\rm and}~~~{\rm Res}_{H}^{G}(U)\cong V\oplus M_{H}$
for some $RG$-supermodule $M_{G}$ and some $RH$-supermodule $M_{H}$. Moreover, as an $RG$-module, $M_{G}$ is a direct sum of modules each with a vertex contained in $P\cap{}^{g}P$ for some $g\in G\backslash H$ and, as an $RH$-module, $M_{H}$ is a direct sum of modules each with a vertex contained in $H\cap{}^{g}P$ for some $g\in G\backslash H$.
The following is a super version of Proposition 3.9.
Proposition 3.12 (Descent of super Green correspondents).
Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$. Let $G$ be a finite group, $P$ a subgroup of $G$, and $H$ a subgroup of $G$ such that $N_{G}(P)\subseteq H$ and $H\nsubseteq G_{\bar{0}}$. Let $U^{\prime}$ be an $\mathcal{O}^{\prime}$-free module-indecomposable $\mathcal{O}^{\prime}G$-supermodule having $P$ as a vertex, and let $V^{\prime}$ be an $\mathcal{O}^{\prime}H$-module which is the super Green correspondent of $U^{\prime}$. Assume that $k$ is finite. Then $U$ descends to $\mathcal{O}$ if an only if $V$ descends to $\mathcal{O}$. A similar statement holds for $k^{\prime}$ and $k$ instead of $\mathcal{O}^{\prime}$ and $\mathcal{O}$ respectively.
[…]
4 On Rickard equivalences and splendid complexes
4.1.
Let $R$ be a commutative ring and $G$ a finite group. A finitely generated $R$-algebra AA is called $G$-graded (resp. fully $G$-graded) if, as an $R$-module, AA is a direct sum $A=\oplus_{g\in G}A_{g}$ satisfying $A_{g}A_{h}\subseteq A_{gh}$ (resp. $A_{g}A_{h}=A_{gh}$) for all $g,h\in G$; if $A^{\times}\cap A_{g}\neq\varnothing$, then $A$ is called a cross product (of $A_{1}$ and $G$). Let $A$, $B$ and $C$ be fully $G$-graded $R$-algebras. A finitely generated AA-$B$-bimodule $M$ is called $G$-graded if, as an $R$-module, $M$ is a direct sum $M=\oplus_{x\in G}M_{x}$ satisfying $A_{g}M_{x}B_{h}=M_{gxh}$ for all $g,h,x\in G$. Hence with this terminology, if $G=\mathbb{Z}/2\mathbb{Z}$, then $A$ and $B$ are exactly $R$-superalgebras and $M$ is exactly an $(A,B)$-bisupermodule. Following [26], we denote by AA-Gr-$B$ the categroy of $G$-graded finitely generated AA-$B$-bimodules and grade preserving AA-$B$-bimodule homomorphisms. Let $K^{b}(\mbox{$A$-Gr-$B$})$ denote the homotopy category of bound complexes over AA-Gr-$B$; see e.g. [23, page 117] for homotopy categories. Let $M$ be a $G$-graded AA-$B$-bimodule and $N$ a $G$-graded $B$-$C$-bimodule. Then $M\otimes_{B}N$ is a $G$-graded AA-$C$-bimodule with $(M\otimes_{B}N)_{x}=\sum_{gh=x}M_{g}\otimes_{B}N_{h}$. Hence for an object $X$ of $K^{b}(\mbox{$A$-Gr-$B$})$ and an object $Y$ of $K^{b}(\mbox{$B$-Gr-$C$})$, the complex $X\otimes_{B}Y$ is an object of $K^{b}(\mbox{$A$-Gr-$C$})$.
Let $A$ and $B$ be symmetric $R$-algebras. If $X$ is a bounded complex of AA-$B$-bimodules that are finitely generated projective as left $A$-modules and as right $B$-modules, such that $X\otimes_{B}X^{*}$ and $A$ are homotopy equivalent as complexes of AA-AA-bimomdules, and that $X^{*}\otimes_{A}X$ and $B$ are homotopy equivalent as complexes of $B$-$B$-bimodules. Then we say that $A$ and $B$ are Rickard equivalent, and $X$ is called a Rickard complex of AA-$B$-bimodules. Here $A$ (resp. $B$) denotes the complex with only one nonzero term $A$ (resp. $B$) in degree 0, and we adopt [23, Definition 1.17.9] for the definition of tensor product of complexes.
Let $A$, $B$ and $C$ be fully $G$-graded symmetric $R$-algebras. If there is a complex $X\in K^{b}(\mbox{$A$-Gr-$B$})$ where each term of $X$ is finitely generated projective as left $A$-module and as right $B$-module, such that $X\otimes_{B}X^{*}$ and $A$ are isomorphic in $K^{b}(\mbox{$A$-Gr-$A$})$ and $X^{*}\otimes_{A}X$ and $B$ are isomorphic in $K^{b}(\mbox{$B$-Gr-$B$})$. Then $A$ and $B$ are graded Rickard equivalent.
Using [22, Lemmas 3.6 and 3.19], it is straightforward to check the following:
Proposition 4.2.
Let $R$ be a commutative ring, and let $A_{1},A_{2},B_{1},B_{2}$ be superalgebras (in other words, $\mathbb{Z}/2\mathbb{Z}$-graded algebras). If $A_{i}$ and $B_{i}$ are $\mathbb{Z}/2\mathbb{Z}$-graded Rickard equivalent via the complexes $X_{i}$ of $(A_{i},B_{i})$-bisupermodules (in other words, $\mathbb{Z}/2\mathbb{Z}$-graded $A_{i}$-$B_{i}$-bimodules), for $i=1,2$, then $A_{1}\otimes_{R}A_{2}$ and $B_{1}\otimes_{R}B_{2}$ are $\mathbb{Z}/2\mathbb{Z}$-graded Rickard equivalent via the complex $X_{1}\boxtimes_{R}X_{2}$ of $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bisupermodules.
Let $R$ be a commutative ring, $A$ and $B$ be fully $G$-graded $R$-algebras. An $(A,B)$-bimodule can be regarded as a module of the $R$-algebra $A\otimes_{R}B^{\rm op}$. We regard $B^{\rm op}$ as a $G$-graded algebra with components $B^{\rm op}_{g}=B_{g^{-1}}$, and $A\otimes_{R}B^{\rm op}$ as a $(G\times G)$-graded $R$-algebra with $(A\otimes_{R}B)_{(g,h)}=A_{g}\otimes_{R}B^{\rm op}_{h}$. Let $\Delta(A,B):=\oplus_{g\in G}A_{g}\otimes_{R}B^{\rm op}_{g}$, which is a fully $G$-graded $R$-algebra and which is also an $R$-subalgebra of $A\otimes_{R}B^{\rm op}$.
Lemma 4.3 ([26, Lemma 2.6]).
Let $R$ be a commutative ring, $G$ a finite group, $A$ and $B$ fully $G$-graded $R$-algebras. There are mutually inverse functors induce a equivalence between the categories $\Delta(A,B)$-mod and AA-Gr-$B$. In particular, the categories $K^{b}(\mbox{$\Delta(A,B)$-{\rm mod}})$ and $K^{b}(\mbox{$A$-{\rm Gr}-$B$})$ are equivalent.
An explicit pair of mutually inverse functors which induces the equivalence in Lemma 4.3 has been given in [26, Lemma 2.6]. The following theorem is a homotopy category version of [26, Theorem 4.7], which is, as [26, Theorem 4.7], a straightforward consequence of [26, 4.2, 4.3].
Theorem 4.4.
Let $R$ be a commutative ring, $G$ a finite group, $A$ and $B$ fully $G$-graded symmetric $R$-algebras. The following are equivalent:
  1. 1.
    There is an object $X\in K^{b}(\mbox{$A$-{\rm Gr}-$B$})$ inducing a graded Rickard equivalent between $A$ and $B$.
  2. 2.
    There is an object $X_{1}\in K^{b}(\Delta(A,B))$, and isomorphisms $\alpha_{1}:X_{1}\otimes_{B_{1}}X_{1}^{*}\to A_{1}$ in $K^{b}(\Delta(A,A))$ and $\beta_{1}:X_{1}^{*}\otimes_{A_{1}}X_{1}\to B_{1}$ in $K^{b}(\Delta(B,B))$.
Moreover, if the above conditions hold, then $X$ and $X_{1}$ correspond to each other via the category equivalence in Lemma 4.3.
If we only know that a Rickard complex for $A_{1}$ and $B_{1}$ extends to $\Delta(A,B)$, then we need $|G|$ to be invertible in $R$:
Theorem 4.5 (A homotopy category version of [26, Theorem 4.8]).
Let $R$ be a commutative ring, $G$ a finite group, $A$ and $B$ fully $G$-graded symmetric $R$-algebras. Assume that $|G|$ is invertible in $R$. Then the following are equivalent:
  1. 1.
    There is an object $X\in K^{b}(\mbox{$A$-{\rm Gr}-$B$})$ inducing a graded Rickard equivalent between $A$ and $B$.
  2. 2.
    There is an object $X_{1}\in K^{b}(\Delta(A,B))$ which, regarded as an object of $K^{b}((A_{1}\otimes_{R}B_{1}^{\rm op}))$, induces a Rickard equivalence between $A_{1}$ and $B_{1}$.
Moreover, if above conditions hold, then $X$ and $X_{1}$ are correspond to each other via the category equivalence in Lemma 4.3.
[…]
4.6.
Let $G$ and $H$ be finite groups. An $(\mathcal{O}G,\mathcal{O}H)$-bimodule $M$ can be regarded as an $\mathcal{O}(G\times H)$-module (and vice versa) via $(g,h)m=gmh^{-1}$, where $g\in G$, $h\in H$ and $m\in M$. If $M$ is indecomposable as an $(\mathcal{O}G,\mathcal{O}H)$-bimodule, then $M$ is indecomposable as an $\mathcal{O}(G\times H)$-module, hence has a vertex (in $G\times H$) and a source.
Since Rickard’s paper [30], it has been customary to use the adjective “splendid” to any of the Rickard/derived or Morita equivalences. In [1, Remark 1.2], Boltje defined a more general version of splendid Rickard equivalences, which is more flexible and preserves all the good properties of the original version of Rickard [30] and also later definitions (see for instance [24, Definition 9.7.5]). We adopt Boltje’s definition. If $\varphi:P\xrightarrow{\cong}Q$ is an isomorphism between subgroups $P\leq G$ and $Q\leq H$, we set
$\Delta\varphi:=\{(u,\varphi(u))\mid u\in P\},$
and call it a twisted diagonal subgroup of $G\times H$. Let $R\in\{\mathcal{O},k\}$. Whenever useful, we regard an $R\Delta\varphi$-module as an $RP$-module and vice versa via the isomorphism $P\cong\Delta\varphi$ sending $u\in P$ to $(u,\varphi(u))$. We say that an RGRG-RHRH-bimodule $M$ is splendid if every indecomposable direct summand of $M$, regarded as an $R(G\times H)$-module, is a $p$-permutation module with twisted diagonal vertex. We say that a complex $X=(X_{i})_{i\in\mathbb{Z}}$ of RGRG-RHRH-bimodule is splendid if $X_{i}$ is splendid for any $i\in\mathbb{Z}$. Using the Mackey formula, one checks that a splendid bimodule is projective as left module and as right module. With this definition of “splendid”, splendid modules and splendid complexes are closed under induction and restriction:
Proposition 4.7.
Let $R\in\{\mathcal{O},k\}$. Let $G$ and $H$ be finite groups, $G_{1}$ a subgroup of $G$ and $H_{1}$ a subgroup of $H$.
  1. 1.
    If $M$ is a splendid (complex of) $RG_{1}$-$RH_{1}$-bimodule(s), then ${\rm Ind}_{G_{1}\times H_{1}}^{G\times H}(M)$ is a splendid (complex of) RGRG-RHRH-bimodule(s).
  2. 2.
    If $M$ is a splendid (complex of) RGRG-RHRH-bimodule(s), then ${\rm Res}_{G_{1}\times H_{1}}^{G\times H}(M)$ is a splendid (complex of) $RG_{1}$-$RH_{1}$-bimodule(s).
  3. 3.
    Assume that $p\nmid|G\times H:G_{1}\times H_{1}|$ and $M$ is a/an (complex of) RGRG-RHRH-bimodule(s). If ${\rm Res}_{G_{1}\times H_{1}}^{G\times H}(M)$ is splendid, then $M$ is splendid.
Proof. It suffices to consider bimodules. If $M$ is a $RG_{1}$-$RH_{1}$-bimodule, then a vertex of ${\rm Ind}_{G_{1}\times H_{1}}^{G\times H}(M)$ is contained in a vertex of $M$, and this proves (i). One easily checks statement (ii) by using the Mackey formula. It remains to prove statement (iii). Since $p\nmid|G\times H:G_{1}\times H_{1}|$, $M$ is relatively $G_{1}\times H_{1}$-projective, hence $M$ is isomorphic to a direct summand of
${\rm Ind}_{G_{1}\times H_{1}}^{G\times H}{\rm Res}_{G_{1}\times H_{1}}^{G\times H}(M);$
[…]
With the definition of “splendid” in 4.6, tensor products and duals of splendid bimodules/complexes are splendid:
Proposition 4.8.
Let $R\in\{\mathcal{O},k\}$. Let $G$, $H$ and $L$ finite groups. If $M$ is a splendid RGRG-RHRH-bimodule and $N$ is a splendid RHRH-RCRC-bimodule, then $M\otimes_{RH}N$ is a splendid RGRG-RLRL-bimodule, and $M^{*}:={\rm Hom}_{R}(M,R)$ is a splendid RHRH-RGRG-bimodule.
[…]
Propositions 4.7 (ii) and 4.8 explain why the definition of “splendid” in 4.6 is more flexible than the versions in [30] and [24, Definition 9.7.5].
Definition 4.9.
Let $R\in\{\mathcal{O},k\}$. Let $G$ and $H$ be finite groups, $b\in Z(RG)$ and $c\in Z(RH)$ be nonzero idempotents. If there is a bounded splendid complex $X$ of finitely generated RGbRGb-RHcRHc-bimodules inducing a Rickard equivalence between $RGb$ and $RHc$. Then we say that $RGb$ and $RHc$ are splendidly Rickard equivalent, and $X$ is called a splendid Rickard complex.
4.10.
Let $R\in\{\mathcal{O},k\}$, and let $G$ and $H$ be finite groups. By definition, an RGRG-RHRH module $M$ is splendid if it is a direct sum of summands to of the $R(G\times H)$-bimodules ${\rm Ind}_{\Delta\varphi}^{G\times H}(R)$, with $\Delta\varphi$ running over the twisted diagonal subgroups of $G\times H$. For any $\varphi:P\cong Q$, note that we have an isomorphism of RGRG-RHRH-bimodules
${\rm Ind}_{\Delta\varphi}^{G\times H}(R)\cong RG\otimes_{RP}{}_{\varphi}RH\cong RG_{\varphi^{-1}}\otimes_{RQ}RH$
sending $(x,y)\otimes 1$ to $(x,y^{-1})$ for all $x\in G$ and $y\in H$. Here $_{\varphi}RH$ is the RPRP-RHRH-bimodule which is equal to $RH$ as a right $RH$-module endowed with the left structure homomorphism $P\xrightarrow{\varphi}Q\to{\rm End}_{R}(RH)$; the notation $RG_{\varphi^{-1}}$ is similarly defined.
More generally, as in [11, Definition 1.3], we can define splendid modules, complexes, and Rickard equivalences for symmetric interior $P$-algebras:
Definition 4.11.
Let $R\in\{\mathcal{O},k\}$ and let $P$ be a finite $p$-group. Let $A$ and $B$ be symmetric interior $P$-algebras over $R$. An AA-$B$-bimodule $M$ is called splendid if it is a direct sum of summands of the AA-$B$-bimodule $A\otimes_{RQ}{}_{\varphi}B\cong A_{\varphi^{-1}}\otimes_{RT}B$, with $Q$, $T$ running over subgroups of $P$ and $\varphi:Q\to T$ running over the set of isomorphisms between $Q$ and $T$. We say that a complex $X=(X_{i})_{i\in\mathbb{Z}}$ of AA-$B$-bimodule is splendid if $X_{i}$ is splendid for any $i\in\mathbb{Z}$. If there is a bounded splendid complex $X$ of finitely generated AA-$B$-bimodules inducing a Rickard equivalence between $A$ and $B$. Then we say that $A$ and $B$ are splendidly Rickard equivalent, and $X$ is called a splendid Rickard complex.
Proposition 4.12.
Let $R\in\{\mathcal{O},k\}$. Let $P$ be a finite $p$-group, and let $A_{1},A_{2},B_{1},B_{2}$ be interior $P$-algebras over $R$, which are also $R$-superalgebras with the images of $P$ contained in their even parts $(A_{1})_{\bar{0}},(A_{2})_{\bar{0}},(B_{1})_{\bar{0}},(B_{2})_{\bar{0}}$. If $M_{i}$ is a splendid (complex of) $A_{i}$-$B_{i}$-bimodule(s) which is also a/an (complex of) $A_{i}$-$B_{i}$-bisupermodule, for $i=1,2$, then $M_{1}\boxtimes_{R}M_{2}$ is a splendid (complex of) $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bimodule(s).
Proof. It suffices to consider bimodules. Let $Q_{1},Q_{2},T_{1},T_{2}$ be subgroups of $P$ and let $\varphi_{1}:Q_{1}\cong T_{1}$, $\varphi_{2}:Q_{2}\to T_{2}$ be group isomorphisms. It suffices to show that
$(A_{1}\otimes_{RQ_{1}}{}_{\varphi_{1}}B_{1})\boxtimes_{R}(A_{2}\otimes_{RQ_{2}}{}_{\varphi_{2}}B_{2})$
is a splendid $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bimodule. By [22, Lemma 3.6], we have an isomorphism of $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bimodules
[…]
5 On spin blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$
Notation 5.1.
Let $n$ be a positive integer. We use $[n]$ to denote the set $\{1,\cdots,n\}$. For conventional convenience we set $0]:=\emptyset$. We denote by $\mathscr{P}(n)$ the set of partitions of $n$ and by $\mathscr{P}_{0}(n)$ the set of strict partitions of $n$, i.e. partitions of $n$ without repeat parts. Set $\mathscr{P}_{0}(0):=\{{\rm empty~partion}\}$. For $\lambda\in\mathscr{P}_{0}(n)$, ${\rm Ab}_{\lambda}$ will denote its $\bar{p}$-abacus ($p$-bar-abacus) display; see [21, §2.3.b] for more details on this. For any non-empty partition $\lambda$, set $h(\lambda):={\rm max}\{k~|~\lambda_{k}>0\}$ to be the length of $\lambda$. We adopt the convention that the 0-th position on the abacus display is always unoccupied, i.e. we always use $h(\lambda)$ beads in ${\rm Ab}_{\lambda}$. Let $\mathscr{P}_{0}:=\bigsqcup_{n\in\mathbb{N}}\mathscr{P}_{0}(n)$. Following [29], we can associate to every $\lambda\in\mathscr{P}_{0}$ its $\bar{p}$-core ${\rm core}(\lambda)\in\mathscr{P}_{0}$. We can also define a non-negative integer called the $\bar{p}$-weight of $\lambda$: ${\rm wt}(\lambda):=(|\lambda|-|{\rm core}(\lambda)|)/p\in\mathbb{N}$.
Assume in this section that $(K,\mathcal{O},k)$ is large enough for $\widetilde{S}_{n}$; this means that $K$ (and hence $\mathcal{O}$) contains a primitive $(2n!)$-th root of unity. We also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$. Let $R\in\{\mathcal{O},k\}$, and $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$. If $n>1$, we consider $R\widetilde{S}_{n}$ as a superalgebra corresponding to $(\widetilde{S}_{n})_{\bar{0}}=\widetilde{A}_{n}$; see 3.10. We set
$e_{z}:=(1-z)/2\in R\widetilde{S}_{n}.$
We will use the adjective “spin” when referring to objects associated to $R\widetilde{S}_{n}e_{z}$ rather than the whole of $R\widetilde{S}_{n}$. For example the spin characters of $\widetilde{S}_{n}$ will refer to the $K$-characters $\chi$ of $\widetilde{S}_{n}$ satisfying $\chi(z)=-\chi(1)$, while spin blocks refers to the blocks of $R\widetilde{S}_{n}e_{z}$. Note that $R\widetilde{S}_{n}e_{z}$ inherits the structure of superalgebra from $R\widetilde{S}_{n}$ since the idempotent $e_{z}$ is in $(R\widetilde{S}_{n})_{\bar{0}}$.
Let $G\leq\widetilde{S}_{n}$ with $z\in G$. Following the notation from 3.10, we set
$G_{\bar{0}}:=G\cap\widetilde{A}_{n}\quad\text{and}\quad G_{\bar{1}}:=G\setminus\widetilde{A}_{n}$ (5.1)
and treat $RGe_{z}$ as a superalgebra in the appropriate way. (Note that unlike in 3.10 we are yet not assuming that $G_{\bar{0}}$ is a proper subgroup of $G$.)
As stated in [22, (4.3)], we have the following isomorphisms of $R$-superalgebras (see 2.7 for the notation $\mathcal{T}_{n}^{R}$):
$\begin{split}R\widetilde{S}_{n}^{+}e_{z}&\cong\mathcal{T}_{n}^{R},~t_{i}e_{z}\mapsto\tau_{i},\ ze_{z}\mapsto-1\\ R\widetilde{S}_{n}^{-}e_{z}&\cong\mathcal{T}_{n}^{R},~t_{i}e_{z}\mapsto(-1)^{i}\sqrt{-1}\tau_{i},\ ze_{z}\mapsto-1.\end{split}$ (5.2)
Remark 5.2.
If $p\equiv 1~{\rm mod}~4$, then $\mathbb{Z}_{p}$ (resp. $\mathbb{F}_{p}$) contains a primitive $4$-th root of unity. Since $R$ is a flat $R_{p}$-module (see [19, Lemma 4.1]), the $R$-superalgebra isomorphisms (5.2) restrict to isomorphisms of $R_{p}$-superalgebras
$\begin{split}R_{p}\widetilde{S}_{n}^{+}e_{z}&\cong\mathcal{T}_{n}^{R_{p}},~t_{i}e_{z}\mapsto\tau_{i},\ ze_{z}\mapsto-1\\ R_{p}\widetilde{S}_{n}^{-}e_{z}&\cong\mathcal{T}_{n}^{R_{p}},~t_{i}e_{z}\mapsto(-1)^{i}\sqrt{-1}\tau_{i},\ ze_{z}\mapsto-1.\end{split}$
If $p\equiv 3~{\rm mod}~4$, we still have the first $R_{p}$-superalgebra isomorphism but the second does not hold, because this time $R_{p}$ does not contain any primitive $4$-th root of unity.
5.3.
[…]
As in [22], if $X\subseteq[n]$, then $S_{X}$ will signify the subgroup of $S_{n}$ consisting of all elements that fix $\X[n]\backslash$ X pointwise and $\widetilde{S}_{X}$ will denote $\pi_{n}^{-1}(S_{X})\leq\widetilde{S}_{n}$. Moreover, for disjoint $X_{1},\cdots,X_{i}\subseteq[n]$, let $\widetilde{S}_{X_{1},\cdots,X_{r}}$ be the subgroup of $\widetilde{S}_{n}$ generated by $\widetilde{S}_{X_{1}},\cdots,\widetilde{S}_{X_{i}}$. In particular,
$\widetilde{S}_{X_{1},\cdots,X_{i}}\cong\widetilde{S}_{X_{1}}\times_{z}\cdots\times_{z}\widetilde{S}_{X_{i}}$
and
$R\widetilde{S}_{X_{1},\cdots,X_{i}}e_{z}\cong R\widetilde{S}_{X_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}_{X_{i}}e_{z}\cong\mathcal{T}_{|X_{1}|}^{R}\otimes_{R}\cdots\otimes_{R}\mathcal{T}_{|X_{i}|}^{R}$ (5.3)
as superalgebras. We also define the corresponding subgroups of $\widetilde{A}_{n}$:
$\widetilde{A}_{X}:=\widetilde{S}_{X}\cap\widetilde{A}_{n}\quad\text{and}\quad\widetilde{A}_{X_{1},\cdots,X_{i}}:=\widetilde{S}_{X_{1},\cdots,X_{i}}\cap\widetilde{A}_{n}.$
In the important special case where $n=n_{1}+\cdots+n_{i}$ and $Y_{l}=[n_{1}+\cdots+n_{l}]\setminus[n_{1}+\cdots+n_{l-1}]$ for $l=1,\cdots,r$, we have the standard Young subgroups
$\widetilde{S}_{n_{1},\cdots,n_{i}}:=\widetilde{S}_{Y_{1},\cdots,Y_{i}}\cong\widetilde{S}_{n_{1}}\times_{z}\cdots\times_{z}\widetilde{S}_{n_{i}}\leq\widetilde{S}_{n}\quad\text{and}\quad\widetilde{A}_{n_{1},\cdots,n_{i}}=\widetilde{S}_{n_{1},\cdots,n_{i}}\cap\widetilde{A}_{n}\leq\widetilde{A}_{n}.$
If some of the $n_{l}$’s are equal to each other, we use the exponential notation, as for example for the standard Young subgroup $\widetilde{S}_{r+ip,p^{d-i}}=\widetilde{S}_{r+ip,p,\cdots,p}$.
Following Schur [32], to every strict $\lambda\in\mathscr{P}_{0}(n)$, one associated canonically a spin character $\chi_{\lambda}$ of $\widetilde{S}_{n}$ such that $\chi_{\lambda}$ is irreducible if $n-h(\lambda)$ is even and $\chi_{\lambda}=\chi_{\lambda}^{+}+\chi_{\lambda}^{-}$ for irreducible characters $\chi_{\lambda}^{\pm}$ if $n-h(\lambda)$ is odd. Moreover,
$\{\chi_{\lambda}\mid\lambda\in\mathscr{P}_{0}(n),~n-h(\lambda)~{\rm is~even}\}\cup\{\chi_{\lambda}^{\pm}\mid\lambda\in\mathscr{P}_{0}(n),~n-h(\lambda)~{\rm is~odd}\}$
is a complete irredundant set of irreducible spin characters of $\widetilde{S}_{n}$.
Theorem 5.4 ([17], see also [21, Theorem 5.2.4]).
With one exception, $\chi_{\lambda}^{\pm}$ and $\chi_{\mu}^{\pm}$ are in the same $p$-block if and only if ${\rm core}(\lambda)={\rm core}(\mu)$. The exception is $\lambda=\mu$ is a $\bar{p}$-core and $n-h(\lambda)$ is odd, in which case $\chi_{\lambda}^{+}$ and $\chi_{\lambda}^{-}$ are in different $p$-blocks.
The blocks of $R\widetilde{S}_{n}$ is described by the following theorem:
Theorem 5.5 ([17, Theorem 1.1], see also [22, Theorem 4.3]).
  1. 1.
    The spin blocks of $R\widetilde{S}_{n}$ are labelled by pairs $(\rho,d)$, where $\rho$ is a $\bar{p}$-core and $d$ is a non-negative integer with $n=|\rho|+dp$. If $d=0$ and $n-h(\rho)$ is odd then $(\rho,0)$ labels two spin blocks of $R\widetilde{S}_{n}$ with corresponding block idempotents $e_{\rho,0}^{+},e_{\rho,0}^{-}\in R\widetilde{S}_{n}e_{z}$. In all cases we denote the block idempotent (or sum of two block idempotents) corresponding to $(\rho,d)$ by $e_{\rho,d}\in R\widetilde{S}_{n}e_{z}$.
  2. 2.
    The character $\chi_{\lambda}^{\pm}$ lies in the (sum of) block(s) $\mathcal{O}\widetilde{S}_{n}e_{\rho,d}$, where $\rho$ is the $\bar{p}$-core of $\lambda$. If $d=0$ and $\rho$ is odd, we choose the labeling so that $\chi_{\rho}^{\pm}$ lies in the block $\mathcal{O}\widetilde{S}_{n}e_{\rho,0}^{\pm}$.
Defect groups and Brauer correspondents of blocks of $R\widetilde{S}_{n}$ are described as follows:
Theorem 5.6 ([5, Theorem A and Corollary 26], see also [22, Theorem 4.3]).
Keep the notation of Theorem 5.5. Fix a $\bar{p}$-core $\rho$ and $d$ a non-negative integer with $n=|\rho|+dp$, and set $U:=[|\rho|]$ and $V:=[n]\setminus U$.
  1. (i)
    If $d=0$, then each block of $R\widetilde{S}_{n}e_{\rho,d}$ has trivial defect group. If $d>0$, then any Sylow $p$-subgroup $D$ of $\widetilde{S}_{V}\leq\widetilde{S}_{n}$ is a defect group of $R\widetilde{S}_{n}e_{\rho,d}$.
  2. (ii)
    If $d=0$, then each block of $R\widetilde{S}_{n}e_{\rho,d}$ is its own Brauer correspondent. If $d>0$, then, setting $D$ to be as in (i), the Brauer correspondent of $R\widetilde{S}_{n}e_{\rho,d}$ in $N_{\widetilde{S}_{n}}(D)$ is $RN_{\widetilde{S}_{n}}(D)e_{\rho,0}$. Here we view $e_{\rho,0}$ as an element of $RN_{\widetilde{S}_{n}}(D)$ via $e_{\rho,0}\in R\widetilde{S}_{U}\hookrightarrow RN_{\widetilde{S}_{n}}(D)$. If $\rho=\varnothing$, we interpret $e_{\varnothing,0}$ as $e_{z}$.
Remark 5.7 ([22, Remark 4.4]).
The $R\widetilde{S}_{n}e_{\rho,d}$’s are precisely the superblocks of $R\widetilde{S}_{n}e_{z}$. That is, the finest decomposition of $R\widetilde{S}_{n}e_{z}$ into a direct sum of two-sided ideals that are invariant under the involution $\sigma_{R\widetilde{S}_{n}}$ (see (2.1) for this notation). This follows from the distribution of characters given in Theorem 5.5 (ii). In particular, we always have $e_{\rho,d}\in R\widetilde{A}_{n}e_{z}$.
The spin blocks of $R\widetilde{A}_{n}$ are described in [18, Proposition 3.16]:
Theorem 5.8.
For each $n>1$, $\bar{p}$-core $\rho$ and $d$ a non-negative integer with $n=|\rho|+dp$, $R\widetilde{A}_{n}e_{\rho,d}$ is a single block of $R\widetilde{A}_{n}$ unless $d=0$ and $\rho$ is even. In this latter case $R\widetilde{A}_{n}e_{\rho,d}$ is a direct sum of two blocks of $R\widetilde{A}_{n}$. If $n=1$, then $R\widetilde{A}_{1}e_{(1),0}=R\widetilde{A}_{1}e_{(1),0}=Re_{z}$ is a single block of $R\widetilde{A}_{1}$.
If $d=0$, once again, the defect group is trivial. In all other cases the defect group of $R\widetilde{A}_{n}e_{\rho,d}$ is the same as that of $R\widetilde{S}_{n}e_{\rho,d}$.
We set
$B^{\rho,d}_{R}:=R\widetilde{S}_{n}e_{\rho,d}~~{\rm and}~~(B^{\rho,d}_{R})_{\bar{0}}:=R\widetilde{A}_{n}e_{\rho,d}.$
It is easy to see that $R\widetilde{A}_{n}e_{\rho,d}$ is indeed the purely even subalgebra of the superalgebra $R\widetilde{S}_{n}e_{\rho,d}$. We will often assume $d>0$ to ensure that blocks and superblocks coincide. However, the corresponding statements for trivial defect blocks are completely elementary.
Proposition 5.9.
Keep the notation of Theorem 5.5. The central idempotent $e_{\rho,d}$ of $R\widetilde{S}_{n}$ is contained in the subring $R_{p}\widetilde{S}_{n}$, hence also contained in $R_{p}\widetilde{A}_{n}$ by Remark 5.7.
This proposition can be deduced from the following lemma, which is a slight generalisation of [16, Proposition 2.2].
Lemma 5.10.
Let $G$ be a finite group, $b$ a central idempotent of $\mathcal{O}G$ and let $\chi:G\to K$ be the character of of a $KGb$-module. Let $\bar{b}$ be the image of $b$ in $kG$. If the values of $\chi$ are contained in $\mathbb{Q}$ (hence in $\mathbb{Z}$), then we have $b\in\mathbb{Z}_{p}G$ and $\bar{b}\in\mathbb{F}_{p}G$.
[…]
We set
$B^{\rho,d}_{R_{p}}:=R_{p}\widetilde{S}_{n}e_{\rho,d}~~{\rm and}~~(B^{\rho,d}_{R_{p}})_{\bar{0}}:=R_{p}\widetilde{A}_{n}e_{\rho,d}.$
As before, it is easy to see that $R_{p}\widetilde{A}_{n}e_{\rho,d}$ is indeed the purely even subalgebra of the superalgebra $R_{p}\widetilde{S}_{n}e_{\rho,d}$.
Fix a $\bar{p}$-core $\rho$ and a non-negative integer $d$. Let $r:=|\rho|$ and $n:=r+dp$. As in Theorem 5.5, we set $U:=[r]$ and $V:=[n]\setminus U$. For $i\in\{1,\cdots,d\}$ we also set $V_{i}:=[r+ip]\setminus[r+(i-1)p].$ In other words, we have:
$\underbrace{1,\cdots,r}_{U}\,\overbrace{\underbrace{r+1,\cdots,r+p}_{V_{1}}\,\,\underbrace{r+p+1,\dots,r+2p}_{V_{2}}\,\cdots\,\underbrace{r+p(d-1)+1,\cdots,r+dp}_{V_{d}}}^{V}$ (5.4)
Let $D_{i}$ is a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$. With this notation, the following now follows easily from Theorem 5.5:
Lemma 5.11 ([22, Lemma 4.6]).
Let $D$ be a defect group of the block algebra $R\widetilde{S}_{n}e_{\rho,d}$. Then $D$ is abelian if and only if $d<p$. In this case we can choose $D=D_{1}\times\dots\times D_{d}$.
5.12.
Keep the notation in (5.4). Let $\iota:S_{d}\to S_{V}$ be the natural permutation action of $S_{d}$ on the set $\{P_{1},\cdots,P_{d}\}$: more precisely, for any $1\leq i\leq d$, $1\leq j\leq p$ and $w\in S_{d}$
$\iota(w)\cdot(r+(i-1)p+j)=r+(w(i)-1)p+j.$
For each $w\in S_{d}$, choose a lift $T_{w}$ of $\iota(w)$ to $\widetilde{S}_{V}$, and set $T_{i}:=T_{s_{i}}$ for all $1\leq i\leq d-1$.
Note that every Sylow $p$-subgroup of $S_{V_{i}}$ lifts uniquely and isomorphically to a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$. Hence we can choose the $D_{i}$’s such that $T_{w}D_{i}T_{w}^{-1}=D_{w(i)}$, for all $1\leq i\leq d$ and $w\in S_{d}$. Indeed, one can first construct Sylow $p$-subgroups of each of the $S_{V_{i}}$’s that get permuted by the $\iota(w)$’s and then lift to $\widetilde{S}_{V_{i}}$. By the construction of the $T_{w}$’s we also have $T_{w}\widetilde{S}_{V_{i}}T_{w}^{-1}=\widetilde{S}_{V_{w(i)}}$, for all $1\leq i\leq d$ and $w\in S_{d}$.
Note that $R\widetilde{S}_{V_{1}}e_{z}\cong R\widetilde{S}_{p}e_{z}$ via $t_{r+m}e_{z}\mapsto t_{m}e_{z}$. We now fix isomorphisms between the $R\widetilde{S}_{V_{i}}e_{z}$’s. For each $1<i\leq d$, we identify $R\widetilde{S}_{V_{i}}e_{z}$ with $R\widetilde{S}_{V_{1}}e_{z}\cong R\widetilde{S}_{p}e_{z}$ via the isomorphism
$R\widetilde{S}_{V_{1}}e_{z}\to R\widetilde{S}_{V_{i}}e_{z},\quad a\mapsto(-1)^{i|a|}T_{(1,i)}aT_{(1,i)}^{-1}.$ (5.5)
(This does not depend on our choice of $T_{(1,i)}$, because $T_{(1,i)}$ is uniquely determined up to multiplication by $z$.) Through these isomorphisms and using (5.3), we can identify the superalgebra
$R\widetilde{S}_{V_{1},\cdots,V_{d}}e_{z}\cong R\widetilde{S}_{V_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}_{V_{d}}e_{z}$ (5.6)
with $(R\widetilde{S}_{p}e_{z})^{\otimes d}$. Clearly (5.5) and (5.6) still hold with $R$ replaced by $R_{p}$ - because all coefficients involved in these isomorphisms are contained in $R_{p}$.
5.13.
Fix a $\bar{p}$-core $\rho$ and a non-negative integer $d$. We adopt the notation in (5.4), in particular, $r:=|\rho|$ and $n:=r+dp$. Let
$G$ $:=\widetilde{S}_{n},$
$L$ $:=\widetilde{S}_{r,p^{d}},$
$N^{\rho,d}$ $:=N_{G}(\widetilde{S}_{V_{1},\cdots,V_{d}}).$
Note that $N^{\rho,d}$ must permute $V_{1},\cdots,V_{d}$. In particular, $N^{\rho,d}\leq\widetilde{S}_{n}$ is the subgroup generated by $L=\widetilde{S}_{U,V_{1},\cdots,V_{d}}$ and the set $\{T_{w}\mid w\in S_{n}\}$ where the notation $T_{w}$ is from 5.12. Recall the notation in (5.1) that we will also use the subgroups $G_{\bar{0}}=\widetilde{A}_{n}$, $N_{\bar{0}}^{\rho,d}=N^{\rho,d}\cap\widetilde{A}_{n}$, etc. We make the following identifications using (5.3):
$\begin{split}RLe_{z}&\cong R\widetilde{S}_{U}e_{z}\otimes_{R}R\widetilde{S}_{V_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}_{V_{d}}e_{z}.\end{split}$ (5.7)
Set $b:=e_{\rho,d}\in RGe_{z}$ and
$f=:e_{\rho,0}\otimes e_{\varnothing,1}^{(1)}\otimes\cdots\otimes e_{\varnothing,1}^{(d)}\in RLe_{z}\subseteq RN^{\rho,d}e_{z}.$
By Proposition 5.9, we have $f\in R_{p}L\subseteq R_{p}N^{\rho,d}$. By [22, page 128], the isomorphism
$RLe_{z}\cong R\widetilde{S}_{r}e_{z}\otimes_{R}(R\widetilde{S}_{p}e_{z})^{\otimes d}$
from (5.7) extends to an isomorphism
$RN^{\rho,d}e_{z}\cong R\widetilde{S}_{r}e_{z}\otimes_{R}(R\widetilde{S}_{p}e_{z}\wr_{s}\mathcal{T}_{d}^{R}),\ T_{i}e_{z}\mapsto\kappa_{i}^{-1}(1\otimes\tau_{i}),$
where $\kappa_{i}$’s are from [22, Lemma 4.7]. More specifically, according to the proof of [22, Lemma 4.7], if $\widetilde{S}_{n}=\widetilde{S}_{n}^{+}$, then $\kappa_{i}\in\{1,-1\}$ for all $1\leq i\leq d-1$, and if $\widetilde{S}_{n}=\widetilde{S}_{n}^{-}$, then $\kappa_{i}\in\{\sqrt{-1},-\sqrt{-1}\}$ for all $1\leq i\leq d-1$. In particular, we have the isomorphisms of superalgebras
$\begin{split}RLf&\cong R\widetilde{S}_{r}e_{\rho,0}\otimes_{R}(R\widetilde{S}_{p}e_{\varnothing,1})^{\otimes d}\cong B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\otimes d}\\ RN^{\rho,d}f&\cong R\widetilde{S}_{r}e_{\rho,0}\otimes_{R}(R\widetilde{S}_{p}e_{\varnothing,1}\wr_{s}\mathcal{T}_{d}^{R})\cong B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R}).\end{split}$ (5.8)
Remark 5.14.
For the group $\widetilde{S}_{n}^{+}$, since $\kappa_{i}=\pm 1\in R_{p}$ and since $R$ is a flat $R_{p}$-module (see [19, Lemma 4.1]), the isomorphisms (5.8) of $R$-superalgebras restricts to isomorphisms of $R_{p}$-superalgebras
$\begin{split}R_{p}Lf&\cong R_{p}\widetilde{S}_{r}^{+}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{+}e_{\varnothing,1})^{\otimes d}\\ R_{p}N^{\rho,d}f&\cong R_{p}\widetilde{S}_{r}^{+}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{+}e_{\varnothing,1}\wr_{s}\mathcal{T}_{d}^{R_{p}}).\end{split}$
For the group $\widetilde{S}_{n}^{-}$, $\kappa_{i}=\pm\sqrt{-1}$. Then $\kappa_{i}\in R_{p}$ if and only if $p\equiv 1~{\rm mod}~4$. Hence if $p\equiv 1~{\rm mod}~4$, the isomorphisms (5.8) of $R$-superalgebras restricts to isomorphisms of $R_{p}$-superalgebras
$\begin{split}R_{p}Lf&\cong R_{p}\widetilde{S}_{r}^{-}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{-}e_{\varnothing,1})^{\otimes d}\\ R_{p}N^{\rho,d}f&\cong R_{p}\widetilde{S}_{r}^{-}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{-}e_{\varnothing,1}\wr_{s}\mathcal{T}_{d}^{R_{p}}).\end{split}$
Lemma 5.15 ([22, Lemmas 4.9 and 4.10]).
Let $0<d<p$. As in 5.12, for $i=1,\cdots,d$, let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$ such that $T_{w}D_{i}T_{w}^{-1}=D_{w(i)}$, for all $i=1,\cdots,d$ and $w\in S_{d}$, and let
$D:=D_{1}\times\cdots\times D_{d}\leq\widetilde{S}_{V}.$
  1. 1.
    Each of $RGb$, RLfRLf and $RN^{\rho,d}f$ is a block with defect group $D$.
  2. 2.
    $RN^{\rho,d}f$ (resp. $RN_{\bar{0}}^{\rho,d}f$) is the Brauer correspondent of $RGb$ (resp. $RG_{\bar{0}}b$) in $N^{\rho,d}$ (resp. $N_{\bar{0}}^{\rho,d}$).
5.16[22, Definition 5.1].
Rouquier cores and RoCK blocks. Let $d\in\mathbb{N}$. A $d$-Rouquier $\bar{p}$}-core $\rho$ is a $\bar{p}$-core such that ${\rm Ab}_{\rho}$ has the following properties:
  1. 1.
    The first runner has at least $d$ beads.
  2. 2.
    The $(i+1)$-th runner has at least d1d-1 more beads than the $i$-th runner, for $1\leq i\leq(p-3)/2$.
If $\rho$ is a $d$-Rouquier $\bar{p}$-core, we refer to $B^{\rho,d}_{R}$ as an RoCK block of weight $d$ over $R$.
For more information on RoCK blocks we refer to [22, §5].
6 On Brauer correspondents, and proofs of Theorems 1.3 and 1.5
As in Section 5, we continue to assume that $K$ (and hence $\mathcal{O}$) contains a primitive $(2n!)$-th root of unity in this section. We also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$. Let $R\in\{\mathcal{O},k\}$ and $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$. In this section we continue to adopt the notation in 5.13; we always assume that $d>0$. Let $D$ be a Sylow $p$-subgroup of $\widetilde{S}_{V}$. Then by Theorem 5.6, $RN_{\widetilde{S}_{n}}(D)e_{\rho,0}$ is the Brauer correspondent of $RGb=R\widetilde{S}_{n}e_{\rho,d}$, and by [21, Theorem 5.2.10], $RN_{\widetilde{A}_{n}}(D)e_{\rho,0}$ is the Brauer correspondent of $RG_{\bar{0}}b=R\widetilde{A}_{n}e_{\rho,d}$. Note that we interpret $e_{\varnothing,0}$ as $e_{z}$. We set
$\begin{split}b^{\rho,d}_{R}&:=RN_{\widetilde{S}_{n}}(D)e_{\rho,0},~~(b^{\rho,d}_{R})_{\bar{0}}:=RN_{\widetilde{A}_{n}}(D)e_{\rho,0},\\ b^{\rho,d}_{R_{p}}&:=R_{p}N_{\widetilde{S}_{n}}(D)e_{\rho,0},~~(b^{\rho,d}_{R_{p}})_{\bar{0}}:=R_{p}N_{\widetilde{A}_{n}}(D)e_{\rho,0}.\end{split}$
Theorem 6.1 ([21, Proposition 5.2.13]).
Suppose that $0<d<p$. Let $P$ be a Sylow $p$-subgroup of $\widetilde{S}_{p}$, let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$ chosen as in 5.12, and let $D=D_{1}\times\cdots\times D_{d}$. Then we have an isomorphism of $k$-superalgebras
$b^{\rho,d}_{k}\cong k\widetilde{S}_{U}e_{\rho,0}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})\cong B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}).$ (6.1)
Theorem 6.2.
Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$. Suppose that $0<d<p$. Let $P$ be a Sylow $p$-subgroup of $\widetilde{S}_{p}$, let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$ chosen as in 5.12, and let $D=D_{1}\times\cdots\times D_{d}$. Then we have an isomorphism of $\mathbb{F}_{p}$-superalgebras
$b^{\rho,d}_{\mathbb{F}_{p}}\cong\mathbb{F}_{p}\widetilde{S}_{U}e_{\rho,0}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}).$ (6.2)
[…]
6.3.
Keep the notation of Theorem 6.2. We have the following isomorphisms all of which are obtained by restricting either the isomorphism (6.2) or the isomorphisms in Remark 5.14 to appropriate subalgebras:
$\begin{split}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}&\cong b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}\\ \mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}&\cong B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}\end{split}$ (6.3)
We realise [21, Lemma 5.2.19] over $\mathbb{F}_{p}$:
Lemma 6.4.
Let $P$ be a Sylow $p$-subgroup of $\widetilde{A}_{p}$. There is a splendid Rickard complex of $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodules that extends to a complex of $(\mathbb{F}_{p}\widetilde{S}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{S}_{p}}(P)e_{\varnothing,0})_{C_{2}}$-modules.
Proof. In [19, Theorem 1.10], Kessar and Linckelmann construct a splendid Rickard complex for blocks with a cyclic defect group over a non-splitting field, hence a splendid Rickard complex $X$ of $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodules. We review the construction of $X$ from the proof of [19, Theorem 1.10]. Note that $|P|=p$ and $O_{p}(\widetilde{A}_{p})$, the largest $p$-normal subgroup of $\widetilde{A}_{n}$, is the trivial subgroup. The complex $X$ is of the form
$X:=\cdots\to 0\to N\xrightarrow{\varphi}M\to 0\to\cdots$ (6.4)
where $M$ is in degree zero. Moreover $M$ is the unique (up to isomorphism) summand of the $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodule $e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}$. Let $Z\to M$ be a projective cover of the $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodule $M$. Then $N$ is a direct summand of $Z$ and $\varphi$ is the restriction of the map $Z\to M$ to $N$; moreover, if we write $Z=N\oplus N^{\prime}$, then $N$ and $N^{\prime}$ do not have nonzero isomorphic direct summand.
As in the proof of [21, Lemma 5.2.19], we transport our setup to the group algebra setting. Set
$\mathcal{A}=\langle\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P),(s,s)\rangle\leq\widetilde{S}_{p}\times N_{\widetilde{S}_{p}}(P),$
for some fixed $s\in N_{\widetilde{S}_{p}}(P)\backslash N_{\widetilde{S}_{p}}(P)$. We have an isomorphism of $\mathbb{F}_{p}$-superalgebras
$\mathbb{F}_{p}\mathcal{A}(e_{\varnothing,1}\otimes e_{\varnothing,0})$ $\cong(\mathbb{F}_{p}\widetilde{S}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{S}_{p}}(P)e_{\varnothing,0})_{C_{2}},$
$(a_{1},a_{2})(e_{\varnothing,1}\otimes e_{\varnothing,0})$ $\mapsto a_{1}e_{\varnothing,1}\otimes a_{2}^{-1}e_{\varnothing,0},$
$(s,s)(e_{\varnothing,1}\otimes e_{\varnothing,0})$ $\mapsto se_{\varnothing,1}\otimes s^{-1}e_{\varnothing,0},$
for all $a_{1}\in\widetilde{A}_{p}$ and $a_{2}\in N_{\widetilde{A}_{p}}(P)$. We identify $\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1}\otimes_{\mathbb{F}_{p}}(\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})^{\rm sop}$ with the subalgebra $\mathbb{F}_{p}(\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P))(e_{\varnothing,1}\otimes e_{\varnothing,0})$ of $\mathbb{F}_{p}\mathcal{A}(e_{\varnothing,1}\otimes e_{\varnothing,0})$.
[…]
$\tilde{X}:=\cdots\to 0\to\tilde{N}\xrightarrow{\tilde{\varphi}}\tilde{M}\to 0\to\cdots$
[…]
In [21, Proposition 5.2.21], Kleshchev and Livesey proved the following result for Brauer correspondents of blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ with abelian defect groups.
Theorem 6.5 ([21, Proposition 5.2.33]).
$kN_{\widetilde{S}_{n}}(D)e_{\rho,0}$, i.e. $b^{\rho,d}_{k}$, is splendid Rickard equivalent to $kNf$, and $kN_{\widetilde{A}_{n}}(D)e_{\rho,0}$, i.e. $(b^{\rho,d}_{k})_{\bar{0}}$, is splendid Rickard equivalent to $kN_{\bar{0}}f$.
Now we realise Theorem 6.5 over $\mathbb{F}_{p}$:
Theorem 6.6.
Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$. Then $\mathbb{F}_{p}N_{G}(D)e_{\rho,0}$, i.e. $b^{\rho,d}_{\mathbb{F}_{p}}$, is splendid Rickard equivalent to $\mathbb{F}_{p}N^{\rho,d}f$, and $\mathbb{F}_{p}N_{G}(D)e_{\rho,0}$, i.e. $(b^{\rho,d}_{\mathbb{F}_{p}})_{\bar{0}}$, is splendid Rickard equivalent to $\mathbb{F}_{p}N_{\bar{0}}^{\rho,d}f$.
Proof. By Remark 5.14, we have
$\mathbb{F}_{p}N^{\rho,d}f\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}).$
Let $P$ be a Sylow $p$-subgroup of $\widetilde{S}_{p}$. The block algebra $\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1}$ has a cyclic defect group $P$ and Brauer correspondent $\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0}$. By Lemma 6.4, there is a splendid Rickard complex $X$ of $((B^{\varnothing,1}_{\mathbb{F}_{p}})_{\bar{0}},(b^{\varnothing,1}_{\mathbb{F}_{p}})_{\bar{0}})$-bimodule that extends to a complex $\tilde{X}$ of $(B^{\varnothing,1}_{\mathbb{F}_{p}},b^{\varnothing,1}_{\mathbb{F}_{p}})_{C_{2}}$-modules. Let $H$ be the subgroup of $G$ generated by $\widetilde{S}_{U}$ and all $\widetilde{A}_{V_{i}}$’s. Now (by Propositions 4.2, 4.12 and Remark 2.3) $X^{\otimes d}$ induces a splendid Rickard equivalence between
$\mathbb{F}_{p}(H\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}\cong\mathbb{F}_{p}\widetilde{A}_{V_{1}}e_{\varnothing,1}\otimes_{\mathbb{F}_{p}}\cdots\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}\widetilde{A}_{V_{d}}e_{\varnothing,1}\cong(B^{\varnothing,1}_{\bar{0}})^{\otimes d}$
and
$\mathbb{F}_{p}(N_{G}(D)\cap H\cap\widetilde{S}_{V})e_{\varnothing,0}\cong\mathbb{F}_{p}N_{\widetilde{A}_{V_{1}}}(D_{1})e_{\varnothing,0}\otimes_{\mathbb{F}_{p}}\cdots\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}N_{\widetilde{A}_{V_{d}}}(D_{d})e_{\varnothing,0}\cong(b^{\varnothing,1}_{\bar{0}})^{\otimes d}.$
We now adopt the identifications (6.3). By Proposition 2.12, $\tilde{X}^{\otimes d}$ extends to a complex $Y_{0}$ of modules for the group algebra of
$\langle(H\cap\widetilde{S}_{V})\times(N_{G}(D)\cap H\cap\widetilde{S}_{V}),(u_{i},u_{i}),(T_{w},T_{w})\rangle\leq(N^{\rho,d}\cap\widetilde{S}_{V})\times(N_{G}(D)\cap\widetilde{S}_{V}),$
where $i$ ranges over $\{1,\cdots,d\}$ with $u_{i}\in N_{\widetilde{S}_{V_{i}}}(D_{i})\backslash N_{\widetilde{A}_{V_{i}}}(D_{i})$, and $w$ ranges over $S_{d}$ with $T_{w}$ defined as in 5.12. By Proposition 4.7 (iii), $Y_{0}$ is a splendid complex. We then induce $Y_{0}$ up to
$(N^{\rho,d}\cap\widetilde{S}_{V})\times(N_{G}(D)\cap\widetilde{S}_{V})$
to obtain a complex $Y$. Since $d<p$, we have $p\nmid|C_{2}\wr S_{d}|$. It follows from Theorem 4.5 that $Y$ induces a $(C_{2}\wr S_{d})$-graded Rickard equivalence between
$\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}\cong B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$
and
$\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}\cong b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}.$
By Proposition 4.7 (i), $Y$ is a splendid complex.
Next consider the group homomorphism
$C_{2}\wr S_{d}\to\mathbb{Z}/2\mathbb{Z},~~g_{i}\mapsto\bar{1},~~s_{j}\mapsto\bar{1}$
for $1\leq i\leq d$ and $1\leq j\leq d-1$, where $g_{i}$ is the generator of the $i$-th $C_{2}$ and $s_{j}$ is the appropriate elementary transposition in $S_{d}$. Let $\mathcal{K}$ be the kernel of this homomorphism. Since $B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$ is a $C_{2}\wr S_{d}$-graded algebra, we use the notation $(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})_{g}$ to denote the $g$-component of $B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$ for any $g\in C_{2}\wr S_{d}$. Then we have
$(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})_{\bar{0}}=\bigoplus_{g\in\mathcal{K}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})_{g}.$
Since $Y$ is a $C_{2}\wr S_{d}$-graded bimodule, $Y$ can be regarded as a bisupermodule ($\mathbb{Z}/2\mathbb{Z}$-graded bimodule) via
$Y_{\bar{0}}=\bigoplus_{g\in\mathcal{K}}Y_{g}~~{\rm and}~~Y_{\bar{1}}=\bigoplus_{g\in(C_{2}\wr S_{d})\backslash\mathcal{K}}Y_{g}.$
By the definition of a $(C_{2}\wr S_{d})$-graded Rickard equivalence, we see that $Y$ (considered as a bisupermodule) induces a $\mathbb{Z}/2\mathbb{Z}$-graded Rickard equivalence between
$\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}\cong B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$
and
$\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}\cong b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}.$
It is straightforward to check that the $(B^{\rho,0}_{\mathbb{F}_{p}},B^{\rho,0}_{\mathbb{F}_{p}})$-bisupermodule $B^{\rho,0}_{\mathbb{F}_{p}}$ induces a Morita self-superequivalence of $B^{\rho,0}_{\mathbb{F}_{p}}$; this bimodule is splendid because it is isomorphic to a direct summand of $\mathbb{F}_{p}\widetilde{S}_{r}\cong{\rm Ind}_{\Delta\widetilde{S}_{r}}^{\widetilde{S}_{r}\times\widetilde{S}_{r}}(\mathbb{F}_{p})$. Hence by Proposition 4.2, the complex $B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y$ of $\mathbb{Z}/2\mathbb{Z}$-graded $(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\rho,1}^{\otimes d})$-$(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\rho,0}^{\otimes d})$-bimodule induces a $\mathbb{Z}/2\mathbb{Z}$-graded Rickard equivalence between
$\mathbb{F}_{p}N^{\rho,d}f\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}$
and
$\mathbb{F}_{p}N_{G}(D)e_{\rho,d}\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}.$
By Proposition 4.12, $B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y$ is a splendid complex of $(\mathbb{F}_{p}N^{\rho,d}f,\mathbb{F}_{p}N_{G}(D)e_{\rho,d})$-bimodules.
Now by Theorem 4.4, $(B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y)_{\bar{0}}$ induces a Rickard equivalence between $\mathbb{F}_{p}N^{\rho,d}_{\bar{0}}f$ and $\mathbb{F}_{p}N_{G}(D)_{\bar{0}}e_{\rho,d}$. Since $(B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y)_{\bar{0}}$ is isomorphic to a direct summand of the complex
${\rm Res}_{N^{\rho,d}_{\bar{0}}\times N_{G}(D)_{\bar{0}}}^{N^{\rho,d}\times N_{G}(D)}(B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y),$
[…]
7 The Morita superequivalence bimodule, and proofs of Theorems 1.8 and 1.9
7.1.
Throughout this section we set $d$ to be an integer with $1\leq d<p$ and $\rho$ a $d$-Rouquier $\bar{p}$-core. As in Section 5, we continue to assume that $K$ (and hence $\mathcal{O}$) contains a primitive $(2n!)$-th root of unity in this section. We also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$. Let $R\in\{\mathcal{O},k\}$ and $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$. We will adopt the notation in 5.13, in particular, $r=|\rho|$, $n=r+dp$,
$G=\widetilde{S}_{n},\quad L=\widetilde{S}_{U,V_{1},\dots,V_{d}},\quad N^{\rho,d}=N_{G}(\widetilde{S}_{V_{1},\cdots,V_{d}}),\quad$
$b=e_{\rho,d}\in RG,\quad f=e_{\rho,0}\otimes e_{\varnothing,1}^{(1)}\otimes\dots\otimes e_{\varnothing,1}^{(d)}\in RLe_{z},$
and the defect group
$D=D_{1}\times\dots\times D_{d}$
is chosen as in Lemma 5.15. We identify $RLf$ with $B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\otimes d}$ and $RN^{\rho,d}f$ with $B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})$ via the isomorphisms (5.8). Let $\ell:=(p-1)/2$.
7.2.
The bisupermodule which induces a Morita superequivalence. We set $M^{R}$ to be the $(B^{\varnothing,1}_{R},B^{\varnothing,1}_{R})$-bisupermodule
$M^{R}:=\Omega_{B^{\varnothing,1}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\rm sop}}^{\ell}(B^{\varnothing,1}_{R}).$
Then we have the $((B^{\varnothing,1}_{R})^{\otimes d},(B^{\varnothing,1}_{R})^{\otimes d})$-bisupermodule $(M^{R})^{\boxtimes d}$. For $1\leq i\leq d$, we identify $(B^{\varnothing,1}_{R})^{(i)}:=R\widetilde{S}_{V_{i}}e_{\varnothing,1}^{(i)}$ with $B^{\varnothing,1}_{R}=R\widetilde{S}_{p}e_{\varnothing,1}$ via (5.5), as in 5.12. Denote by $(M^{R})^{(i)}$ the $i$-th factor in $(M^{R})^{\boxtimes d}$. That is, $(M^{R})^{(i)}$ is a $((B^{\varnothing,1}_{R})^{(i)},(B^{\varnothing,1}_{R})^{(i)})$-bisupermodule that we identify with $M^{R}$. Denote by $(M^{R})_{S_{d}}^{\boxtimes d}$ the $(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R},B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})_{S_{d}}$-supermodule defined as in (2.5).
Recalling the identification $RLf\cong B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\otimes d}$, define the $(RLf,RLf)$-bisupermodule
$M_{L}^{R}:=B^{\rho,0}_{R}\boxtimes_{R}(M^{R})^{\boxtimes d}.$
We have also identified $RN^{\rho,d}f$ with $B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})$. Recalling (2.6), define the $(RN^{\rho,d}f,RN^{\rho,d}f)$-bisupermodule
$M_{N^{\rho,d}}^{R}:=B^{\rho,0}_{R}\boxtimes_{R}(M^{R}\wr_{s}\mathcal{T}_{d}^{R})=B^{\rho,0}_{R}\boxtimes_{R}\Big({\rm Ind}_{(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R},B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})_{S_{d}}}^{(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})^{\rm sop}}(M^{R})^{\boxtimes d}_{S_{d}}\Big).$
By [22, Lemma 7.2 (ii)], the $(RN^{\rho,d}f,RN^{\rho,d}f)$-bimodule $M_{N^{\rho,d}}^{R}$ is module-indecomposable with vertex $\Delta D:=\{(u,u)\mid u\in D\}$, and by [22, Lemma 4.8 (ii)], we have $N_{G\times N^{\rho,d}}(\Delta D)\leq N^{\rho,d}\times N^{\rho,d}$. Hence we can now define $X^{R}$ to be the super Green correspondent of $M_{N^{\rho,d}}^{R}$ in $G\times N^{\rho,d}$; see Theorem 3.11 for the definition of the super Green correspondence. By [22, Lemma 7.5], $X^{R}$ is an $(RGb,RN^{\rho,d}f)$-bisupermodule.
Theorem 7.3 ([22, Theorem 8.13]).
The bisupermodules $X^{R}$ and $(X^{R})^{*}$ induce a Morita superequivalence between $RGb$ and $RN^{\rho,d}f$. $X^{R}_{\bar{0}}$ and $(X^{R}_{\bar{0}})^{*}$ induce a Morita superequivalence between $RG_{\bar{0}}b$ and $RN_{\bar{0}}^{\rho,d}f$. Both the bimodules $X^{R}$ and $X^{R}_{\bar{0}}$ have endopermutation sources.
Theorem 7.4.
Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$. Then $R_{p}Gf$ and $R_{p}Nf$ are Morita superequivalent. More precisely, there is an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$, together with its dual $(X^{R_{p}})^{*}$, inducing a Morita superequivalence between $R_{p}Gb$ and $R_{p}N^{\rho,d}f$ and fulfilling $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$. The bimodule $X^{\mathbb{F}_{p}}$ has endopermutation source.
Proof. By Propositions 3.5 and 3.4, it suffices to show that there exists an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$ such that $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$. By Remark (5.14) we have an isomorphism of $R_{p}$-superalgebras
$R_{p}N^{\rho,d}f\cong B^{\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}}).$
We let $M^{R_{p}}$ be the $(B^{\varnothing,1}_{R_{p}},B^{\varnothing,1}_{R_{p}})$-bimodule
$M^{R_{p}}:=\Omega_{B^{\varnothing,1}_{R_{p}}\otimes_{R_{p}}(B^{\varnothing,1}_{R_{p}})^{\rm sop}}^{\ell}(B^{\varnothing,1}_{R_{p}}).$
Then by Proposition 3.8 we have $M^{R}\cong R\otimes_{R_{p}}M^{R_{p}}$. Define the $(R_{p}N^{\rho,d}f,R_{p}N^{\rho,d}f)$-bisupermodule
$\begin{split}M_{N^{\rho,d}}^{R_{p}}&:=B^{\rho,0}_{R_{p}}\boxtimes_{R_{p}}(M^{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})\\ &=B^{\rho,0}_{R_{p}}\boxtimes_{R_{p}}\Big({\rm Ind}_{(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}},B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})_{S_{d}}}^{(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})\otimes_{R_{p}}(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})^{\rm sop}}(M^{R_{p}})^{\boxtimes d}_{S_{d}}\Big).\end{split}$
Then we have $M_{N^{\rho,d}}^{R}\cong R\otimes_{R_{p}}M_{N^{\rho,d}}^{R_{p}}$ as $(RN^{\rho,d}f,RN^{\rho,d}f)$-bisupermodules. Now by Proposition 3.9, there is an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$ such that $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$, as desired.
[…]
By [22, Lemma 3.10], Theorem 7.4 has the following corollary:
Corollary 7.5.
With the notation of Theorem 7.4, the $(R_{p}G_{\bar{0}}b,R_{p}N_{\bar{0}}^{\rho,d}f)$-bimodule $X_{\bar{0}}^{R_{p}}$ and its dual induce a Morita equivalence between $R_{p}G_{\bar{0}}b$ and $R_{p}N_{\bar{0}}^{\rho,d}f$.
Corollary 7.5 answers a question of Kleshchev and Livesey proposed in [22, Remark 8.15] in the highest generality of their question.
Theorem 7.6.
With the notation of Theorem 7.4, $\mathbb{Z}_{p}Gb$ is splendidly Rickard equivalent to $\mathbb{Z}_{p}N^{\rho,d}f$.
Proof. Since there is a finite extension of $\mathbb{Q}_{p}$ containing a primitive $(2n!)$-th root of unity and containing $\sqrt{-1}$, we may assume that $k$ is finite. By [22, Lemma 9.1], the $(\mathcal{O}Gb,\mathcal{O}N^{\rho,d}f)$-bimodule $X^{\mathcal{O}}$ is isomorphic to a direct summand of
$\mathcal{O}Gi\otimes_{\mathcal{O}D}{\rm Ind}_{\Delta D}^{D\times D}(U)\otimes j\mathcal{O}N^{\rho,d}$
for some source idempotents $i\in(\mathcal{O}Gb)^{D}$ and $j\in(\mathcal{O}N^{\rho,d}f)^{D}$, where $U$ is an endopermutation $\mathcal{O}\Delta D$-module. We can regard $U$ as an endopermutation $\mathcal{O}D$-module via the isomorphism $D\cong\Delta D$. By [22, Lemma 9.1], $U$ has an $N_{G}(D)$-stable endosplit $p$-permutation resolution $\mathcal{Y}$; see [30, Section 7] or [24, Definition 7.11.1] for this notion. It follows that $\mathcal{Y}$ is $\mathcal{F}$-stable, where $\mathcal{F}$ is the fusion system on $D$ determined by the source idempotent $i$ of $\mathcal{O}Gb$. By [24, Theorem 9.11.5 (i)], $X^{\mathcal{O}}$ has an endosplit $p$-permutation resolution $Y^{\mathcal{O}}$, and $Y^{\mathcal{O}}$ induces a splendid Rickard equivalence between $\mathcal{O}Gb$ and $\mathcal{O}N^{\rho,d}f$. Let $X^{k}=k\otimes_{\mathcal{O}}X^{\mathcal{O}}$, $Y^{k}=k\otimes_{\mathcal{O}}Y^{\mathcal{O}}$ and $X^{\mathbb{F}_{p}}=\mathbb{F}_{p}\otimes_{\mathbb{Z}_{p}}X^{\mathbb{Z}_{p}}$. By definition, one checks that $Y^{k}$ is an endosplit $p$-permutation resolution of $X^{k}$. By Theorem 7.4, we have $X^{k}\cong k\otimes_{\mathbb{F}_{p}}X^{\mathbb{F}_{p}}$. Then by [28, Theorem 1.3 (b)], there is a splendid complex of $(\mathbb{F}_{p}G\bar{b},\mathbb{F}_{p}N^{\rho,d}\bar{f})$-bimodules $Y^{\mathbb{F}_{p}}$ such that $Y^{k}\cong k\otimes_{\mathbb{F}_{p}}Y^{\mathbb{F}_{p}}$. By [19, Proposition 4.5 (a)], $Y^{\mathbb{F}_{p}}$ induce a splendid Rickard equivalence between $\mathbb{F}_{p}G\bar{b}$ and $\mathbb{F}_{p}N^{\rho,d}\bar{f}$. By lifting property of splendid Rickard equivalences ([30, Theorem 5.2]), $Y^{\mathbb{F}_{p}}$ lifts to a splendid Rickard complex $Y^{\mathbb{Z}_{p}}$ of $(\mathbb{Z}_{p}Gb,\mathbb{Z}_{p}N^{\rho,d}f)$-bimodules.
[…]
Theorem 7.7.
Let $G=\widetilde{S}_{n}^{+}$ (and hence $G_{\bar{0}}:=\widetilde{A}_{n}$). Then $\mathbb{Z}_{p}G_{\bar{0}}b$ is splendidly Rickard equivalent to $\mathbb{Z}_{p}N^{\rho,d}_{\bar{0}}f$.
Proof. Since $p\nmid|G\times N^{\rho,d}:G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}|$, any $p$-subgroup of $G\times N^{\rho,d}$ is contained in $G_{\bar{0}}\times N^{\rho,d}_{\bar{0}}$. Since ${\rm Res}_{G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}}^{G\times N^{\rho,d}}(X^{\mathcal{O}})=X_{\bar{0}}^{\mathcal{O}}\oplus X^{\mathcal{O}}_{\bar{1}}$, $X^{\mathcal{O}}_{\bar{0}}$ is isomorphic to a direct summand of
${\rm Res}_{G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}}^{G\times N^{\rho,d}}{\rm Ind}_{\Delta D}^{G\times N^{\rho,d}}(U)\cong\bigoplus_{t\in G\times N^{\rho,d}/G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}}{\rm Ind}_{{}^{t}(\Delta D)}^{G_{\bar{0}}\times N^{\rho,d}_{\bar{0}}}({}^{t}U),$
where the isomorphism is by using the Mackey formula. Again since $p\nmid|G\times N^{\rho,d}:G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}|$, any vertex of the bimodule $X_{\bar{0}}^{\mathcal{O}}$ has the same order as a vertex of the bimodule $X^{\mathcal{O}}$. It follows that $X_{\bar{0}}^{\mathcal{O}}$ has an endopermutation $)\mathcal{O}({}^{t}(\Delta D$))-source $^{t}U$ for some $t\in G\times N^{\rho,d}/G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}$. Write $t=(t_{1},t_{2})$ with $t_{1}\in G$ and $t_{2}\in N^{\rho,d}$; then $^{t}(\Delta D)=\{({}^{t_{1}}u,{}^{t_{2}}u)\mid u\in D\}$. Since $U$ is $N_{G}(D)$-stable, we see that $^{t_{1}}U$ is $N_{G_{\bar{0}}}({}^{t_{1}}D)$-stable when regarded as an $\mathcal{O}{}^{t_{1}}D$-module. Write $P:={}^{t_{1}}D$, $Q:={}^{t_{2}}D$ and $\varphi:P\cong Q$ the group isomorphism induced by conjugation action of $t_{2}t_{1}^{-1}$. By [24, Proposition 9.7.1], $P$ is a defect group of $\mathcal{O}G_{\bar{0}}b$ and $Q$ is a defect group of $\mathcal{O}N^{\rho,d}_{\bar{0}}f$. Note that $X_{\bar{0}}^{\mathcal{O}}$ is isomorphic to a direct summand of
$\mathcal{O}G_{\bar{0}}b\otimes_{\mathcal{O}P}{\rm Ind}_{\Delta\varphi}^{P\times Q}({}^{t_{1}}U)\otimes_{\mathcal{O}Q}\mathcal{O}N^{\rho,d}_{\bar{0}}f.$
Then using the same argument as in [24, page 320], $X_{\bar{0}}^{\mathcal{O}}$ is isomorphic to a direct summand of
$\mathcal{O}G_{\bar{0}}i_{0}\otimes_{\mathcal{O}P}{\rm Ind}_{\Delta\varphi}^{P\times Q}({}^{t_{1}}U)\otimes_{\mathcal{O}Q}j_{0}\mathcal{O}N^{\rho,d}_{\bar{0}}$
[…]
Remark 7.8.
By [22, Theorem 9.7], when $p\equiv 1~{\rm mod}~4$, the $\mathcal{O}D$-source $U$ of $X^{\mathcal{O}}$ is the trivial module. Hence $X^{\mathcal{O}}$ induces a splendid Morita equivalence between $\mathcal{O}Gb$ and $\mathcal{O}N^{\rho,d}d$, and $X^{\mathcal{O}}_{\bar{0}}$ induces a splendid Morita equivalence between $\mathcal{O}G_{\bar{0}}b$ and $\mathcal{O}N^{\rho,d}_{\bar{0}}f$. So by Theorem 7.4 we immediately have the following:
Theorem 7.9.
If $p\equiv 1~{\rm mod}~4$, then $X^{\mathbb{Z}_{p}}$ induces a splendid Morita equivalence between $\mathbb{Z}_{p}Gb$ and $\mathbb{Z}_{p}N^{\rho,d}d$, and $X^{\mathbb{Z}_{p}}_{\bar{0}}$ induces a splendid Morita equivalence between $\mathbb{Z}_{p}G_{\bar{0}}b$ and $\mathbb{Z}_{p}N^{\rho,d}_{\bar{0}}f$.
We can now construct splendid Rickard equivalences between $\mathbb{Z}_{p}Gb$ (resp. $\mathbb{Z}_{p}G_{\bar{0}}b$) and its Brauer correspondent in the normaliser of a defect group.
Theorem 7.10.
Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$. Let $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$. Then $R_{p}Gb$ and $R_{p}G_{\bar{0}}b$ are splendidly Rickard equivalent to their Brauer correspondent in $R_{p}N_{G}(D)$ and $R_{p}N_{G_{0}}(D)$, respectively.
Proof. By the lifting theorem of splendid Rickard equivalences [30, Theorem 5.2], it suffices to consider $R_{p}=\mathbb{F}_{p}$. By Theorem 6.6, there is a splendid Rickard equivalence between the Brauer correspondent of $\mathbb{F}_{p}Gb$ (resp. $\mathbb{F}_{p}G_{\bar{0}}$) in $\mathbb{F}_{p}N_{G}(D)$ (resp. $\mathbb{F}_{p}N_{G_{0}}(D)$) and $\mathbb{F}_{p}N^{\rho,d}f$ (resp. $\mathbb{F}_{p}N^{\rho,d}_{\bar{0}}f$). Now the statements follows from Theorems 7.6, 7.7 and Proposition 4.8.
[…]
Acknowledgements. We thank Prof. Pham Huu Tiep for suggesting us to consider spin blocks of symmetric and alternating groups in SUSTech, May 2024, and thank Prof. Alexander Kleshchev for helpful explanation.
The first author is supported by National Key R & D Profram of China (No. 2020YFE020 4200) and NSFC (Nos. 12431001, 12350710787). The second author is supported by Fundamental Research Funds for the Central Universities (CCNU24XJ028/25JC025/25JCPT031), and NSFC (Nos. 12501024, 12471016).

### v5

@@
 Recently, Kleshchev and Livesey [22, 23] proved the existence of RoCK $p$-blocks for double covers of symmetric and alternating groups over large enough coefficient rings.
-They proved that RoCK blocks of double covers are Morita equivalent to standard “local” blocks via bimodules with endopermutation source.
-Based on this, Kleshchev and Livesey proved that these RoCK blocks are splendidly Rickard equivalent to their Brauer correspondents.
+They proved that these RoCK blocks of double covers are Morita equivalent to standard “local” blocks via bimodules with endopermutation source.
+Based on this, Kleshchev and Livesey proved that RoCK blocks are splendidly Rickard equivalent to their Brauer correspondents.
 The analogous result for blocks of symmetric groups, a theorem of Chuang and Kessar, was an important step in Chuang and Rouquier ultimately proving Broué’s abelian defect group conjecture for symmetric groups.
-In this paper we show that in most cases the Morita equivalences and splendid Rickard equivalences constructed by Kleshchev and Livesey descend to the ring $\mathbb{Z}_{p}$ of $p$-adic integers, hence prove Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture for most of these RoCK blocks.
+In this paper we show that the Morita and splendid Rickard equivalences constructed by Kleshchev and Livesey descend to the ring $\mathbb{Z}_{p}$ of $p$-adic integers, hence prove Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture for these RoCK blocks.
 keywords:
@@
 For an arbitrary complete discrete valuation ring $\mathcal{O}$ and a block idempotent $b$ of a finite group $G$ over $\mathcal{O}$ with an abelian defect group, there is a splendid Rickard equivalence between $\mathcal{O}Gb$ and its Brauer correspondent.
-Comparing to the refined version, the usual Broué conjecture is with the assumption that the complete discrete valuation rings is large enough.
-Conjecture 1.1 has been proved for blocks with cyclic defect groups in [19], blocks with a Klein four defect group in [12], blocks of alternating groups in [13], blocks of ${\rm SL}_{n}(q)$ and ${\rm GL}_{n}(q)$ in defining characteristic (see [15]), and unipotent blocks ${\rm GL}_{n}(q)$ in non-defining characteristic (see [16]).
+Comparing to the refined version, the usual Broué conjecture is with the assumption that the complete discrete valuation ring is large enough for $G$.
+Hence Conjecture 1.1 contains more local-global information and has more applications; see [20, Corollary 1.9], [1, Theorem 1.4] and [15, Theorem 1].
+Conjecture 1.1 has been proved for blocks with cyclic defect groups in [20], blocks with a Klein four defect group in [13], blocks of alternating groups in [14], blocks of ${\rm SL}_{n}(q)$ and ${\rm GL}_{n}(q)$ in defining characteristic (see [16]), and unipotent blocks of ${\rm GL}_{n}(q)$ in non-defining characteristics (see [17]).
 Conjecture 1.1 has also been verified for certain blocks of $p$-nilpotent and $p$-solvable groups in [3], [27] and [28].
-Since any field is a splitting field for symmetric groups, Conjecture 1.1 has been proved for blocks of symmetric groups by Chuang and Rouquier [6, Theorem 7.6].
-For a non-negative integer $n$, denote by $S_{n}$ the symmetric group on $n$ letters and by $\widetilde{S}_{n}$ one of the two double covers $\widetilde{S}_{n}^{+}$ and $\widetilde{S}_{n}^{-}$ of $S_{n}$.
+Since any field is a splitting field for symmetric groups, Conjecture 1.1 - which in the symmetric group case is equivalent to the usual Broué conjecture - has been proved by Chuang and Rouquier [7, Theorem 7.6].
+For a non-negative integer $n$, denote by $S_{n}$ the symmetric group on $n$ letters and by $\widetilde{S}^{\eta}_{n}$ ($\eta\in\{\pm\}$) one of the two double covers $\widetilde{S}_{n}^{+}$ and $\widetilde{S}_{n}^{-}$ of $S_{n}$; see 2.13.
 Denote by $A_{n}$ the corresponding alternating group and by $\widetilde{A}_{n}$ its double cover.
-In this paper we provide additional evidence for Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture by considering blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$.
+In this paper we provide additional evidence for Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture by considering blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$.
 Let $\mathcal{O}$ be a complete discrete valuation ring.
-Denote by $z$ the central element in $\widetilde{S}_{n}$ of order $2$.
-Since $\mathcal{O}\widetilde{S}_{n}(1+z)/2$ is isomorphic to $\mathcal{O}S_{n}$ and $\mathcal{O}\widetilde{A}_{n}(1+z)/2$ is isomorphic to $\mathcal{O}A_{n}$, Conjecture 1.1 is true for blocks of $\mathcal{O}\widetilde{S}_{n}(1+z)/2$ and of $\mathcal{O}\widetilde{A}_{n}(1+z)/2$.
-So we restrict our attention in this paper to blocks of $\mathcal{O}\widetilde{S}_{n}(1-z)/2$ and of $\mathcal{O}\widetilde{A}_{n}(1-z)/2$, which are called the spin blocks of the symmetric and alternating groups, respectively.
+Denote by $z$ the central element in $\widetilde{S}^{\eta}_{n}$ of order $2$.
+Since $\mathcal{O}\widetilde{S}^{\eta}_{n}(1+z)/2$ is isomorphic to $\mathcal{O}S_{n}$ and $\mathcal{O}\widetilde{A}_{n}(1+z)/2$ is isomorphic to $\mathcal{O}A_{n}$, Conjecture 1.1 is true for blocks of $\mathcal{O}\widetilde{S}^{\eta}_{n}(1+z)/2$ and of $\mathcal{O}\widetilde{A}_{n}(1+z)/2$.
+So we restrict our attention in this paper to blocks of $\mathcal{O}\widetilde{S}^{\eta}_{n}(1-z)/2$ and of $\mathcal{O}\widetilde{A}_{n}(1-z)/2$, which are called the spin blocks of the symmetric and alternating groups, respectively.
 For block with trivial defect group, Conjecture 1.1 is trivially true, hence we only need to consider Conjecture 1.1 for block with nontrivial defect groups.
@@
 Let $p:={\rm char}(k)$, a prime number.
-For the rest of this introduction we assume that $K$ has characteristic 0, and when dealing with a finite group $G$, we will assume that $K$ (and consequently $\mathcal{O}$) contains a primitive $|G|$-th root of unity; we also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$).
+For the rest of this introduction we assume that $K$ has characteristic 0, and when dealing with a finite group $G$, we will assume that $K$ (and consequently $\mathcal{O}$) contains a primitive $|G|$-th root of unity; following [22, 23] we also assume that 1-1 and $2$ have square roots (say $\sqrt{-1}$ and $\sqrt{2}$) in $K$ (and hence $\mathcal{O}$).
 Let $\mathbb{Z}_{p}$ denote the ring of $p$-adic integers, $\mathbb{Q}_{p}$ the field of $p$-adic numbers and $\mathbb{F}_{p}$ the field with $p$ elements.
-Since $z$ is a central element of order $2$, if $p=2$, then $\widetilde{S}_{n}$ (resp.
-$\widetilde{A}_{n}$) satisfies Conjecture 1.1 if and only if $S_{n}\cong\widetilde{S}_{n}/\langle z\rangle$ (resp.
+Since $z$ is a central element of order $2$, if $p=2$, then $\widetilde{S}^{\eta}_{n}$ (resp.
+$\widetilde{A}_{n}$) satisfies Conjecture 1.1 if and only if $S_{n}\cong\widetilde{S}^{\eta}_{n}/\langle z\rangle$ (resp.
 $A_{n}\cong\widetilde{A}_{n}/\langle z\rangle$) satisfies Conjecture 1.1; see [31, Lemma 10.2.11] (we note that the blanket assumption in [31] that the coefficient ring is large enough is not used in the proof of [31, Lemma 10.2.11]).
-Hence Conjecture 1.1 holds for $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ if $p=2$.
+Hence Conjecture 1.1 holds for $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$ if $p=2$.
 So from now on we assume that $p\geq 3$.
-Let $R\in\{\mathcal{O},k\}$, $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$ and let $e_{z}:=(1-z)/2\in R\widetilde{S}_{n}$.
-From [21, 22] we see that it is important to consider $R\tilde{S}_{n}$ as a superalgebra via
-$(R\widetilde{S}_{n})_{\bar{0}}:=R\widetilde{A}_{n}~~~{\rm and}~~~(R\widetilde{S}_{n})_{\bar{1}}:=R(\widetilde{S}_{n}\backslash\widetilde{A}_{n}).$
+Let $(R,R_{p})\in\{(\mathcal{O},\mathbb{Z}_{p}),(k,\mathbb{F}_{p})\}$ and let $e_{z}:=(1-z)/2\in R\widetilde{S}^{\eta}_{n}$.
+From [22, 23] we see that it is important to consider $R\widetilde{S}^{\eta}_{n}$ as a superalgebra via
+$(R\widetilde{S}^{\eta}_{n})_{\bar{0}}:=R\widetilde{A}_{n}~~~{\rm and}~~~(R\widetilde{S}^{\eta}_{n})_{\bar{1}}:=R(\widetilde{S}^{\eta}_{n}\backslash\widetilde{A}_{n}).$
 Recall from [23] that the spin superblock algebras (i.e.
-the indecomposable superideals) of $R\tilde{S}_{n}e_{z}$ are labeled by pairs $(\rho,d)$, where $\rho$ is a $\bar{p}$-core ($p$-bar core) and $d\in\mathbb{N}$ (called the weight of the block) such that $n=|\rho|+d\rho$.
-Denote such a superblock algebra by $B^{\rho,d}_{R}$ and denote by $e_{\rho,d}$ the central idempotent of $R\widetilde{S}_{n}$ such that $B^{\rho,d}_{R}=R\widetilde{S}_{n}e_{\rho,d}$.
-We prove that the idempotent $e_{\rho,d}$ is contained in the subring $R_{p}\widetilde{S}_{n}$ (see Proposition 5.9 below).
-Hence we can define $B^{\rho,d}_{R_{p}}:=R_{p}\widetilde{S}_{n}e_{\rho,d}$.
+the indecomposable superideals) of $R\widetilde{S}^{\eta}_{n}e_{z}$ are labeled by pairs $(\rho,d)$, where $\rho$ is a $\bar{p}$-core ($p$-bar core) and $d\in\mathbb{N}$ (called the weight of the block) such that $n=|\rho|+dp$.
+Denote such a superblock algebra by $B^{\eta,\rho,d}_{R}$ and denote by $e_{\rho,d}$ the central idempotent of $R\widetilde{S}^{\eta}_{n}$ such that $B^{\eta,\rho,d}_{R}=R\widetilde{S}^{\eta}_{n}e_{\rho,d}$.
+We prove (in Proposition 6.8 below) that the idempotent $e_{\rho,d}$ is contained in the subring $R_{p}\widetilde{S}^{\eta}_{n}$.
+Hence we can define $B^{\eta,\rho,d}_{R_{p}}:=R_{p}\widetilde{S}^{\eta}_{n}e_{\rho,d}$.
 Note that the case $d=0$ corresponds to the defect zero situations, so we assume that $d>0$ in the rest of this section.
-In this case $B^{\rho,d}_{R}$ is in fact a spin block algebra (and not just a superblock algebra) of $R\widetilde{S}_{n}$.
-We denote by $D^{\rho,d}$ a defect group of $B^{\rho,d}_{R}$.
-The defect group $D^{\rho,d}$ of $B^{\rho,d}_{R}$ is abelian if and only if $d<p$.
+In this case $B^{\eta,\rho,d}_{R}$ is in fact a spin block algebra (and not just a superblock algebra) of $R\widetilde{S}^{\eta}_{n}$.
+We denote by $D^{\rho,d}$ a defect group of $B^{\eta,\rho,d}_{R}$.
+The defect group $D^{\rho,d}$ of $B^{\eta,\rho,d}_{R}$ is abelian if and only if $d<p$.
 Moreover, if $d<p$, then $D^{\rho,d}\cong(C_{p})^{\times d}$.
-By [22, Remark 4.4], the idempotent $e_{\rho,d}$ is contained in the subalgebra $R\widetilde{A}_{n}e_{z}$, and hence (by Proposition 5.9) in $R_{p}\widetilde{A}_{n}e_{z}$.
-By [22, Theorem 4.5] (or [18, Propsition 3.16]), as an element of $R\widetilde{A}_{n}$, $e_{\rho,d}$ is a single block unless $d=0$ and $\rho$ is even.
-So $B^{\rho,d}_{R}$ can be regarded as an $R$-superalgebra with even part $(B^{\rm\rho,d}_{R})_{\bar{0}}:=R\widetilde{A}_{n}e_{\rho,d}$.
-We denote $R_{p}\widetilde{A}_{n}e_{\rho,d}$ by $(B^{\rm\rho,d}_{R_{p}})_{\bar{0}}$.
-It is known that $D^{\rho,d}$ is also a defect group of $B^{\rho,d}_{\bar{0}}$ (see [18, Proposition 3.16] or Theorem 5.8 below).
-Let $b^{\rho,d}_{R}$ (resp.
-$(b^{\rho,d}_{R})_{\bar{0}}$) be the Brauer correspondent of $B^{\rho,d}_{R}$ (resp.
-$(B^{\rho,d}_{R})_{\bar{0}}$) in $RN_{\widetilde{S}_{n}}(D^{\rho,d})$ (resp.
+By [23, Remark 4.4], the idempotent $e_{\rho,d}$ is contained in the subalgebra $R\widetilde{A}_{n}e_{z}$, and hence (by Proposition 6.8) in $R_{p}\widetilde{A}_{n}e_{z}$.
+By [23, Theorem 4.5] (or [19, Propsition 3.16]), as an element of $R\widetilde{A}_{n}$, $e_{\rho,d}$ is a single block unless $d=0$ and $\rho$ is even.
+So $B^{\eta,\rho,d}_{R}$ can be regarded as an $R$-superalgebra with even part $(B^{\eta,\rho,d}_{R})_{\bar{0}}:=R\widetilde{A}_{n}e_{\rho,d}$.
+We denote $R_{p}\widetilde{A}_{n}e_{\rho,d}$ by $(B^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$.
+It is known that $D^{\rho,d}$ is also a defect group of $B^{\eta,\rho,d}_{\bar{0}}$ (see [19, Proposition 3.16] or Theorem 6.7 below).
+Let $B^{\eta,\rho,d}_{R}$ (resp.
+$(b^{\eta,\rho,d}_{R})_{\bar{0}}$) be the Brauer correspondent of $B^{\eta,\rho,d}_{R}$ (resp.
+$(B^{\eta,\rho,d}_{R})_{\bar{0}}$) in $RN_{\widetilde{S}^{\eta}_{n}}(D^{\rho,d})$ (resp.
 $RN_{\widetilde{A}_{n}}(D^{\rho,d})$).
-Similarly, we let $b^{\rho,d}_{R_{p}}$ (resp.
-$(b^{\rho,d}_{R_{p}})_{\bar{0}}$) be the Brauer correspondent of $B^{\rho,d}_{R_{p}}$ (resp.
-$(B^{\rho,d}_{R_{p}})_{\bar{0}}$) in $R_{p}N_{\widetilde{S}_{n}}(D^{\rho,d})$ (resp.
+Similarly, we let $b^{\eta,\rho,d}_{R_{p}}$ (resp.
+$(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$) be the Brauer correspondent of $B^{\eta,\rho,d}_{R_{p}}$ (resp.
+$(B^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$) in $R_{p}N_{\widetilde{S}^{\eta}_{n}}(D^{\rho,d})$ (resp.
 $R_{p}N_{\widetilde{A}_{n}}(D^{\rho,d})$).
-Note that $(b^{\rho,d}_{R})_{\bar{0}}$ (resp.
-$(b^{\rho,d}_{R_{p}})_{\bar{0}}$) is indeed the even part of the superalgebra $B^{\rho,d}_{R}$ (resp.
-$B^{\rho,d}_{R_{p}}$).
-Let $\mathcal{T}_{n}^{R}:=R\widetilde{S}_{n}e_{z}$.
-For a superalgebra $A$, a twisted wreath superproduct $A\wr_{s}\mathcal{T}_{d}^{R}$ is defined in [22, §3.7] (see also 2.7 below).
-In [21], Kleshchev and Livesey proved the following structure theorem of $b^{\rho,d}_{k}$:
-Theorem 1.2 ([21, Theorem A]).
+Note that $(b^{\eta,\rho,d}_{R})_{\bar{0}}$ (resp.
+$(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$) is indeed the even part of the superalgebra $b^{\eta,\rho,d}_{R}$ (resp.
+$b^{\eta,\rho,d}_{R_{p}}$).
+If $d>0$, we have $b^{\eta,\rho,d}_{R}=RN_{\widetilde{S}^{\eta}_{n}}(D^{\rho,d})e_{\rho,0}$, $(b^{\eta,\rho,d}_{R})_{\bar{0}}=RN_{\widetilde{A}_{n}}(D^{\rho,d})e_{\rho,0}$, $b^{\eta,\rho,d}_{R_{p}}=R_{p}N_{\widetilde{S}^{\eta}_{n}}(D^{\rho,d})e_{\rho,0}$ and $(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}=R_{p}N_{\widetilde{A}_{n}}(D^{\rho,d})e_{\rho,0}$; see Theorem 6.5 (ii) below.
+Let $\mathcal{T}_{n,R}$ be the twisted group superalgebra defined by Kleshchev and Livesey [22, 23]; see 2.7 below.
+For a superalgebra $A$, the twisted wreath superproduct $A\wr_{s}\mathcal{T}_{n,R}$ is defined in [23, §3.7] (see also 2.7 below).
+In order to consider the refined Broué conjecture, for $\eta\in\{\pm\}$, we define (in 3.1) the $\eta$-twisted group superalgebra $\mathcal{T}^{\eta}_{n,R}$ and the $\eta$-twisted wreath superproduct $A\wr_{s}\mathcal{T}^{\eta}_{n,R}$.
+We prove in Proposition 3.2 that $\mathcal{T}_{n,R}\cong\mathcal{T}^{\eta}_{n,R}\cong R\otimes_{R_{p}}\mathcal{T}^{\eta}_{n,R_{p}}$ and $A\wr_{s}\mathcal{T}_{n,R}\cong A\wr_{s}\mathcal{T}^{\eta}_{n,R}$ as $R$-superalgebras.
+Note that if $\sqrt{-1}\notin R_{p}$, then it seems that $\mathcal{T}^{+}_{n,R_{p}}$ and $\mathcal{T}^{-}_{n,R_{p}}$ are not isomorphic as $R$-algebras.
+In [22], Kleshchev and Livesey proved the following structure theorem of $b^{\eta,\rho,d}_{k}$:
+Theorem 1.2 ([22, Theorem A]).
 Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Then
   1. 1.
-    $b^{\rho,d}_{k}\cong B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})$ as $k$-superalgebras.
+    $b^{\eta,\rho,d}_{k}\cong B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k})$ as $k$-superalgebras.
   2. 2.
-    $(b^{\rho,d}_{k})_{\bar{0}}\cong(B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}))_{\bar{0}}$ as $k$-algebras.
-Based on the above theorem we prove a similar structure theorem of $b^{\rho,d}_{\mathbb{F}_{p}}$:
+    $(b^{\eta,\rho,d}_{k})_{\bar{0}}\cong(B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k}))_{\bar{0}}$ as $k$-algebras.
+Based on the above theorem we prove a similar structure theorem of $b^{\eta,\rho,d}_{\mathbb{F}_{p}}$:
 Theorem 1.3 (Theorem 7.2).
-Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$. Then
-  1. 1.
-    $b^{\rho,d}_{\mathbb{F}_{p}}\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})$ as $\mathbb{F}_{p}$-superalgebras.
-  2. 2.
-    $(b^{\rho,d}_{\mathbb{F}_{p}})_{\bar{0}}\cong(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}))_{\bar{0}}$ as $\mathbb{F}_{p}$-algebras.
-In [21], Kleshchev and Livesey proved the following analogue of “the Rickard and Marcus step for symmetric groups” for double covers of symmetric and alternating groups:
-Theorem 1.4 ([21, Theorem C]).
 Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Then
   1. 1.
-    $B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})$ is derived equivalent to $B^{\rho,0}_{k}\otimes_{k}(B^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})$.
+    $b^{\eta,\rho,d}_{\mathbb{F}_{p}}\cong B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})$ as $\mathbb{F}_{p}$-superalgebras.
   2. 2.
-    $(B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}))_{\bar{0}}$ is derived equivalent to $(B^{\rho,0}_{k}\otimes_{k}(B^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}))_{\bar{0}}.$
+    $(b^{\eta,\rho,d}_{\mathbb{F}_{p}})_{\bar{0}}\cong(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}))_{\bar{0}}$ as $\mathbb{F}_{p}$-algebras.
+In [22], Kleshchev and Livesey proved the following analogue of “the Rickard and Mǎrcuş step for symmetric groups” for double covers of symmetric and alternating groups:
+Theorem 1.4 ([22, Theorem C]).
+Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Then
+  1. 1.
+    $B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k})$ is derived equivalent to $B^{\eta,\rho,0}_{k}\otimes_{k}(B^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k})$.
+  2. 2.
+    $(B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k}))_{\bar{0}}$ is derived equivalent to $(B^{\eta,\rho,0}_{k}\otimes_{k}(B^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k}))_{\bar{0}}.$
 We prove a similar result over $\mathbb{F}_{p}$:
 Theorem 1.5 (Theorem 7.6).
-Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$. Then
+Let $0<d<p$ and $\rho$ be any $\bar{p}$-core. Then
   1. 1.
-    $B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})$ is splendidly Rickard equivalent to $B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})$.
+    $B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})$ is splendidly Rickard equivalent to $B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})$.
   2. 2.
-    $(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}))_{\bar{0}}$ is splendidly Rickard equivalent to $(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}))_{\bar{0}}$.
-See Definitions 4.9 and 4.11 below for splendid Rickard equivalences between general interior $P$-algebras.
-In [21], Kleshchev and Livesey defined the notion of a $d$-Rouquier $\bar{p}$-core for $d\in\mathbb{N}$ (see 5.16 below) and proved the following result:
-Theorem 1.6 ([21, Theorem D], [22, Theorem A]).
+    $(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}))_{\bar{0}}$ is splendidly Rickard equivalent to $(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}))_{\bar{0}}$.
+See Definitions 5.8 and 5.13 below for splendid Rickard equivalences between general interior $P$-algebras.
+In [22], Kleshchev and Livesey defined the notion of a $d$-Rouquier $\bar{p}$-core for $d\in\mathbb{N}$ (see 6.16 below) and proved the following result:
+Theorem 1.6 ([22, Theorem D], [23, Theorem A and Lemma 9.2]).
 Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core.
-Then $B^{\rho,d}_{R}$ is Morita superequivalent to $B^{\rho,0}_{R}\otimes_{R}(B^{\emptyset,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})$ via a bimodule with endopermutation source.
-For the definition of a Morita superequivalence, see [22, §3.3] or 2.4 below.
-Following [22] we call the block $e_{\rho,d}$ of $R\widetilde{S}_{n}$ in Theorem 1.6 an RoCK block of $\widetilde{S}_{n}$.
+Then $B^{\eta,\rho,d}_{R}$ is Morita superequivalent to $B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d,R})$ via a bimodule $X^{R}$ with an endopermutation source $U^{R}$.
+For the definition of a Morita superequivalence, see [23, §3.3] or 2.4 below.
+Following [23] we call the block $e_{\rho,d}$ of $R\widetilde{S}^{\eta}_{n}$ in Theorem 1.6 an RoCK block of $\widetilde{S}^{\eta}_{n}$.
 Since $e_{\rho,d}$ is also a block idempotent of $R\widetilde{A}_{n}$, following [23] we also called the the block $e_{\rho,d}$ of $R\widetilde{A}_{n}$ an RoCK block of $\widetilde{A}_{n}$.
-Using Theorem 1.6, Kleshchev and Livesey proved Broué’s abelian defect group conjecture for RoCK blocks of $R\widetilde{S}_{n}$ and $R\widetilde{A}_{n}$:
-Theorem 1.7 ([22, Theorem B]).
+Using Theorem 1.6, Kleshchev and Livesey proved Broué’s abelian defect group conjecture for RoCK blocks of $R\widetilde{S}^{\eta}_{n}$ and $R\widetilde{A}_{n}$:
+Theorem 1.7 ([23, Theorem B]).
 Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core.
-Then $B^{\rho,d}_{R}$ and $(B^{\rho,d}_{R})_{\bar{0}}$ are splendidly Rickard equivalent to $b^{\rho,d}_{R}$ and $(b^{\rho,d}_{R})_{\bar{0}}$ respectively.
-In this paper we prove that in most cases Theorems 1.6 and 1.7 descend to $R_{p}$:
+Then $B^{\eta,\rho,d}_{R}$ and $(B^{\eta,\rho,d}_{R})_{\bar{0}}$ are splendidly Rickard equivalent to $b^{\eta,\rho,d}_{R}$ and $(b^{\eta,\rho,d}_{R})_{\bar{0}}$ respectively.
+In this paper we prove that Theorems 1.6 and 1.7 descend to $R_{p}$:
 Theorem 1.8 (Theorem 8.4).
 Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core.
-Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$.
-Then $B^{\rho,d}_{R_{p}}$ is Morita superequivalent to $B^{\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\emptyset,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})$.
-Corollary 7.5 (a corollary of Theorem 1.8) answers a question of Kleshchev and Livesey proposed in [22, Remark 8.15] in the highest generality of their question.
+Then $B^{\eta,\rho,d}_{R_{p}}$ is Morita superequivalent to $B^{\eta,\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})$ via a bimodule $X^{R_{p}}$ with an endopermutation source $U^{R_{p}}$, such that $R\otimes_{R_{p}}X^{R_{p}}$ and $R\otimes_{R_{p}}U^{R_{p}}$ are isomorphic to the $X^{R}$ and $U^{R}$ in Theorem 1.6, respectively.
+Corollary 8.5 (a corollary of Theorem 1.8) answers a question of Kleshchev and Livesey proposed in [23, Remark 8.15] in the highest generality of their question.
 Theorem 1.9 (Theorem 8.10).
 Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core.
-Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$.
-Then $B^{\rho,d}_{R_{p}}$ and $(B^{\rho,d}_{R_{p}})_{\bar{0}}$ are splendidly Rickard equivalent to $b^{\rho,d}_{R_{p}}$ and $(b^{\rho,d}_{R_{p}})_{\bar{0}}$ respectively.
-In particular, Conjecture 1.1 holds for RoCK blocks of $\widetilde{S}_{n}^{+}$ and $\widetilde{A}_{n}$, and if $p\equiv 1~{\rm mod}~4$, Conjecture 1.1 also holds for RoCK blocks of $\widetilde{S}_{n}^{-}$.
+Then $B^{\eta,\rho,d}_{R_{p}}$ and $(B^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$ are splendidly Rickard equivalent to $b^{\eta,\rho,d}_{R_{p}}$ and $(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$ respectively.
+In particular, Conjecture 1.1 holds for RoCK blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$.
 Besides Theorems 1.3, 1.5 and 1.8, our proof of Theorem 1.9 makes use of a recent theorem of Miller [28, Theorem 1.3 (b)], which gives a descent criterion of an endosplit $p$-permutation resolution of a module.
-Remark 1.10.
-For the group $\widetilde{S}_{n}^{-}$, if $p\equiv 3~{\rm mod}~4$, Theorems 1.2, 1.4, 1.6 and 1.7 do not descend to $R_{p}$, but descend to the subring of $R$ generated by $R_{p}$ and a primitive $4$-th root of unity $\sqrt{-1}$.
-The reason is that by (5.2) below we have an isomorphism of superalgebras
-$R\widetilde{S}_{n}^{-}e_{z}\cong\mathcal{T}_{n}^{R},~t_{i}e_{z}\mapsto(-1)^{i}\sqrt{-1}\tau_{i},\ ze_{z}\mapsto-1,$
-but the element $\sqrt{-1}\in R$ is not in the subring $R_{p}$.
-Using [1, Theorem 1.4] and [14, Theorem 1], we obtain a corollary of Theorem 1.9:
+Using [1, Theorem 1.4] and [15, Theorem 1], we obtain a corollary of Theorem 1.9:
 Corollary 1.10.
-Refinements of the Alperin–McKay conjecture by Isaacs–Navarro, Navarro, Turull and Boltje (see [1, (1.5), (1.6), (1.7), (1.8)]) and the blockwise Navarro/Galois Alperin weight conjecture (see [14, Conjecture 2]) hold for RoCK blocks of $\widetilde{S}_{n}^{+}$ and $\widetilde{A}_{n}$, and if $p\equiv 1~{\rm mod}~4$, these conjectures also hold for RoCK blocks of $\widetilde{S}_{n}^{-}$.
-The blockwise Navarro Alperin weight conjecture has been proved for all blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ in [8] using different method.
-Let $\mathbb{P}$ be the set of prime numbers, and $\mathbb{P}_{1}:=\{p\in\mathbb{P}\mid p\equiv 1~{\rm mod}~4\}$.
-Now consider the following three families of group-prime pairs: $\Omega_{1}:=\{(\widetilde{S}_{n}^{+},p)\mid n\in\mathbb{N},~p\in\mathbb{P}\}$, $\Omega_{2}:=\{(\widetilde{A}_{n},p)\mid n\in\mathbb{N},~p\in\mathbb{P}\}$, $\Omega_{3}:=\{(\widetilde{S}_{n}^{-},p)\mid n\in\mathbb{N},~p\in\mathbb{P}_{1}\}$.
+Refinements of the Alperin–McKay conjecture by Isaacs–Navarro, Navarro, Turull and Boltje (see [1, (1.5), (1.6), (1.7), (1.8)]) and the blockwise Navarro/Galois Alperin weight conjecture (see [15, Conjecture 2]) hold for RoCK blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$.
+The blockwise Navarro Alperin weight conjecture has been proved for all blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$ in [9] using different method.
 For a partition $\lambda$ we denote by $h(\lambda)$ the number of nonzero parts of $\lambda$.
-Now by the results (Theorems 1.3, 1.5, 1.8) in this paper, to prove Conjecture 1.1 for all spin blocks of these three families, it suffices to prove the following refined version of the Kessar–Schaps conjecture [20]:
+Now by Theorem 1.9 and Theorem 5.10 below, to prove Conjecture 1.1 for all spin blocks of symmetric and alternating groups, it suffices to prove the following refined version of the Kessar–Schaps conjecture [21]:
 Conjecture 1.11.
-Arbitrary spin blocks $B^{\rho,d}_{R_{p}}$ and $B^{\rho^{\prime},d^{\prime}}_{R_{p}}$ of symmetric groups (or spin blocks $(B^{\rho,d}_{R_{p}})_{\bar{0}}$ and $(B^{\rho^{\prime},d^{\prime}}_{R_{p}})_{\bar{0}}$ of alternating groups) are splendidly Rickard equivalent if and only if $d=d^{\prime}$ and $|\rho|-h(\rho)\equiv|\rho^{\prime}|-h(\rho^{\prime})~({\rm mod}~2)$.
+Arbitrary spin blocks $B^{\eta,\rho,d}_{R_{p}}$ and $B^{\eta,\rho^{\prime},d^{\prime}}_{R_{p}}$ of symmetric groups (or spin blocks $(B^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$ and $(B^{\eta,\rho^{\prime},d^{\prime}}_{R_{p}})_{\bar{0}}$ of alternating groups) are splendidly Rickard equivalent if and only if $d=d^{\prime}$ and $|\rho|-h(\rho)\equiv|\rho^{\prime}|-h(\rho^{\prime})~({\rm mod}~2)$.
 If one replaces “splendidly Rickard equivalent” with “derived equivalent” and replaces $R_{p}$ with $k$, the above conjecture is the usual Kessar–Schaps conjecture.
@@
 This paper is organised as follows: In Section 2 we review some relevant results on superalgebras and supermodules which are necessary for later use.
+In Section 3 we prove descent results for twisted group superalgebras and twisted wreath superproducts.
 In Section 4, we prove general descent results for supermodules and superequivalences.
-In Section 4, we formulate convenient reference for results on graded Rickard equivalences and splendid complexes.
-In Section 5 we review some block theory of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$.
-Base on [21], we prove in Section 6 the structure results for Brauer correspondents of spin blocks over $\mathbb{F}_{p}$.
+In Section 5, we formulate convenient reference for some useful results on graded Rickard equivalences and splendid complexes.
+In Section 6 we review some block theory of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$.
+Based on [22], we prove in Section 7 the structure results for Brauer correspondents of spin blocks over $\mathbb{F}_{p}$.
 In Section 8, we analyse the Morita superequivalence bimodule for an RoCK block constructed by Kleshchev and Livesey [23], and then realise this bisupermodule over $\mathbb{Z}_{p}$; we prove the refined Broué conjecture for RoCK spin blocks in the end of Section 8.
@@
 For an algebra $A$, we denote by $A^{\rm op}$ the opposite algebra of $A$, by $A^{\times}$ the set of units in $A$, by $A$-mod the category of finitely generated $A$-modules, and by $K^{b}(A)$ the homotopy category of bound complexes over $A$-mod (see e.g.
-[23, page 117]).
+[24, page 117]).
 Unless otherwise stated, modules in this paper are left modules.
@@
 $kG$) we usually mean an indecomposable direct factor $B$ of $\mathcal{O}G$ (resp.
-$kG$).
-Some times we also mean the correspondent primitive central idempotent.
+$kG$); some times we also mean the corresponding primitive central idempotent.
 Let $R\in\{\mathcal{O},k\}$ and $A$ a finitely generated $R$-algebra.
 A projective cover $(P,\varphi)$ of a finitely generated $A$-module $M$ is a projective $A$-module $P$ and a surjective $A$-module homomorphism $\varphi:P\to M$ such that no proper direct summand of $P$ surjects onto $M$.
-The assumption on $A$ ensure that projective covers of $M$ always exist and are unique up to isomorphism.
+The assumption on $A$ ensures that projective covers of $M$ always exist and are unique up to isomorphism.
 The $A$-module $\Omega_{A}(M):={\rm ker}(\varphi)$ is called the Heller translate of $M$.
@@
 Superalgebras and supermodules.
-We review some definitions and notation from [22, §3].
+We review some definitions and notation from [23, §3].
 Let $R$ be a commutative ring.
@@
 Let $A$ be an $R$-superalgebra.
-If $A_{\bar{1}}\cap A^{\times}\neq\emptyset$, then $A$ is called a superalgebra with superunit, and any $u\in A_{\bar{1}}\cap A^{\times}$ is called a superunit.
+If $A_{\bar{1}}\cap A^{\times}\neq\varnothing$, then $A$ is called a superalgebra with superunit, and any $u\in A_{\bar{1}}\cap A^{\times}$ is called a superunit.
 An $A$-supermodule $M$ is an $A$-module which is also an $R$-superspace such that $A_{\varepsilon}M_{\delta}\subseteq M_{\varepsilon+\delta}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$.
-If $B$ is another $R$-superalgebra, an AA-$B$-bisupermodule module $M$ is an AA-$B$-bimodule which is also an $R$-superspace such that $A_{\varepsilon}M_{\delta}\subseteq M_{\varepsilon+\delta}$ and $M_{\delta}B_{\varepsilon}\subseteq M_{\varepsilon+\delta}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$.
+If $B$ is another $R$-superalgebra, an AA-$B$-bisupermodule module $M$ is an AA-$B$-bimodule which is also an $R$-superspace such that $A_{\varepsilon}M_{\delta}\subseteq M_{\varepsilon+\delta}$ and $M_{\delta}B_{\varepsilon}\subseteq M_{\delta+\varepsilon}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$.
 A homomorphism $f:M\to N$ of $A$-supermodules is an $A$-module homomorphism that is also a homomorphism of $R$-superspaces; a homomorphism of bisupermodules is defined similarly.
@@
 Remark 2.3.
-Note that any finite generated algebra (module, bimodule) $A$ can be regarded as an superalgebra (a supermodule, a bisupermodule) with $A$ as the even part and 0 as the odd part.
-So the above tensor products, as well as most other results in this paper, for superalgebras, supermodules and bisupermodules apply to usual algebras, modules and bimodules.
+Note that any finitely generated algebra (module, bimodule) $A$ can be regarded as a superalgebra (supermodule, bisupermodule) with $A$ as the even part and 0 as the odd part.
+So the above tensor products (as well as most other results in this paper) for superalgebras, supermodules and bisupermodules apply to usual algebras, modules and bimodules.
 2.4.
@@
 $M\otimes_{B}N\cong A$ as $(A,A)$-bisupermodules and $N\otimes_{A}M\cong B$ as $(B,B)$-bisupermodules.
+Here $M\otimes_{B}N$ is an $(A,A)$-bisupermodule with $(M\otimes_{B}N)_{\bar{0}}=M_{\bar{0}}\otimes_{B}N_{\bar{0}}\oplus M_{\bar{1}}\otimes_{B}N_{\bar{1}}$ and $(M\otimes_{B}N)_{\bar{1}}=M_{\bar{0}}\otimes_{B}N_{\bar{1}}\oplus M_{\bar{1}}\otimes_{B}N_{\bar{0}}$.
 2.5.
@@
 We treat $M^{*}$ as a right $A$-supermodule via $(f\cdot a)(m):=f(am)$, for all $a\in A$, $f\in M^{*}$ and $m\in M$, hence also as an $A^{\rm sop}$-module via $(a\cdot f)(m):=(-1)^{|a||f|}f(am)$.
-For more properties of dual supermodules and relevant results on Morita superequivalences, we refer to [22, §3.7].
+For more properties of dual supermodules and relevant results on Morita superequivalences, we refer to [23, §3.7].
 2.6.
@@
 Note that $A_{1_{G}}$ is always a subsuperalgebra of $A$, and $(A_{1_{G}})^{\rm sop}=(A^{\rm sop})_{1_{G}}$.
-If $A$ and $B$ are $G$-graded crossed $R$-superproducts, following [22, (3.5)] we define
+If $A$ and $B$ are $G$-graded crossed $R$-superproducts, following [23, (3.5)] we define
 $(A,B)_{G}:=\sum_{g\in G}A_{g}\otimes_{R}B_{g^{-1}}=\sum_{g\in G}A_{g}\otimes_{R}(B^{\rm sop})_{g}\subseteq A\otimes_{R}B^{\rm sop}.$ (2.2)
@@
 Let $R\in\{\mathcal{O},k\}$ and let $n$ be a positive integer.
-The twisted group superalgebra $\mathcal{T}_{n}^{R}$ over $R$ of the symmetric group $S_{n}$ is the $R$-superalgebra given by odd generators $\tau_{1},\cdots,\tau_{n-1}$ subject to the relations
+In [22, 23], Kleshchev and Livesey defined the twisted group superalgebra $\mathcal{T}_{n,R}$ over $R$ of the symmetric group $S_{n}$ to be the $R$-superalgebra given by odd generators $\tau_{1},\cdots,\tau_{n-1}$ subject to the relations
 $\tau_{r}^{2}=1,\quad\tau_{r}\tau_{s}=-\tau_{s}\tau_{r}\text{ if }|r-s|>1,\quad(\tau_{r}\tau_{r+1})^{3}=1.$
 For each $w\in S_{n}$, choosing a reduced expression $w=s_{r_{1}}\cdots s_{r_{k}}$ in terms of simple transpositions, we define $\tau_{w}:=\tau_{r_{1}}\cdots\tau_{r_{k}}$ (which depends up to a sign on the choice of a reduced expression).
-It is well-known that $\{\tau_{w}\mid w\in S_{n}\}$ is an $R$-basis of $\mathcal{T}_{n}^{R}$ (one way to see that is to use the first isomorphism in (5.2) below).
-Given an $R$-superalgebra AA, following [21, Section 5.1.a], we define the twisted wreath superproduct $A\wr_{s}\mathcal{T}_{n}^{R}$ as the free product $A^{\otimes n}\star\mathcal{T}_{n}^{R}$ of superalgebras subject to the relations
+It is well-known that $\{\tau_{w}\mid w\in S_{n}\}$ is an $R$-basis of $\mathcal{T}_{n,R}$ (one way to see that is to use the first isomorphism in (6.2) below).
+Given an $R$-superalgebra AA, Kleshchev and Livesey [22, Section 5.1.a] defined the twisted wreath superproduct $A\wr_{s}\mathcal{T}_{n,R}$ as the free product $A^{\otimes n}\star\mathcal{T}_{n,R}$ of superalgebras subject to the relations
 $\tau_{r}\,(a_{1}\otimes\dots\otimes a_{n})=(-1)^{\sum_{u\neq r,r+1}|a_{u}|}\left({}^{s_{r}}(a_{1}\otimes\dots\otimes a_{n})\,\tau_{r}\right)$
@@
 Let $R\in\{\mathcal{O},k\}$, $n$ a positive integer and $A$ an $R$-superalgebra. Then
-$A\wr_{s}\mathcal{T}_{n}^{R}=\bigoplus_{w\in S_{n}}A^{\otimes n}\tau_{w},$ (2.3)
-where $A^{\otimes n}$ denotes $\underbrace{A\otimes_{R}\cdots\otimes_{R}A}_{n}$. Hence $A\wr_{s}\mathcal{T}_{n}^{R}$ is an $S_{n}$-graded crossed $R$-superproduct.
+$A\wr_{s}\mathcal{T}_{n,R}=\bigoplus_{w\in S_{n}}A^{\otimes n}\tau_{w},$ (2.3)
+where $A^{\otimes n}$ denotes $\underbrace{A\otimes_{R}\cdots\otimes_{R}A}_{n}$. Hence $A\wr_{s}\mathcal{T}_{n,R}$ is an $S_{n}$-graded crossed $R$-superproduct.
 Proof.
@@
 Let $R^{\prime}\in\{\mathcal{O}^{\prime},k^{\prime}\}$ and let $A^{\prime}=R^{\prime}\otimes_{R}A$.
-Then by [22, (3.8)], $A^{\prime}\wr_{s}\mathcal{T}_{n}^{R^{\prime}}$ is an $S_{n}$-graded crossed $R^{\prime}$-superproduct via
-$A^{\prime}\wr_{s}\mathcal{T}_{n}^{R^{\prime}}=\bigoplus_{w\in S_{n}}(A^{\prime})^{\otimes n}\tau_{w}.$ (2.4)
+Then by [23, (3.8)], $A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$ is an $S_{n}$-graded crossed $R^{\prime}$-superproduct via
+$A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}=\bigoplus_{w\in S_{n}}(A^{\prime})^{\otimes n}\tau_{w}.$ (2.4)
 […]
 Remark 2.9.
-To explain the equality (2.4), [22, (3.8)] uses the element $\sqrt{-2}$; see [22, Proposition 3.22].
+To explain the equality (2.4), [23, (3.8)] uses the element $\sqrt{-2}$; see [23, Proposition 3.22].
 But after the explanation, $\sqrt{-2}$ can be throw away - because $\sqrt{-2}$ would not involved in this equality.
@@
 Keep the notation of Proposition 2.8.
-Given the $S_{n}$-grading in (2.3), recall the definition of $(A\wr_{s}\mathcal{T}_{n}^{R},A\wr_{s}\mathcal{T}_{n}^{R})_{S_{n}}$ from (2.2).
+Given the $S_{n}$-grading in (2.3), recall the definition of $(A\wr_{s}\mathcal{T}_{n,R},A\wr_{s}\mathcal{T}_{n,R})_{S_{n}}$ from (2.2).
 Let $M$ be an $(A,A)$-bisupermodule.
-Following [22, (3.9)], we denote by $M^{\boxtimes n}_{S_{n}}$ the $(A\wr_{s}\mathcal{T}_{d}^{R},A\wr_{s}\mathcal{T}_{n}^{R})_{S_{n}}$-supermodule given by extending the $(A^{\otimes n},A^{\otimes n})$-bisupermodule $M^{\boxtimes n}:=\underbrace{M\boxtimes_{R}\cdots\boxtimes_{R}M}_{n}$, via
+Following [23, (3.9)], we denote by $M^{\boxtimes n}_{S_{n}}$ the $(A\wr_{s}\mathcal{T}_{n,R},A\wr_{s}\mathcal{T}_{n,R})_{S_{n}}$-supermodule given by extending the $(A^{\otimes n},A^{\otimes n})$-bisupermodule $M^{\boxtimes n}:=\underbrace{M\boxtimes_{R}\cdots\boxtimes_{R}M}_{n}$, via
 […]
 for all $m_{1},\cdots,m_{n}\in M$ and $1\leq r\leq n-1$.
-Following [22, (3.10)] we now define
-$M\wr_{s}\mathcal{T}_{n}^{R}:={\rm Ind}_{(A\wr_{s}\mathcal{T}_{n}^{R},A\wr_{s}\mathcal{T}_{n}^{R})_{S_{n}}}^{(A\wr_{s}\mathcal{T}_{n}^{R})\otimes_{R}(A\wr_{s}\mathcal{T}_{n}^{R})^{\rm sop}}(M^{\boxtimes n}_{S_{n}}).$ (2.6)
+Following [23, (3.10)] we now define
+$M\wr_{s}\mathcal{T}_{n,R}:={\rm Ind}_{(A\wr_{s}\mathcal{T}_{n,R},A\wr_{s}\mathcal{T}_{n,R})_{S_{n}}}^{(A\wr_{s}\mathcal{T}_{n,R})\otimes_{R}(A\wr_{s}\mathcal{T}_{n,R})^{\rm sop}}(M^{\boxtimes n}_{S_{n}}).$ (2.6)
 2.11.
 Let $R\in\{\mathcal{O},k\}$, $n$ a positive integer and $A$ an $R$-superalgebra with a superunit $u$.
-If $C_{2}$ is a cyclic group with generator $g$, we can give $A$ the structure of a $C_{2}$-graded crossed $R$-superproduct with graded components $A_{g^{\varepsilon}}=A_{\bar{0}}u^{\varepsilon}=A_{\varepsilon}$ for $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$.
-For any $n\in\mathbb{Z}_{>0}$, following [21, 5.1.c], we consider the wreath product
+If $C_{2}$ is a cyclic group of order $2$ with generator $g$, we can give $A$ the structure of a $C_{2}$-graded crossed $R$-superproduct with graded components $A_{g^{\varepsilon}}=A_{\bar{0}}u^{\varepsilon}=A_{\varepsilon}$ for $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$.
+For any $n\in\mathbb{Z}_{>0}$, following [22, 5.1.c], we consider the wreath product
 $C_{2}\wr S_{n}=\{g_{1}^{\varepsilon_{1}}\cdots g_{n}^{\varepsilon_{n}}w\mid\varepsilon_{1},\dots,\varepsilon_{n}\in\mathbb{Z}/2\mathbb{Z},\,w\in S_{n}\}$
 where $g_{r}$ is the generator of the $r$-th $C_{2}$ factor in the base group $C_{2}^{\times n}$.
-Then we can give $A\wr_{s}\mathcal{T}_{n}^{R}$ the structure of a $(C_{2}\wr S_{n})$-graded crossed $R$-superproduct with graded components
-$(A\wr_{s}\mathcal{T}_{n})_{g_{1}^{\varepsilon_{1}}\cdots g_{n}^{\varepsilon_{n}}w}=A_{\bar{0}}^{\otimes n}u_{1}^{\varepsilon_{1}}\dots u_{n}^{\varepsilon_{n}}\tau_{w}$
+Then we can give $A\wr_{s}\mathcal{T}_{n,R}$ the structure of a crossed product of $A_{\bar{0}}^{\otimes n}$ and $C_{2}\wr S_{n}$ (see 5.1) with graded components
+$(A\wr_{s}\mathcal{T}_{n,R})_{g_{1}^{\varepsilon_{1}}\cdots g_{n}^{\varepsilon_{n}}w}=A_{\bar{0}}^{\otimes n}u_{1}^{\varepsilon_{1}}\dots u_{n}^{\varepsilon_{n}}\tau_{w}$
 where each $u_{r}$ is $1^{\otimes(r-1)}\otimes u\otimes 1^{\otimes(n-r)}$.
@@
 $(A,B)_{C_{2}}=(A_{\bar{0}}\otimes_{R}B_{\bar{0}}^{\rm sop})\oplus(A_{\bar{1}}\otimes_{R}B_{\bar{1}}^{\rm sop})\subseteq A\otimes_{R}B^{\rm sop},$
-$(A\wr_{s}\mathcal{T}_{n}^{R},B\wr_{s}\mathcal{T}_{n}^{R})_{C_{2}\wr S_{n}}\subseteq(A\wr_{s}\mathcal{T}_{n}^{R})\otimes_{R}(B\wr_{s}\mathcal{T}_{n}^{R})^{\rm sop}.$
-Setting $U_{r}:=(u_{A})_{r}\otimes(u_{B}^{-1})_{r}$, for $1\leq r\leq n$, we have that $A_{\bar{0}}^{\otimes n}\otimes_{R}(B_{\bar{0}}^{\otimes n})^{\rm sop}$ together with the $U_{r}$’s generate a subsuperalgebra of $(A\wr_{s}\mathcal{T}_{n}^{R},B\wr_{s}\mathcal{T}_{n}^{R})_{C_{2}\wr_{s}S_{n}}$ isomorphic to $((A,B)_{C_{2}})^{\otimes n}$.
+$(A\wr_{s}\mathcal{T}_{n,R},B\wr_{s}\mathcal{T}_{n,R})_{C_{2}\wr S_{n}}\subseteq(A\wr_{s}\mathcal{T}_{n,R})\otimes_{R}(B\wr_{s}\mathcal{T}_{n,R})^{\rm sop}.$
+Setting $U_{r}:=(u_{A})_{r}\otimes(u_{B}^{-1})_{r}$, for $1\leq r\leq n$, we have that $A_{\bar{0}}^{\otimes n}\otimes_{R}(B_{\bar{0}}^{\otimes n})^{\rm sop}$ together with the $U_{r}$’s generate a subsuperalgebra of $(A\wr_{s}\mathcal{T}_{n,R},B\wr_{s}\mathcal{T}_{n,R})_{C_{2}\wr_{s}S_{n}}$ isomorphic to $((A,B)_{C_{2}})^{\otimes n}$.
 We identify the two superalgebras.
-Then $(A\wr_{s}\mathcal{T}_{n}^{R},B\wr_{s}\mathcal{T}_{n}^{R})_{C_{2}\wr_{s}S_{n}}$ is generated by $((A,B)_{C_{2}})^{\otimes n}$ and $T_{r}:=\tau_{r}\otimes\tau_{r}^{-1}$, for $1\leq r\leq n-1$ subject to the relations listed in [21, (5.1.13)].
-Proposition 2.12 (a slight generalisation of [21, Proposition 5.1.14]).
+Then $(A\wr_{s}\mathcal{T}_{n,R},B\wr_{s}\mathcal{T}_{n,R})_{C_{2}\wr_{s}S_{n}}$ is generated by $((A,B)_{C_{2}})^{\otimes n}$ and $\{T_{r}:=\tau_{r}\otimes\tau_{r}^{-1}\mid 1\leq r\leq n-1\}$ subject to the relations listed in [22, (5.1.13)].
+Proposition 2.12 (a slight generalisation of [22, Proposition 5.1.14]).
 Let $R\in\{\mathcal{O},k\}$, let $n$ be a positive integer, and let $A$ and $B$ be $R$-superalgebras with superunit.
-If $X$ is a complex of $A_{\bar{0}}$-$B_{\bar{0}}$-bimodules that extends to a complex of $(A,B)_{C_{2}}$-modules, then $X^{\otimes n}$ extends to a complex of $(A\wr_{s}\mathcal{T}_{n},B\wr_{s}\mathcal{T}_{n})_{C_{2}\wr S_{n}}$-modules.
+If $X$ is a complex of $A_{\bar{0}}$-$B_{\bar{0}}$-bimodules that extends to a complex of $(A,B)_{C_{2}}$-modules, then $X^{\otimes n}$ extends to a complex of $(A\wr_{s}\mathcal{T}_{n,R},B\wr_{s}\mathcal{T}_{n,R})_{C_{2}\wr S_{n}}$-modules.
 […]
 2.13.
-The groups $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$. The double covers of $S_{n}$ (where $n$ is a non-negative integer) are given by
+The groups $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$. The double covers of $S_{n}$ (where $n$ is a non-negative integer) are given by
 $\widetilde{S}_{n}^{+}:=\langle z,t_{1},\cdots,t_{n-1}~|~z^{2}=1,~t_{i}z=zt_{i},~t_{i}^{2}=1,~t_{i}t_{j}=zt_{j}t_{i}~{\rm if}~|i-j|>1,~(t_{i}t_{i+1})^{3}=1\rangle,$
 $\widetilde{S}_{n}^{-}:=\langle z,t_{1},\cdots,t_{n-1}~|~z^{2}=1,~t_{i}z=zt_{i},~t_{i}^{2}=z,~t_{i}t_{j}=zt_{j}t_{i}~{\rm if}~|i-j|>1,~(t_{i}t_{i+1})^{3}=1\rangle.$
-We will use $\widetilde{S}_{n}$ to simultaneously represent both double covers.
-Denote by $\pi_{n}$ the surjective group homomorphism $\widetilde{S}_{n}\to S_{n}$, $z\mapsto 1$, $t_{i}\mapsto s_{i}$, where $s_{i}$ denotes the transposition $(i,i+1)\in S_{n}$.
+We will use $\widetilde{S}^{\eta}_{n}$ ($\eta\in\{\pm\}$) to simultaneously represent both double covers.
+Denote by $\pi_{n}$ the surjective group homomorphism $\widetilde{S}^{\eta}_{n}\to S_{n}$, $z\mapsto 1$, $t_{i}\mapsto s_{i}$, where $s_{i}$ denotes the transposition $(i,i+1)\in S_{n}$.
 Then the double cover of $A_{n}$ is given by $\widetilde{A}_{n}:=\pi_{n}^{-1}(A_{n})$.
+3 Descent of twisted wreath superproducts
+3.1.
+Let $R\in\{\mathcal{O},k\}$ and let $n$ be a positive integer.
+For $\eta\in\{\pm\}$, we define the $\eta$-twisted group superalgebra $\mathcal{T}^{\eta}_{n,R}$ of the symmetric group $S_{n}$ over $R$ to be the $R$-superalgebra given by odd generators $\tau_{1},\cdots,\tau_{n-1}$ subject to the relations
+$\xi_{r}^{2}=\eta(-1)^{\frac{(p-1)}{2}},\quad\xi_{r}\xi_{s}=-\xi_{s}\xi_{r}\text{ if }|r-s|>1,\quad(\xi_{r}\xi_{r+1})^{3}=1.$
+For each $w\in S_{n}$, choosing a reduced expression $w=s_{r_{1}}\cdots s_{r_{k}}$ in terms of simple transpositions, we define $\xi_{w}:=\xi_{r_{1}}\cdots\xi_{r_{k}}$ (which depends up to a sign on the choice of a reduced expression).
+Proposition 3.2 (Descent of $\mathcal{T}_{n,R}$).
+Let $(R,R_{p})\in\{(\mathcal{O},\mathbb{Z}_{p}),(k,\mathbb{F}_{p})\}$.
+If $\sqrt{-1}\in R$, then there is an isomorphism of $R$-superalgebras $\mathcal{T}^{\eta}_{n,R}\cong\mathcal{T}_{n,R}$ sending $\eta^{i}(-1)^{\frac{(p-1)i}{2}}\sqrt{\eta(-1)^{\frac{(p-1)}{2}}}\xi_{i}$ to $\tau_{i}$.
+Hence in this case, we have an isomorphism of $R$-superalgebras $R\otimes_{R_{p}}\mathcal{T}^{\eta}_{n,R_{p}}\cong\mathcal{T}_{n,R}$ sending $\eta^{i}(-1)^{\frac{(p-1)i}{2}}\sqrt{\eta(-1)^{\frac{(p-1)}{2}}}\otimes\xi_{i}$ to $\tau_{i}$.
+In particular, $\{\xi_{w}\mid w\in S_{n}\}$ is an $R_{p}$-basis of $\mathcal{T}_{n,R_{p}}^{\eta}$.
+[…]
+Definition 3.3.
+Given an $R$-superalgebra AA, we define the $\eta$-twisted wreath superproduct $A\wr_{s}\mathcal{T}^{\eta}_{n,R}$ as the free product $A^{\otimes n}\star\mathcal{T}^{\eta}_{n,R}$ of superalgebras subject to the relations
+$\xi_{r}\,(a_{1}\otimes\dots\otimes a_{n})=(-1)^{\sum_{u\neq r,r+1}|a_{u}|}\left({}^{s_{r}}(a_{1}\otimes\dots\otimes a_{n})\,\xi_{r}\right)$
+for all $a_{1},\cdots,a_{n}\in A$ and $1\leq r\leq n-1$, where $s_{r}$ is the transposition $(r,r+1)\in S_{n}$.
+Proposition 3.4.
+Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$ such that $\sqrt{-1}\in\mathcal{O}^{\prime}$.
+Let $(R^{\prime},R)\in\{(\mathcal{O}^{\prime},\mathcal{O}),(k^{\prime},k)\}$, $n$ a positive integer, and $A^{\prime}$ an $R^{\prime}$-superalgebra.
+Then the following hold:
+  1. 1.
+    The isomorphism $\mathcal{T}^{\eta}_{n,R^{\prime}}\cong\mathcal{T}_{n,R^{\prime}}$ in Proposition 3.2 extends to an isomorphism of $R^{\prime}$-superalgebras $A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}}\cong A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$.
+  2. 2.
+    If $A^{\prime}\cong R^{\prime}\otimes_{R}A$ for some $R$-superalgebra AA, then $R^{\prime}\otimes_{R}(A\wr_{s}\mathcal{T}^{\eta}_{n,R})\cong A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$.
+  3. 3.
+    We have a decomposition
+    $A\wr_{s}\mathcal{T}^{\eta}_{n,R}=\bigoplus_{w\in S_{n}}A^{\otimes n}\xi_{w}.$ (3.1)
+    Hence $A\wr_{s}\mathcal{T}^{\eta}_{n,R}$ is an $S_{n}$-graded crossed $R$-superproduct.
+Proof.
+The $\eta^{i}(-1)^{\frac{(p-1)i}{2}}\sqrt{\eta(-1)^{\frac{(p-1)}{2}}}\xi_{i}$-conjugation action and the $\tau_{i}$-conjugation action on an element $(a_{1}\otimes\dots\otimes a_{n})\in A^{\prime\otimes n}$ are the same, hence statement (i) holds.
+By statement (i) we have
+$R^{\prime}\otimes_{R}(A\wr_{s}\mathcal{T}^{\eta}_{n,R})\cong(R^{\prime}\otimes_{R}A)\wr_{s}(R^{\prime}\otimes_{R}\mathcal{T}^{\eta}_{n,R})\cong A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}}\cong A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$
+as $R^{\prime}$-superalgebras, whence (ii). By Proposition 2.8, we have
+$A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}=\bigoplus_{w\in S_{n}}A^{\prime\otimes n}\tau_{w}.$
+Since $\eta^{i}(-1)^{\frac{(p-1)i}{2}}\sqrt{\eta(-1)^{\frac{(p-1)}{2}}}$ is invertible in $R^{\prime}$, by statement (ii), we have
+$A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}}=\bigoplus_{w\in S_{n}}A^{\prime\otimes n}\xi_{w}.$
+[…]
+3.5.
+Keep the notation of Proposition 3.4.
+Given the $S_{n}$-grading in (3.1), recall the definition of $(A\wr_{s}\mathcal{T}^{\eta}_{n,R},A\wr_{s}\mathcal{T}^{\eta}_{n,R})_{S_{n}}$ from (2.2).
+Let $M$ be an $(A,A)$-bisupermodule.
+Analogous to (2.5), we denote by $M^{\boxtimes n}_{S_{n}}$ the $(A\wr_{s}\mathcal{T}^{\eta}_{n,R},A\wr_{s}\mathcal{T}^{\eta}_{n,R})_{S_{n}}$-supermodule given by extending the $(A^{\otimes n},A^{\otimes n})$-bisupermodule $M^{\boxtimes n}:=\underbrace{M\boxtimes_{R}\cdots\boxtimes_{R}M}_{n}$, via
+[…]
+for all $m_{1},\cdots,m_{n}\in M$ and $1\leq r\leq n-1$.
+Analogous to (2.6), we now define
+$M\wr_{s}\mathcal{T}^{\eta}_{n,R}:={\rm Ind}_{(A\wr_{s}\mathcal{T}^{\eta}_{n,R},A\wr_{s}\mathcal{T}^{\eta}_{n,R})_{S_{n}}}^{(A\wr_{s}\mathcal{T}^{\eta}_{n,R})\otimes_{R}(A\wr_{s}\mathcal{T}^{\eta}_{n,R})^{\rm sop}}(M^{\boxtimes n}_{S_{n}}).$ (3.3)
+Proposition 3.6.
+Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$ such that $\sqrt{-1}\in\mathcal{O}^{\prime}$.
+Let $(R^{\prime},R)\in\{(\mathcal{O}^{\prime},\mathcal{O}),(k^{\prime},k)\}$, $n$ a positive integer, $A^{\prime}$ an $R^{\prime}$-superalgebra and $M^{\prime}$ an $(A^{\prime},A^{\prime})$-bisupermodule.
+Through the isomorphism $A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}}\cong A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$, we regard the $(A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}},A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}})$-bisupermodule $M^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$ as an $(A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}},A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}})$-bisupermodule.
+Then the following hold:
+  1. 1.
+    We have an isomorphism of $(A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}},A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}})$-bisupermodules $M^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}}\cong M^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$.
+  2. 2.
+    If $A^{\prime}\cong R^{\prime}\otimes_{R}A$ for some $R$-superalgebra AA and $M^{\prime}\cong R^{\prime}\otimes_{R}M$ for some $(A,A)$-bisupermodule $M$, then $R^{\prime}\otimes_{R}(M\wr_{s}\mathcal{T}^{\eta}_{n,R})\cong M^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$ as $(A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}},A^{\prime}\wr_{s}\mathcal{T}^{\eta}_{n,R^{\prime}})$-bisupermodules.
+[…]
+3.7.
+Let $R\in\{\mathcal{O},k\}$, $n$ a positive integer and $A$ an $R$-superalgebra with a superunit $u$.
+If $C_{2}$ is a cyclic group of order $2$ with generator $g$, we can give $A$ the structure of a $C_{2}$-graded crossed $R$-superproduct with graded components $A_{g^{\varepsilon}}=A_{\bar{0}}u^{\varepsilon}=A_{\varepsilon}$ for $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$.
+For any $n\in\mathbb{Z}_{>0}$, again we consider the wreath product
+$C_{2}\wr S_{n}=\{g_{1}^{\varepsilon_{1}}\cdots g_{n}^{\varepsilon_{n}}w\mid\varepsilon_{1},\dots,\varepsilon_{n}\in\mathbb{Z}/2\mathbb{Z},\,w\in S_{n}\}$
+where $g_{r}$ is the generator of the $r$-th $C_{2}$ factor in the base group $C_{2}^{\times n}$.
+Then we can give $A\wr_{s}\mathcal{T}^{\eta}_{n,R}$ the structure of a crossed product of $A_{\bar{0}}^{\otimes n}$ and $C_{2}\wr S_{n}$ (see 5.1) with graded components
+$(A\wr_{s}\mathcal{T}^{\eta}_{n,R})_{g_{1}^{\varepsilon_{1}}\cdots g_{n}^{\varepsilon_{n}}w}=A_{\bar{0}}^{\otimes n}u_{1}^{\varepsilon_{1}}\dots u_{n}^{\varepsilon_{n}}\tau_{w}$
+where each $u_{r}$ is $1^{\otimes(r-1)}\otimes u\otimes 1^{\otimes(n-r)}$.
+Let $A$ now be a superalgebra with a superunit $u_{A}$ and $B$ be a superalgebra with a superunit $u_{B}$. Recalling (2.2), we have the subsuperalgebras
+$(A,B)_{C_{2}}=(A_{\bar{0}}\otimes_{R}B_{\bar{0}}^{\rm sop})\oplus(A_{\bar{1}}\otimes_{R}B_{\bar{1}}^{\rm sop})\subseteq A\otimes_{R}B^{\rm sop},$
+$(A\wr_{s}\mathcal{T}^{\eta}_{n,R},B\wr_{s}\mathcal{T}^{\eta}_{n,R})_{C_{2}\wr S_{n}}\subseteq(A\wr_{s}\mathcal{T}^{\eta}_{n,R})\otimes_{R}(B\wr_{s}\mathcal{T}^{\eta}_{n,R})^{\rm sop}.$
+Setting $U_{r}:=(u_{A})_{r}\otimes(u_{B}^{-1})_{r}$, for $1\leq r\leq n$, we have that $A_{\bar{0}}^{\otimes n}\otimes_{R}(B_{\bar{0}}^{\otimes n})^{\rm sop}$ together with the $U_{r}$’s generate a subsuperalgebra of $(A\wr_{s}\mathcal{T}^{\eta}_{n,R},B\wr_{s}\mathcal{T}^{\eta}_{n,R})_{C_{2}\wr_{s}S_{n}}$ isomorphic to $((A,B)_{C_{2}})^{\otimes n}$.
+We identify the two superalgebras.
+Then $(A\wr_{s}\mathcal{T}^{\eta}_{n,R},B\wr_{s}\mathcal{T}^{\eta}_{n,R})_{C_{2}\wr_{s}S_{n}}$ is generated by $((A,B)_{C_{2}})^{\otimes n}$ and $\{\Xi_{r}:=\xi_{r}\otimes\xi_{r}^{-1}\mid 1\leq r\leq n-1\}$ subject to the following relations:
+[…]
+where $\underline{a}=a_{1}\otimes\dots\otimes a_{n}\in A_{\bar{0}}^{\otimes n}$ and $\underline{b}=b_{1}\otimes\dots\otimes b_{n}\in(B_{\bar{0}}^{\otimes n})^{\rm op}$.
+Proposition 3.8 (a slight modification of [22, Proposition 5.1.14]).
+Let $R\in\{\mathcal{O},k\}$, let $n$ be a positive integer, and let $A$ and $B$ be $R$-superalgebras with superunit.
+If $X$ is a complex of $A_{\bar{0}}$-$B_{\bar{0}}$-bimodules that extends to a complex of $(A,B)_{C_{2}}$-modules, then $X^{\otimes n}$ extends to a complex of $(A\wr_{s}\mathcal{T}^{\eta}_{n,R},B\wr_{s}\mathcal{T}^{\eta}_{n,R})_{C_{2}\wr S_{n}}$-modules.
+[…]
 4 Descent and supermodules
@@
 […]
-3.2.
+4.2.
 Let $R\subseteq R^{\prime}$ be two commutative rings.
@@
 Let $U^{\prime}$ be an $R^{\prime}$-supermodule.
-We said that $U^{\prime}$ is defined over $R$ or $U^{\prime}$ descends to $R$, if there is an $A$-supermodule $U$ such that $U^{\prime}\cong R^{\prime}\otimes_{R}U$ as $A^{\prime}$-supermodules.
+We say that $U^{\prime}$ is defined over $R$ or $U^{\prime}$ descends to $R$, if there is an $A$-supermodule $U$ such that $U^{\prime}\cong R^{\prime}\otimes_{R}U$ as $A^{\prime}$-supermodules.
 In this case, by the definition of an isomorphism of $R^{\prime}$-supermodules, we have $U^{\prime}_{\bar{0}}\cong R^{\prime}\otimes_{R}U_{\bar{0}}$ and $U^{\prime}_{\bar{1}}\cong R^{\prime}\otimes_{R}U_{\bar{1}}$ as $A^{\prime}_{\bar{0}}$-modules (recall from Proposition 4.1 that $A^{\prime}_{\bar{0}}=R^{\prime}\otimes_{R}A_{\bar{0}}$).
@@
 Let $A$ be an $R$-superalgebra with a superunit, and let $A^{\prime}:=R^{\prime}\otimes_{R}A$.
-Let $U^{\prime}$ be an $R^{\prime}$-supermodule.
+Let $U^{\prime}$ be an $A^{\prime}$-supermodule.
 If the $A^{\prime}_{\bar{0}}$-module $U^{\prime}_{\bar{0}}$ descends to $R$, then $U^{\prime}$ descends to $R$.
 Proof.
-By assumption there is an $A_{\bar{0}}$-module $U_{0}$ such that $U^{\prime}_{\bar{0}}\cong R^{\prime}\otimes_{R}U_{\bar{0}}$ as $A^{\prime}_{\bar{0}}$-module.
+By assumption, there is an $A_{\bar{0}}$-module $U_{0}$ such that $U^{\prime}_{\bar{0}}\cong R^{\prime}\otimes_{R}U_{\bar{0}}$ as $A^{\prime}_{\bar{0}}$-module.
 Let $U:=A\otimes_{A_{\bar{0}}}U_{\bar{0}}$.
 It is easy to check that $U$ is an $A$-supermodule.
-By [22, Lemma 3.7 (i)], we have $U^{\prime}\cong A^{\prime}\otimes_{A^{\prime}_{\bar{0}}}U^{\prime}_{\bar{0}}$.
+By [23, Lemma 3.7 (i)], we have $U^{\prime}\cong A^{\prime}\otimes_{A^{\prime}_{\bar{0}}}U^{\prime}_{\bar{0}}$.
 Hence we have
@@
 Then ${\rm Hom}_{R^{\prime}}(R^{\prime}\otimes_{R}U,R^{\prime})\cong R^{\prime}\otimes_{R}{\rm Hom}_{R}(U,R)$ as right $A^{\prime}$-supermodules.
-Proof. By [23, Corollary 1.12.12], we have an $R^{\prime}$-module isomorphism
+Proof. By [24, Corollary 1.12.12], we have an $R^{\prime}$-module isomorphism
 $R^{\prime}\otimes_{R}{\rm Hom}_{R}(U,R)\cong{\rm Hom}_{R^{\prime}}(R^{\prime}\otimes_{R}U,R^{\prime})$
@@
 Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$.
-Let $A$ and $B$ be $\mathcal{O}$-superalgebras, and let $A^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A$, $B^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}B$.
-Let $M$ be an $(A,B)$-bisupermodule, NN a $(B,A)$-bisupermodule, $M^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}M$ and $N^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}N$.
-Then $M^{\prime}$ and $N^{\prime}$ induce a Morita superequivalence between $A^{\prime}$ and $B^{\prime}$ if and only if $M$ and $N$ induce a Morita superequivalence between $A$ and $B$.
+Let $A$ and $B$ be symmetric $\mathcal{O}$-superalgebras, and let $A^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A$, $B^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}B$.
+Let $M$ be an $(A,B)$-bisupermodule, and let $M^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}M$.
+Then $M^{\prime}$ and $M^{\prime*}$ induce a Morita superequivalence between $A^{\prime}$ and $B^{\prime}$ if and only if $M$ and $M^{*}$ induce a Morita superequivalence between $A$ and $B$.
 (Note that in this paper, all superspaces are defined to be finitely generated.) A similar statement holds with $\mathcal{O}^{\prime}$ and $\mathcal{O}$ replaced by $k^{\prime}$ and $k$ respectively.
-Proof. One direction of the implications is trivial. Now assume that
-$M^{\prime}\otimes_{B^{\prime}}N^{\prime}\cong A^{\prime}~~{\rm and}~~N^{\prime}\otimes_{A^{\prime}}M^{\prime}\cong B^{\prime}$ (3.1)
-as $(A^{\prime},A^{\prime})$-bisupermodules and $(B^{\prime},B^{\prime})$-bisupermodules, respectively. Then by [19, Proposition 4.5],
-$M\otimes_{B}N\cong A~~{\rm and}~~N\otimes_{A}M\cong B$ (3.2)
+Proof.
+One direction of the implications is trivial.
+Now assume that we have an $(A^{\prime},A^{\prime})$-bisupermodule isomorphism $f^{\prime}:M^{\prime}\otimes_{B^{\prime}}M^{\prime*}\cong A^{\prime}$ and a $(B^{\prime},B^{\prime})$-bisupermodule isomorphism $g^{\prime}:M^{\prime*}\otimes_{A^{\prime}}M^{\prime}\cong B^{\prime}$.
+Then by [20, Proposition 4.5], we have an $(A,A)$-bimodule isomorphism $f:M\otimes_{B}M^{*}\cong A$ and a $(B,B)$-bimodule isomorphism $g:M^{*}\otimes_{A}M\cong B$.
+Moreover, by the proof of [20, Proposition 4.5], we have $f^{\prime}={\rm id}_{\mathcal{O}^{\prime}}\otimes f$ as $(A^{\prime},A^{\prime})$-bimodule isomorphisms, and $g^{\prime}={\rm id}_{\mathcal{O}^{\prime}}\otimes g$ as $(B^{\prime},B^{\prime})$-bimodule isomorphisms.
+By Propositions 4.1 and 4.4, we have
+$A^{\prime}_{\bar{0}}=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A_{\bar{0}},~~A^{\prime}_{\bar{1}}=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}A_{\bar{1}},$
+$(M^{\prime}\otimes_{B^{\prime}}M^{\prime*})_{\bar{0}}=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}(M\otimes_{B}M^{\prime*})_{\bar{0}}$
+and
+$(M^{\prime}\otimes_{B^{\prime}}M^{\prime*})_{\bar{1}}=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}(M\otimes_{B}M^{*})_{\bar{1}}.$
 […]
 Remark 4.6.
-Proposition 3.5 can be easily extend to the case for cross products (see 4.1) instead of superalgebras, and graded Morita equivalences (see [26, Definition 3.3]) or graded Rickard equivalences (see 4.1) instead of Morita superequivalences.
+Proposition 4.5 can easily be extended to the case of crossed products (see 5.1) instead of superalgebras, and to graded Morita equivalences (see [26, Definition 3.3]) or graded Rickard equivalences (see 5.1) instead of Morita superequivalences.
 The proof is almost the same.
-3.7.
+4.7.
 Projective supermodules and Heller translates.
 Let $R\in\{\mathcal{O},k\}$ and let $A$ be a superalgebra with a superunit.
-Following [22, §3.4], we call an $A$-supermodule $P$ a projective supermodule if $P$ is isomorphic to a direct summand of $A^{\oplus n}$ for some $n$.
+Following [23, §3.4], we call an $A$-supermodule $P$ a projective supermodule if $P$ is isomorphic to a direct summand of $A^{\oplus n}$ for some $n$.
 A projective cover of an $A$-supermodule is defined exactly like for modules.
@@
 […]
-We refer to [23, Theorem 5.2.1] for the definition of Green correspondents. The following result is a descent criterion for Green correspondents.
-Proposition 3.9.
+We refer to [24, Theorem 5.2.1] for the definition of Green correspondents. The following is a descent criterion for Green correspondents.
+Proposition 4.9 (Descent of Green correspondents).
 Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$.
@@
 Let $U^{\prime}$ be an $\mathcal{O}^{\prime}$-free indecomposable $\mathcal{O}^{\prime}G$-module having $P$ as a vertex, and let $V^{\prime}$ be an $\mathcal{O}^{\prime}H$-module which is the Green correspondent of $U^{\prime}$ with respect to $P$.
-Assume that $k$ is finite.
+Assume that $\mathcal{O}^{\prime}$ is finitely generated as $\mathcal{O}$-module.
 Then $U^{\prime}$ descends to $\mathcal{O}$ if an only if $V^{\prime}$ descends to $\mathcal{O}$.
@@
 Assume that $U^{\prime}$ descends to $\mathcal{O}$, that is, there exists an $\mathcal{O}G$-module $U$ such that $U^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}U$.
-By definition, $V^{\prime}$ is the unique (up to isomorphism) direct summand of ${\rm Res}_{H}^{G}(U^{\prime})$ with vertex $P$.
-By [19, Lemma 6.1], we deduce that there is an indecomposable direct summand of $V$ of ${\rm Res}_{P}^{G}(U)$ such that $V^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}V$.
+By definition, $V^{\prime}$ is up to isomorphism the unique direct summand of ${\rm Res}_{H}^{G}(U^{\prime})$ with vertex $P$.
+By [20, Lemma 5.1], we deduce that there is an indecomposable direct summand $V$ of ${\rm Res}_{P}^{G}(U)$ such that $V^{\prime}\cong\mathcal{O}^{\prime}\otimes_{\mathcal{O}}V$.
 […]
-3.10.
+4.10.
 Super group algebra.
@@
 $(RG)_{\bar{0}}:=RG_{\bar{0}}~~~{\rm and}~~~(RG)_{\bar{1}}:=\Big\{\sum_{g\in G\backslash G_{\bar{0}}}\alpha_{g}g\mid\alpha_{g}\in R\Big\}.$
-For example $G$ and $G_{\bar{0}}$ can be chosen to be $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ ($n\geq 2$) respectively.
+For example $G$ and $G_{\bar{0}}$ can be chosen to be $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$ ($n\geq 2$) respectively.
 For any subgroup $H$ of $G$ we will use $H_{\bar{0}}$ to denote $H\cap G_{\bar{0}}$ and $RH$ will inherit its superalgebra structure from that of $RG$.
@@
 The same applies to relative projectivity with respect to a subgroup and other standard notions of finite group representation theory.
-In [22, Theorem 3.39], Kleshchev and Livesey introduced a super version of the usual Green correspondence:
-Theorem 3.11 ([22, Theorem 3.39]).
+In [23, Theorem 3.39], Kleshchev and Livesey introduced a super version of the usual Green correspondence:
+Theorem 4.11 ([23, Theorem 3.39]).
 Let $P$ be a $p$-subgroup of $G$ and $H$ a subgroup of $G$ such that $N_{G}(P)\subseteq H$ and $H\nsubseteq G_{\bar{0}}$.
@@
 Let $U^{\prime}$ be an $\mathcal{O}^{\prime}$-free module-indecomposable $\mathcal{O}^{\prime}G$-supermodule having $P$ as a vertex, and let $V^{\prime}$ be an $\mathcal{O}^{\prime}H$-module which is the super Green correspondent of $U^{\prime}$.
-Assume that $k$ is finite.
+Assume that $\mathcal{O}^{\prime}$ is finitely generated as $\mathcal{O}$-module.
 Then $U$ descends to $\mathcal{O}$ if an only if $V$ descends to $\mathcal{O}$.
@@
 5 On Rickard equivalences and splendid complexes
-4.1.
+5.1.
 Let $R$ be a commutative ring and $G$ a finite group.
@@
 fully $G$-graded) if, as an $R$-module, AA is a direct sum $A=\oplus_{g\in G}A_{g}$ satisfying $A_{g}A_{h}\subseteq A_{gh}$ (resp.
-$A_{g}A_{h}=A_{gh}$) for all $g,h\in G$; if $A^{\times}\cap A_{g}\neq\varnothing$, then $A$ is called a cross product (of $A_{1}$ and $G$).
+$A_{g}A_{h}=A_{gh}$) for all $g,h\in G$; if $A^{\times}\cap A_{g}\neq\varnothing$, then $A$ is called a crossed product (of $A_{1}$ and $G$).
 Let $A$, $B$ and $C$ be fully $G$-graded $R$-algebras.
@@
 Let $K^{b}(\mbox{$A$-Gr-$B$})$ denote the homotopy category of bound complexes over AA-Gr-$B$; see e.g.
-[23, page 117] for homotopy categories.
+[24, page 117] for homotopy categories.
 Let $M$ be a $G$-graded AA-$B$-bimodule and $N$ a $G$-graded $B$-$C$-bimodule.
@@
 $B$) denotes the complex with only one nonzero term $A$ (resp.
-$B$) in degree 0, and we adopt [23, Definition 1.17.9] for the definition of tensor product of complexes.
+$B$) in degree 0, and we adopt [24, Definition 1.17.9] for the definition of tensor product of complexes.
 Let $A$, $B$ and $C$ be fully $G$-graded symmetric $R$-algebras.
-If there is a complex $X\in K^{b}(\mbox{$A$-Gr-$B$})$ where each term of $X$ is finitely generated projective as left $A$-module and as right $B$-module, such that $X\otimes_{B}X^{*}$ and $A$ are isomorphic in $K^{b}(\mbox{$A$-Gr-$A$})$ and $X^{*}\otimes_{A}X$ and $B$ are isomorphic in $K^{b}(\mbox{$B$-Gr-$B$})$.
-Then $A$ and $B$ are graded Rickard equivalent.
-Using [22, Lemmas 3.6 and 3.19], it is straightforward to check the following:
+If there is a complex (resp.
+module) $X$ in $K^{b}(\mbox{$A$-Gr-$B$})$ (resp.
+AA-Gr-$B$) where each term of $X$ is finitely generated projective as left $A$-module and as right $B$-module, such that $X\otimes_{B}X^{*}$ and $A$ are isomorphic in $K^{b}(\mbox{$A$-Gr-$A$})$ (resp.
+AA-Gr-AA) and $X^{*}\otimes_{A}X$ and $B$ are isomorphic in $K^{b}(\mbox{$B$-Gr-$B$})$ (resp.
+(resp.
+$B$-Gr-$B$)).
+Then $A$ and $B$ are graded Rickard (resp.
+Morita) equivalent.
+Note that if $G=\mathbb{Z}/2\mathbb{Z}$, then a graded Morita equivalence is exactly a Morita superequivalence.
+Using [23, Lemmas 3.6 and 3.19], it is straightforward to check the following:
 Proposition 5.2.
@@
 Let $\Delta(A,B):=\oplus_{g\in G}A_{g}\otimes_{R}B^{\rm op}_{g}$, which is a fully $G$-graded $R$-algebra and which is also an $R$-subalgebra of $A\otimes_{R}B^{\rm op}$.
-Lemma 4.3 ([26, Lemma 2.6]).
-Let $R$ be a commutative ring, $G$ a finite group, $A$ and $B$ fully $G$-graded $R$-algebras.
-There are mutually inverse functors induce a equivalence between the categories $\Delta(A,B)$-mod and AA-Gr-$B$.
-In particular, the categories $K^{b}(\mbox{$\Delta(A,B)$-{\rm mod}})$ and $K^{b}(\mbox{$A$-{\rm Gr}-$B$})$ are equivalent.
-An explicit pair of mutually inverse functors which induces the equivalence in Lemma 4.3 has been given in [26, Lemma 2.6].
 The following theorem is a homotopy category version of [26, Theorem 4.7], which is, as [26, Theorem 4.7], a straightforward consequence of [26, 4.2, 4.3].
@@
     There is an object $X_{1}\in K^{b}(\Delta(A,B))$, and isomorphisms $\alpha_{1}:X_{1}\otimes_{B_{1}}X_{1}^{*}\to A_{1}$ in $K^{b}(\Delta(A,A))$ and $\beta_{1}:X_{1}^{*}\otimes_{A_{1}}X_{1}\to B_{1}$ in $K^{b}(\Delta(B,B))$.
-Moreover, if the above conditions hold, then $X$ and $X_{1}$ correspond to each other via the category equivalence in Lemma 4.3.
+Moreover, if the above conditions hold, then $X_{1}$ is the $1$-component of $X$, and
+$X\cong{\rm Ind}_{\Delta(A,B)}^{A\otimes_{R}B^{\rm op}}(X_{1})\cong A\otimes_{A_{1}}X_{1}\cong X_{1}\otimes_{B_{1}}B.$
 If we only know that a Rickard complex for $A_{1}$ and $B_{1}$ extends to $\Delta(A,B)$, then we need $|G|$ to be invertible in $R$:
@@
   2. 2.
-    There is an object $X_{1}\in K^{b}(\Delta(A,B))$ which, regarded as an object of $K^{b}((A_{1}\otimes_{R}B_{1}^{\rm op}))$, induces a Rickard equivalence between $A_{1}$ and $B_{1}$.
-Moreover, if above conditions hold, then $X$ and $X_{1}$ are correspond to each other via the category equivalence in Lemma 4.3.
+    There is an object $X_{1}\in K^{b}(\Delta(A,B))$ which, regarded as an object of $K^{b}(A_{1}\otimes_{R}B_{1}^{\rm op})$, induces a Rickard equivalence between $A_{1}$ and $B_{1}$.
+Moreover, if above conditions hold, then $X_{1}$ is the $1$-component of $X$, and
+$X\cong{\rm Ind}_{\Delta(A,B)}^{A\otimes_{R}B^{\rm op}}(X_{1})\cong A\otimes_{A_{1}}X_{1}\cong X_{1}\otimes_{B_{1}}B.$
 […]
-4.6.
+$5.$
 Let $G$ and $H$ be finite groups.
@@
 Since Rickard’s paper [30], it has been customary to use the adjective “splendid” to any of the Rickard/derived or Morita equivalences.
-In [1, Remark 1.2], Boltje defined a more general version of splendid Rickard equivalences, which is more flexible and preserves all the good properties of the original version of Rickard [30] and also later definitions (see for instance [24, Definition 9.7.5]).
+In [1, Remark 1.2], Boltje defined a more general version of splendid Rickard equivalences, which is more flexible and preserves all the good properties of the original version of Rickard [30] and also later definitions (see for instance [25, Definition 9.7.5]).
 We adopt Boltje’s definition.
@@
   3. 3.
-    Assume that $p\nmid|G\times H:G_{1}\times H_{1}|$ and $M$ is a/an (complex of) RGRG-RHRH-bimodule(s). If ${\rm Res}_{G_{1}\times H_{1}}^{G\times H}(M)$ is splendid, then $M$ is splendid.
+    Let $Y$ be a subgroup of $G\times H$ containing $G_{1}\times H_{1}$, and $M$ a/an (complex of) $RY$-module(s).
+Assume that $p\nmid|Y:G_{1}\times H_{1}|$ and ${\rm Res}_{G_{1}\times H_{1}}^{Y}(M)$ is splendid.
+Then ${\rm Ind}_{Y}^{G\times H}(M)$ is splendid.
 Proof.
@@
 It remains to prove statement (iii).
-Since $p\nmid|G\times H:G_{1}\times H_{1}|$, $M$ is relatively $G_{1}\times H_{1}$-projective, hence $M$ is isomorphic to a direct summand of
-${\rm Ind}_{G_{1}\times H_{1}}^{G\times H}{\rm Res}_{G_{1}\times H_{1}}^{G\times H}(M);$
+Since $p\nmid|Y:G_{1}\times H_{1}|$, $M$ is relatively $G_{1}\times H_{1}$-projective, hence $M$ is isomorphic to a direct summand of
+${\rm Ind}_{G_{1}\times H_{1}}^{Y}{\rm Res}_{G_{1}\times H_{1}}^{Y}(M);$
+see e.g. [24, Theorem 2.6.2]. It follows that ${\rm Ind}_{Y}^{G\times H}(M)$ is isomorphic to a direct summand of
+${\rm Ind}_{G_{1}\times H_{1}}^{G\times H}{\rm Res}_{G_{1}\times H_{1}}^{Y}(M),$
 […]
-With the definition of “splendid” in 4.6, tensor products and duals of splendid bimodules/complexes are splendid:
+With the definition of “splendid” in 5.5, tensor products and duals of splendid bimodules/complexes are splendid:
 Proposition 5.7.
@@
 Let $G$, $H$ and $L$ finite groups.
-If $M$ is a splendid RGRG-RHRH-bimodule and $N$ is a splendid RHRH-RCRC-bimodule, then $M\otimes_{RH}N$ is a splendid RGRG-RLRL-bimodule, and $M^{*}:={\rm Hom}_{R}(M,R)$ is a splendid RHRH-RGRG-bimodule.
+If $M$ is a splendid RGRG-RHRH-bimodule and $N$ is a splendid RHRH-RLRL-bimodule, then $M\otimes_{RH}N$ is a splendid RGRG-RLRL-bimodule, and $M^{*}:={\rm Hom}_{R}(M,R)$ is a splendid RHRH-RGRG-bimodule.
 […]
-Propositions 4.7 (ii) and 4.8 explain why the definition of “splendid” in 4.6 is more flexible than the versions in [30] and [24, Definition 9.7.5].
+Propositions 5.6 (ii) and 5.7 explain why the definition of “splendid” in 5.5 is more flexible than the versions in [30] and [25, Definition 9.7.5].
 Definition 5.8.
@@
 Then we say that $RGb$ and $RHc$ are splendidly Rickard equivalent, and $X$ is called a splendid Rickard complex.
-4.10.
+Theorem 5.9 (a slight generalisation of [30, Theorem 5.2]).
+Let $G$ and $H$ be finite groups, $b\in Z(\mathcal{O}G)$ and $c\in Z(\mathcal{O}H)$ be nonzero idempotents.
+Let $\bar{b}$ and $\bar{c}$ be the image of $b$ and $c$ in $kG$ and $kH$ respectively.
+If $\bar{X}$ is a splendid Rickard complex of $kG\bar{b}$-$kH\bar{c}$-bimodules.
+Then there is a splendid Rickard complex $X$ of $\mathcal{O}Gb$-$\mathcal{O}Hc$-bimodules such that $\bar{X}\cong k\otimes_{\mathcal{O}}X$.
+Moreover, $X$ is unique up to isomorphism.
+[…]
+The following theorem should be a useful tool for the (refined) Broué conjecture.
+Theorem 5.10.
+Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$.
+Assume that $\mathcal{O}^{\prime}$ is finitely generated as $\mathcal{O}$-module.
+Let $(R^{\prime},R)\in\{(\mathcal{O}^{\prime},\mathcal{O}),(k^{\prime},k)\}$.
+Let $G$ and $H$ be finite groups.
+Assume that $K^{\prime}$ contains a primitive $|G\times H|$-th root of unity.
+Let $b$ be a block of $R^{\prime}G$ and $c$ a block of $R^{\prime}H$.
+Assume that $b\in RG$ and $c\in RH$.
+Assume that $RGb$ and $RHc$ are splendidly Rickard equivalent.
+Then the Brauer correspondent of $RGb$ and the Brauer correspondent of $RHc$ are splendidly Morita equivalent.
+Proof.
+By Theorem 5.9, we may assume that $(R^{\prime},R)=(k^{\prime},k)$.
+Let $X$ be a splendid Rickard complex of $(kGb,kHc)$-bimodules.
+Then $X^{\prime}:=k^{\prime}\otimes_{k}X$ is a splendid Rickard complex of $(k^{\prime}Gb,k^{\prime}Hc)$-bimodules.
+By [2, Theorem 1.6 (a)], the element
+$\gamma:=\sum_{i\in\mathbb{Z}}(-1)^{i}[k^{\prime}\otimes_{k}X_{i}]\in T^{\Delta}(k^{\prime}Gb,k^{\prime}Hc)$
+is a $p$-permutation equivalence between $k^{\prime}Gb$ and $k^{\prime}Hc$; see [2] for the definition of a $p$-permutation equivalence and the meaning of the notation $T^{\Delta}(k^{\prime}Gb,k^{\prime}Hc)$.
+Let $(\Delta\varphi,e\otimes f^{*})$ be a maximal $\gamma$-Brauer pair, where $\Delta\varphi$ is a twisted diagonal subgroup of $G\times H$ (with $\varphi:P\xrightarrow{\cong}Q$ an isomorphism between subgroups $P\leq G$ and $Q\leq H$).
+By [2, Theorem 10.11], $(P,e)$ is a maximal $k^{\prime}Gb$-Brauer pair and $(Q,f)$ is a maximal $k^{\prime}Hc$-Brauer pair.
+Let $b^{\prime}:=\sum_{g\in[N_{G}(P)/N_{G}(P,e)]}{}^{g}e$ and $c^{\prime}:=\sum_{h\in[N_{H}(Q)/N_{H}(Q,f)]}{}^{h}f$.
+Then by [25, Theorem 6.7.6 (iii)], $b^{\prime}$ is the Brauer correspondent of $b$ in $k^{\prime}N_{G}(P)$ and $c^{\prime}$ is the Brauer correspondent of $c$ in $k^{\prime}N_{H}(Q)$.
+Since $b\in kG$ and $c\in kH$, we have $b^{\prime}\in kN_{G}(P)$ and $c^{\prime}\in kN_{H}(Q)$.
+Let $M^{\prime}$ be the maximal module of $\gamma$ (see [2, Definition 14.2]).
+So by [2, Definition 14.2], $M^{\prime}$ is up to isomorphism the unique indecomposable direct summand of the $(k^{\prime}Gb,k^{\prime}Hc)$-bimodule $\bigoplus_{i\in\mathbb{Z}}(k^{\prime}\otimes_{k}X_{i})$ who has vertex $\Delta\varphi$.
+Then by [20, Lemma 5.1], we deduce that there is an indecomposable direct summand $M$ of the $(kGb,kHc)$-bimodule $\bigoplus_{i\in\mathbb{Z}}X_{i}$ with vertex $\Delta\varphi$, such that $M^{\prime}\cong k^{\prime}\otimes_{k}M$.
+By [4, (3.4)], the $kN_{G\times H}(\Delta\varphi)$-module ${\rm Br}_{\Delta\varphi}^{k^{\prime}(G\times H)}(M^{\prime})$ is the Green correspondent of the $k^{\prime}(G\times H)$-module $M^{\prime}$.
+(With the notation of [4], ${\rm Br}_{\Delta\varphi}^{k^{\prime}(G\times H)}(M^{\prime})$ is denoted by $M^{\prime}(\Delta\varphi)$.) Then by [2, Theorem 1.3 (d)], the $p$-permutation $k^{\prime}N_{G}(P)b^{\prime}$-$k^{\prime}N_{H}(Q)c^{\prime}$-bimodule
+${\rm Ind}_{N_{G\times H}(\Delta\varphi)}^{N_{G}(P)\times N_{H}(Q)}({\rm Br}_{\Delta\varphi}^{k^{\prime}(G\times H)}(M^{\prime}))$
+induces a Morita equivalence between $k^{\prime}N_{G}(P)b^{\prime}$ and $k^{\prime}N_{H}(Q)c^{\prime}$.
+Now by [20, Proposition 4.5 (c)], the $p$-permutation $kN_{G}(P)b^{\prime}$-$kN_{H}(Q)c^{\prime}$-bimodule
+${\rm Ind}_{N_{G\times H}(\Delta\varphi)}^{N_{G}(P)\times N_{H}(Q)}({\rm Br}_{\Delta\varphi}^{k(G\times H)}(M))$
+[…]
+Remark 5.11.
+In fact, Theorem 5.10 holds in greater generality: we can assume that $RGb$ and $RHc$ are $p$-permutation equivalent to obtain the same result. The proof is entirely analogous.
+5.12.
 Let $R\in\{\mathcal{O},k\}$, and let $G$ and $H$ be finite groups.
@@
 Here $_{\varphi}RH$ is the RPRP-RHRH-bimodule which is equal to $RH$ as a right $RH$-module endowed with the left structure homomorphism $P\xrightarrow{\varphi}Q\to{\rm End}_{R}(RH)$; the notation $RG_{\varphi^{-1}}$ is similarly defined.
-More generally, as in [11, Definition 1.3], we can define splendid modules, complexes, and Rickard equivalences for symmetric interior $P$-algebras:
+More generally, as in [12, Definition 1.3], we can define splendid modules, complexes, and Rickard equivalences for symmetric interior $P$-algebras:
 Definition 5.13.
@@
 $(A_{1}\otimes_{RQ_{1}}{}_{\varphi_{1}}B_{1})\boxtimes_{R}(A_{2}\otimes_{RQ_{2}}{}_{\varphi_{2}}B_{2})$
-is a splendid $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bimodule. By [22, Lemma 3.6], we have an isomorphism of $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bimodules
+is a splendid $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bimodule. By [23, Lemma 3.6], we have an isomorphism of $(A_{1}\otimes_{R}A_{2},B_{1}\otimes_{R}B_{2})$-bimodules
 […]
 […]
-5 On spin blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$
-Notation 5.1.
+6 On spin blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$
+Notation 6.1.
 Let $n$ be a positive integer.
 We use $[n]$ to denote the set $\{1,\cdots,n\}$.
-For conventional convenience we set $0]:=\emptyset$.
-We denote by $\mathscr{P}(n)$ the set of partitions of $n$ and by $\mathscr{P}_{0}(n)$ the set of strict partitions of $n$, i.e.
+For conventional convenience we set $0]:=\varnothing$.
+We denote by $\mathscr{P}_{0}(n)$ the set of strict partitions of $n$, i.e.
 partitions of $n$ without repeat parts.
 Set $\mathscr{P}_{0}(0):=\{{\rm empty~partion}\}$.
-For $\lambda\in\mathscr{P}_{0}(n)$, ${\rm Ab}_{\lambda}$ will denote its $\bar{p}$-abacus ($p$-bar-abacus) display; see [21, §2.3.b] for more details on this.
+For $\lambda\in\mathscr{P}_{0}(n)$, ${\rm Ab}_{\lambda}$ will denote its $\bar{p}$-abacus ($p$-bar-abacus) display; see [22, §2.3.b] for more details on this.
 For any non-empty partition $\lambda$, set $h(\lambda):={\rm max}\{k~|~\lambda_{k}>0\}$ to be the length of $\lambda$.
@@
 We can also define a non-negative integer called the $\bar{p}$-weight of $\lambda$: ${\rm wt}(\lambda):=(|\lambda|-|{\rm core}(\lambda)|)/p\in\mathbb{N}$.
-Assume in this section that $(K,\mathcal{O},k)$ is large enough for $\widetilde{S}_{n}$; this means that $K$ (and hence $\mathcal{O}$) contains a primitive $(2n!)$-th root of unity.
-We also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$.
-Let $R\in\{\mathcal{O},k\}$, and $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$.
-If $n>1$, we consider $R\widetilde{S}_{n}$ as a superalgebra corresponding to $(\widetilde{S}_{n})_{\bar{0}}=\widetilde{A}_{n}$; see 3.10.
+Assume in this section that $(K,\mathcal{O},k)$ is large enough for $\widetilde{S}^{\eta}_{n}$; this means that $K$ (and hence $\mathcal{O}$) contains a primitive $(2n!)$-th root of unity.
+Following [22, 23] we also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$.
+Let $(R,R_{p})\in\{(\mathcal{O},\mathbb{Z}_{p}),(k,\mathbb{F}_{p})\}$.
+If $n>1$, we consider $R\widetilde{S}^{\eta}_{n}$ as a superalgebra corresponding to $(\widetilde{S}^{\eta}_{n})_{\bar{0}}=\widetilde{A}_{n}$; see 4.10.
 We set
-$e_{z}:=(1-z)/2\in R\widetilde{S}_{n}.$
-We will use the adjective “spin” when referring to objects associated to $R\widetilde{S}_{n}e_{z}$ rather than the whole of $R\widetilde{S}_{n}$.
-For example the spin characters of $\widetilde{S}_{n}$ will refer to the $K$-characters $\chi$ of $\widetilde{S}_{n}$ satisfying $\chi(z)=-\chi(1)$, while spin blocks refers to the blocks of $R\widetilde{S}_{n}e_{z}$.
-Note that $R\widetilde{S}_{n}e_{z}$ inherits the structure of superalgebra from $R\widetilde{S}_{n}$ since the idempotent $e_{z}$ is in $(R\widetilde{S}_{n})_{\bar{0}}$.
-Let $G\leq\widetilde{S}_{n}$ with $z\in G$. Following the notation from 3.10, we set
+$e_{z}:=(1-z)/2\in R\widetilde{S}^{\eta}_{n}.$
+We will use the adjective “spin” when referring to objects associated to $R\widetilde{S}^{\eta}_{n}e_{z}$ rather than the whole of $R\widetilde{S}^{\eta}_{n}$.
+For example the spin characters of $\widetilde{S}^{\eta}_{n}$ will refer to the $K$-characters $\chi$ of $\widetilde{S}^{\eta}_{n}$ satisfying $\chi(z)=-\chi(1)$, while spin blocks refers to the blocks of $R\widetilde{S}^{\eta}_{n}e_{z}$.
+Note that $R\widetilde{S}^{\eta}_{n}e_{z}$ inherits the structure of superalgebra from $R\widetilde{S}^{\eta}_{n}$ since the idempotent $e_{z}$ is in $(R\widetilde{S}^{\eta}_{n})_{\bar{0}}$.
+Let $G\leq\widetilde{S}^{\eta}_{n}$ with $z\in G$. Following the notation from 4.10, we set
 $G_{\bar{0}}:=G\cap\widetilde{A}_{n}\quad\text{and}\quad G_{\bar{1}}:=G\setminus\widetilde{A}_{n}$ (6.1)
-and treat $RGe_{z}$ as a superalgebra in the appropriate way. (Note that unlike in 3.10 we are yet not assuming that $G_{\bar{0}}$ is a proper subgroup of $G$.)
-As stated in [22, (4.3)], we have the following isomorphisms of $R$-superalgebras (see 2.7 for the notation $\mathcal{T}_{n}^{R}$):
-$\begin{split}R\widetilde{S}_{n}^{+}e_{z}&\cong\mathcal{T}_{n}^{R},~t_{i}e_{z}\mapsto\tau_{i},\ ze_{z}\mapsto-1\\ R\widetilde{S}_{n}^{-}e_{z}&\cong\mathcal{T}_{n}^{R},~t_{i}e_{z}\mapsto(-1)^{i}\sqrt{-1}\tau_{i},\ ze_{z}\mapsto-1.\end{split}$ (5.2)
-Remark 5.2.
-If $p\equiv 1~{\rm mod}~4$, then $\mathbb{Z}_{p}$ (resp.
-$\mathbb{F}_{p}$) contains a primitive $4$-th root of unity.
-Since $R$ is a flat $R_{p}$-module (see [19, Lemma 4.1]), the $R$-superalgebra isomorphisms (5.2) restrict to isomorphisms of $R_{p}$-superalgebras
-$\begin{split}R_{p}\widetilde{S}_{n}^{+}e_{z}&\cong\mathcal{T}_{n}^{R_{p}},~t_{i}e_{z}\mapsto\tau_{i},\ ze_{z}\mapsto-1\\ R_{p}\widetilde{S}_{n}^{-}e_{z}&\cong\mathcal{T}_{n}^{R_{p}},~t_{i}e_{z}\mapsto(-1)^{i}\sqrt{-1}\tau_{i},\ ze_{z}\mapsto-1.\end{split}$
-If $p\equiv 3~{\rm mod}~4$, we still have the first $R_{p}$-superalgebra isomorphism but the second does not hold, because this time $R_{p}$ does not contain any primitive $4$-th root of unity.
-5.3.
+and treat $RGe_{z}$ as a superalgebra in the appropriate way. (Note that unlike in 4.10 we are yet not assuming that $G_{\bar{0}}$ is a proper subgroup of $G$.)
+As stated in [23, (4.3)], we have the following isomorphisms of $R$-superalgebras (see 2.7 for the notation $\mathcal{T}_{n,R}$):
+$\begin{split}R\widetilde{S}_{n}^{+}e_{z}&\cong\mathcal{T}_{n,R},~t_{i}e_{z}\mapsto\tau_{i},\ ze_{z}\mapsto-1\\ R\widetilde{S}_{n}^{-}e_{z}&\cong\mathcal{T}_{n,R},~t_{i}e_{z}\mapsto(-1)^{i}\sqrt{-1}\tau_{i},\ ze_{z}\mapsto-1.\end{split}$ (6.2)
+6.2.
 […]
-As in [22], if $X\subseteq[n]$, then $S_{X}$ will signify the subgroup of $S_{n}$ consisting of all elements that fix $\X[n]\backslash$ X pointwise and $\widetilde{S}_{X}$ will denote $\pi_{n}^{-1}(S_{X})\leq\widetilde{S}_{n}$.
-Moreover, for disjoint $X_{1},\cdots,X_{i}\subseteq[n]$, let $\widetilde{S}_{X_{1},\cdots,X_{r}}$ be the subgroup of $\widetilde{S}_{n}$ generated by $\widetilde{S}_{X_{1}},\cdots,\widetilde{S}_{X_{i}}$.
+As in [23], if $X\subseteq[n]$, then $S_{X}$ will signify the subgroup of $S_{n}$ consisting of all elements that fix $\X[n]\backslash$ X pointwise and $\widetilde{S}^{\eta}_{X}$ will denote $\pi_{n}^{-1}(S_{X})\leq\widetilde{S}^{\eta}_{n}$.
+Moreover, for disjoint $X_{1},\cdots,X_{i}\subseteq[n]$, let $\widetilde{S}^{\eta}_{X_{1},\cdots,X_{i}}$ be the subgroup of $\widetilde{S}^{\eta}_{n}$ generated by $\widetilde{S}^{\eta}_{X_{1}},\cdots,\widetilde{S}^{\eta}_{X_{i}}$.
 In particular,
-$\widetilde{S}_{X_{1},\cdots,X_{i}}\cong\widetilde{S}_{X_{1}}\times_{z}\cdots\times_{z}\widetilde{S}_{X_{i}}$
+$\widetilde{S}^{\eta}_{X_{1},\cdots,X_{i}}\cong\widetilde{S}^{\eta}_{X_{1}}\times_{z}\cdots\times_{z}\widetilde{S}^{\eta}_{X_{i}}$
 and
-$R\widetilde{S}_{X_{1},\cdots,X_{i}}e_{z}\cong R\widetilde{S}_{X_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}_{X_{i}}e_{z}\cong\mathcal{T}_{|X_{1}|}^{R}\otimes_{R}\cdots\otimes_{R}\mathcal{T}_{|X_{i}|}^{R}$ (5.3)
+$R\widetilde{S}^{\eta}_{X_{1},\cdots,X_{i}}e_{z}\cong R\widetilde{S}^{\eta}_{X_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}^{\eta}_{X_{i}}e_{z}$ (6.3)
 as superalgebras. We also define the corresponding subgroups of $\widetilde{A}_{n}$:
-$\widetilde{A}_{X}:=\widetilde{S}_{X}\cap\widetilde{A}_{n}\quad\text{and}\quad\widetilde{A}_{X_{1},\cdots,X_{i}}:=\widetilde{S}_{X_{1},\cdots,X_{i}}\cap\widetilde{A}_{n}.$
+$\widetilde{A}_{X}:=\widetilde{S}^{\eta}_{X}\cap\widetilde{A}_{n}\quad\text{and}\quad\widetilde{A}_{X_{1},\cdots,X_{i}}:=\widetilde{S}^{\eta}_{X_{1},\cdots,X_{i}}\cap\widetilde{A}_{n}.$
 In the important special case where $n=n_{1}+\cdots+n_{i}$ and $Y_{l}=[n_{1}+\cdots+n_{l}]\setminus[n_{1}+\cdots+n_{l-1}]$ for $l=1,\cdots,r$, we have the standard Young subgroups
-$\widetilde{S}_{n_{1},\cdots,n_{i}}:=\widetilde{S}_{Y_{1},\cdots,Y_{i}}\cong\widetilde{S}_{n_{1}}\times_{z}\cdots\times_{z}\widetilde{S}_{n_{i}}\leq\widetilde{S}_{n}\quad\text{and}\quad\widetilde{A}_{n_{1},\cdots,n_{i}}=\widetilde{S}_{n_{1},\cdots,n_{i}}\cap\widetilde{A}_{n}\leq\widetilde{A}_{n}.$
-If some of the $n_{l}$’s are equal to each other, we use the exponential notation, as for example for the standard Young subgroup $\widetilde{S}_{r+ip,p^{d-i}}=\widetilde{S}_{r+ip,p,\cdots,p}$.
-Following Schur [32], to every strict $\lambda\in\mathscr{P}_{0}(n)$, one associated canonically a spin character $\chi_{\lambda}$ of $\widetilde{S}_{n}$ such that $\chi_{\lambda}$ is irreducible if $n-h(\lambda)$ is even and $\chi_{\lambda}=\chi_{\lambda}^{+}+\chi_{\lambda}^{-}$ for irreducible characters $\chi_{\lambda}^{\pm}$ if $n-h(\lambda)$ is odd.
+$\widetilde{S}^{\eta}_{n_{1},\cdots,n_{i}}:=\widetilde{S}^{\eta}_{Y_{1},\cdots,Y_{i}}\cong\widetilde{S}^{\eta}_{n_{1}}\times_{z}\cdots\times_{z}\widetilde{S}^{\eta}_{n_{i}}\leq\widetilde{S}^{\eta}_{n}\quad\text{and}\quad\widetilde{A}_{n_{1},\cdots,n_{i}}=\widetilde{S}^{\eta}_{n_{1},\cdots,n_{i}}\cap\widetilde{A}_{n}\leq\widetilde{A}_{n}.$
+If some of the $n_{l}$’s are equal to each other, we use the exponential notation, as for example for the standard Young subgroup $\widetilde{S}^{\eta}_{r+ip,p^{d-i}}=\widetilde{S}^{\eta}_{r+ip,p,\cdots,p}$.
+Following Schur [32], to every strict $\lambda\in\mathscr{P}_{0}(n)$, one associated canonically a spin character $\chi_{\lambda}$ of $\widetilde{S}^{\eta}_{n}$ such that $\chi_{\lambda}$ is irreducible if $n-h(\lambda)$ is even and $\chi_{\lambda}=\chi_{\lambda}^{+}+\chi_{\lambda}^{-}$ for irreducible characters $\chi_{\lambda}^{\pm}$ if $n-h(\lambda)$ is odd.
 Moreover,
 $\{\chi_{\lambda}\mid\lambda\in\mathscr{P}_{0}(n),~n-h(\lambda)~{\rm is~even}\}\cup\{\chi_{\lambda}^{\pm}\mid\lambda\in\mathscr{P}_{0}(n),~n-h(\lambda)~{\rm is~odd}\}$
-is a complete irredundant set of irreducible spin characters of $\widetilde{S}_{n}$.
-Theorem 5.4 ([17], see also [21, Theorem 5.2.4]).
+is a complete irredundant set of irreducible spin characters of $\widetilde{S}^{\eta}_{n}$.
+Theorem 6.3 ([18], see also [22, Theorem 5.2.4]).
 With one exception, $\chi_{\lambda}^{\pm}$ and $\chi_{\mu}^{\pm}$ are in the same $p$-block if and only if ${\rm core}(\lambda)={\rm core}(\mu)$.
 The exception is $\lambda=\mu$ is a $\bar{p}$-core and $n-h(\lambda)$ is odd, in which case $\chi_{\lambda}^{+}$ and $\chi_{\lambda}^{-}$ are in different $p$-blocks.
-The blocks of $R\widetilde{S}_{n}$ is described by the following theorem:
-Theorem 5.5 ([17, Theorem 1.1], see also [22, Theorem 4.3]).
+The blocks of $R\widetilde{S}^{\eta}_{n}$ is described by the following theorem:
+Theorem 6.4 ([18, Theorem 1.1], see also [23, Theorem 4.3]).
   1. 1.
-    The spin blocks of $R\widetilde{S}_{n}$ are labelled by pairs $(\rho,d)$, where $\rho$ is a $\bar{p}$-core and $d$ is a non-negative integer with $n=|\rho|+dp$.
-If $d=0$ and $n-h(\rho)$ is odd then $(\rho,0)$ labels two spin blocks of $R\widetilde{S}_{n}$ with corresponding block idempotents $e_{\rho,0}^{+},e_{\rho,0}^{-}\in R\widetilde{S}_{n}e_{z}$.
-In all cases we denote the block idempotent (or sum of two block idempotents) corresponding to $(\rho,d)$ by $e_{\rho,d}\in R\widetilde{S}_{n}e_{z}$.
+    The spin blocks of $R\widetilde{S}^{\eta}_{n}$ are labelled by pairs $(\rho,d)$, where $\rho$ is a $\bar{p}$-core and $d$ is a non-negative integer with $n=|\rho|+dp$.
+If $d=0$ and $n-h(\rho)$ is odd then $(\rho,0)$ labels two spin blocks of $R\widetilde{S}^{\eta}_{n}$ with corresponding block idempotents $e_{\rho,0}^{+},e_{\rho,0}^{-}\in R\widetilde{S}^{\eta}_{n}e_{z}$.
+In all cases we denote the block idempotent (or sum of two block idempotents) corresponding to $(\rho,d)$ by $e_{\rho,d}\in R\widetilde{S}^{\eta}_{n}e_{z}$.
   2. 2.
-    The character $\chi_{\lambda}^{\pm}$ lies in the (sum of) block(s) $\mathcal{O}\widetilde{S}_{n}e_{\rho,d}$, where $\rho$ is the $\bar{p}$-core of $\lambda$.
-If $d=0$ and $\rho$ is odd, we choose the labeling so that $\chi_{\rho}^{\pm}$ lies in the block $\mathcal{O}\widetilde{S}_{n}e_{\rho,0}^{\pm}$.
-Defect groups and Brauer correspondents of blocks of $R\widetilde{S}_{n}$ are described as follows:
-Theorem 5.6 ([5, Theorem A and Corollary 26], see also [22, Theorem 4.3]).
+    The character $\chi_{\lambda}^{\pm}$ lies in the (sum of) block(s) $\mathcal{O}\widetilde{S}^{\eta}_{n}e_{\rho,d}$, where $\rho$ is the $\bar{p}$-core of $\lambda$.
+If $d=0$ and $\rho$ is odd, we choose the labeling so that $\chi_{\rho}^{\pm}$ lies in the block $\mathcal{O}\widetilde{S}^{\eta}_{n}e_{\rho,0}^{\pm}$.
+Defect groups and Brauer correspondents of blocks of $R\widetilde{S}^{\eta}_{n}$ are described as follows:
+Theorem 6.5 ([6, Theorem A and Corollary 26], see also [23, Theorem 4.3]).
 Keep the notation of Theorem 6.4. Fix a $\bar{p}$-core $\rho$ and $d$ a non-negative integer with $n=|\rho|+dp$, and set $U:=[|\rho|]$ and $V:=[n]\setminus U$.
   1. (i)
-    If $d=0$, then each block of $R\widetilde{S}_{n}e_{\rho,d}$ has trivial defect group.
-If $d>0$, then any Sylow $p$-subgroup $D$ of $\widetilde{S}_{V}\leq\widetilde{S}_{n}$ is a defect group of $R\widetilde{S}_{n}e_{\rho,d}$.
+    If $d=0$, then each block of $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$ has trivial defect group.
+If $d>0$, then any Sylow $p$-subgroup $D$ of $\widetilde{S}^{\eta}_{V}\leq\widetilde{S}^{\eta}_{n}$ is a defect group of $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$.
   2. (ii)
-    If $d=0$, then each block of $R\widetilde{S}_{n}e_{\rho,d}$ is its own Brauer correspondent.
-If $d>0$, then, setting $D$ to be as in (i), the Brauer correspondent of $R\widetilde{S}_{n}e_{\rho,d}$ in $N_{\widetilde{S}_{n}}(D)$ is $RN_{\widetilde{S}_{n}}(D)e_{\rho,0}$.
-Here we view $e_{\rho,0}$ as an element of $RN_{\widetilde{S}_{n}}(D)$ via $e_{\rho,0}\in R\widetilde{S}_{U}\hookrightarrow RN_{\widetilde{S}_{n}}(D)$.
+    If $d=0$, then each block of $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$ is its own Brauer correspondent.
+If $d>0$, then, setting $D$ to be as in (i), the Brauer correspondent of $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$ in $N_{\widetilde{S}^{\eta}_{n}}(D)$ is $RN_{\widetilde{S}^{\eta}_{n}}(D)e_{\rho,0}$.
+Here we view $e_{\rho,0}$ as an element of $RN_{\widetilde{S}^{\eta}_{n}}(D)$ via $e_{\rho,0}\in R\widetilde{S}^{\eta}_{U}\hookrightarrow RN_{\widetilde{S}^{\eta}_{n}}(D)$.
 If $\rho=\varnothing$, we interpret $e_{\varnothing,0}$ as $e_{z}$.
-Remark 5.7 ([22, Remark 4.4]).
-The $R\widetilde{S}_{n}e_{\rho,d}$’s are precisely the superblocks of $R\widetilde{S}_{n}e_{z}$.
-That is, the finest decomposition of $R\widetilde{S}_{n}e_{z}$ into a direct sum of two-sided ideals that are invariant under the involution $\sigma_{R\widetilde{S}_{n}}$ (see (2.1) for this notation).
+Remark 6.6 ([23, Remark 4.4]).
+The $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$’s are precisely the superblocks of $R\widetilde{S}^{\eta}_{n}e_{z}$.
+That is, the finest decomposition of $R\widetilde{S}^{\eta}_{n}e_{z}$ into a direct sum of two-sided ideals that are invariant under the involution $\sigma_{R\widetilde{S}^{\eta}_{n}}$ (see (2.1) for this notation).
 This follows from the distribution of characters given in Theorem 6.4 (ii).
 In particular, we always have $e_{\rho,d}\in R\widetilde{A}_{n}e_{z}$.
-The spin blocks of $R\widetilde{A}_{n}$ are described in [18, Proposition 3.16]:
+The spin blocks of $R\widetilde{A}_{n}$ are described in [19, Proposition 3.16]:
 Theorem 6.7.
-For each $n>1$, $\bar{p}$-core $\rho$ and $d$ a non-negative integer with $n=|\rho|+dp$, $R\widetilde{A}_{n}e_{\rho,d}$ is a single block of $R\widetilde{A}_{n}$ unless $d=0$ and $\rho$ is even.
+For each $n>1$, a $\bar{p}$-core $\rho$ and a non-negative integer $d$ with $n=|\rho|+dp$, $R\widetilde{A}_{n}e_{\rho,d}$ is a single block of $R\widetilde{A}_{n}$ unless $d=0$ and $\rho$ is even.
 In this latter case $R\widetilde{A}_{n}e_{\rho,d}$ is a direct sum of two blocks of $R\widetilde{A}_{n}$.
 If $n=1$, then $R\widetilde{A}_{1}e_{(1),0}=R\widetilde{A}_{1}e_{(1),0}=Re_{z}$ is a single block of $R\widetilde{A}_{1}$.
-If $d=0$, once again, the defect group is trivial. In all other cases the defect group of $R\widetilde{A}_{n}e_{\rho,d}$ is the same as that of $R\widetilde{S}_{n}e_{\rho,d}$.
+If $d=0$, once again, the defect group is trivial. In all other cases the defect group of $R\widetilde{A}_{n}e_{\rho,d}$ is the same as that of $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$.
 We set
-$B^{\rho,d}_{R}:=R\widetilde{S}_{n}e_{\rho,d}~~{\rm and}~~(B^{\rho,d}_{R})_{\bar{0}}:=R\widetilde{A}_{n}e_{\rho,d}.$
-It is easy to see that $R\widetilde{A}_{n}e_{\rho,d}$ is indeed the purely even subalgebra of the superalgebra $R\widetilde{S}_{n}e_{\rho,d}$.
+$B^{\eta,\rho,d}_{R}:=R\widetilde{S}^{\eta}_{n}e_{\rho,d}~~{\rm and}~~(B^{\eta,\rho,d}_{R})_{\bar{0}}:=R\widetilde{A}_{n}e_{\rho,d}.$
+It is easy to see that $R\widetilde{A}_{n}e_{\rho,d}$ is indeed the purely even subalgebra of the superalgebra $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$.
 We will often assume $d>0$ to ensure that blocks and superblocks coincide.
@@
 Keep the notation of Theorem 6.4.
-The central idempotent $e_{\rho,d}$ of $R\widetilde{S}_{n}$ is contained in the subring $R_{p}\widetilde{S}_{n}$, hence also contained in $R_{p}\widetilde{A}_{n}$ by Remark 5.7.
-This proposition can be deduced from the following lemma, which is a slight generalisation of [16, Proposition 2.2].
+The central idempotent $e_{\rho,d}$ of $R\widetilde{S}^{\eta}_{n}$ is contained in the subring $R_{p}\widetilde{S}^{\eta}_{n}$, hence also contained in $R_{p}\widetilde{A}_{n}$ by Remark $6.$
+This proposition can be deduced from the following lemma, which is a slight generalisation of [17, Proposition 2.2].
 Lemma 6.9.
@@
 We set
-$B^{\rho,d}_{R_{p}}:=R_{p}\widetilde{S}_{n}e_{\rho,d}~~{\rm and}~~(B^{\rho,d}_{R_{p}})_{\bar{0}}:=R_{p}\widetilde{A}_{n}e_{\rho,d}.$
-As before, it is easy to see that $R_{p}\widetilde{A}_{n}e_{\rho,d}$ is indeed the purely even subalgebra of the superalgebra $R_{p}\widetilde{S}_{n}e_{\rho,d}$.
+$B^{\eta,\rho,d}_{R_{p}}:=R_{p}\widetilde{S}^{\eta}_{n}e_{\rho,d}~~{\rm and}~~(B^{\eta,\rho,d}_{R_{p}})_{\bar{0}}:=R_{p}\widetilde{A}_{n}e_{\rho,d}.$
+As before, it is easy to see that $R_{p}\widetilde{A}_{n}e_{\rho,d}$ is indeed the purely even subalgebra of the superalgebra $R_{p}\widetilde{S}^{\eta}_{n}e_{\rho,d}$.
 Fix a $\bar{p}$-core $\rho$ and a non-negative integer $d$.
@@
 $\underbrace{1,\cdots,r}_{U}\,\overbrace{\underbrace{r+1,\cdots,r+p}_{V_{1}}\,\,\underbrace{r+p+1,\dots,r+2p}_{V_{2}}\,\cdots\,\underbrace{r+p(d-1)+1,\cdots,r+dp}_{V_{d}}}^{V}$ (6.4)
-Let $D_{i}$ is a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$. With this notation, the following now follows easily from Theorem 5.5:
-Lemma 5.11 ([22, Lemma 4.6]).
-Let $D$ be a defect group of the block algebra $R\widetilde{S}_{n}e_{\rho,d}$. Then $D$ is abelian if and only if $d<p$. In this case we can choose $D=D_{1}\times\dots\times D_{d}$.
-5.12.
+Let $D_{i}$ is a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$. With this notation, the following lemma follows easily from Theorem 6.4:
+Lemma 6.10 ([23, Lemma 4.6]).
+Let $D$ be a defect group of the block algebra $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$. Then $D$ is abelian if and only if $d<p$. In this case we can choose $D=D_{1}\times\dots\times D_{d}$.
+6.11.
 Keep the notation in (6.4).
-Let $\iota:S_{d}\to S_{V}$ be the natural permutation action of $S_{d}$ on the set $\{P_{1},\cdots,P_{d}\}$: more precisely, for any $1\leq i\leq d$, $1\leq j\leq p$ and $w\in S_{d}$
+Let $\iota:S_{d}\to S_{V}$ be the natural permutation action of $S_{d}$ on the set $\{V_{1},\cdots,V_{d}\}$: more precisely, for any $1\leq i\leq d$, $1\leq j\leq p$ and $w\in S_{d}$
 $\iota(w)\cdot(r+(i-1)p+j)=r+(w(i)-1)p+j.$
-For each $w\in S_{d}$, choose a lift $T_{w}$ of $\iota(w)$ to $\widetilde{S}_{V}$, and set $T_{i}:=T_{s_{i}}$ for all $1\leq i\leq d-1$.
-Note that every Sylow $p$-subgroup of $S_{V_{i}}$ lifts uniquely and isomorphically to a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$.
+For each $w\in S_{d}$, choose a lift $T_{w}$ of $\iota(w)$ to $\widetilde{S}^{\eta}_{V}$, and set $T_{i}:=T_{s_{i}}$ for all $1\leq i\leq d-1$.
+Note that every Sylow $p$-subgroup of $S_{V_{i}}$ lifts uniquely and isomorphically to a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$.
 Hence we can choose the $D_{i}$’s such that $T_{w}D_{i}T_{w}^{-1}=D_{w(i)}$, for all $1\leq i\leq d$ and $w\in S_{d}$.
-Indeed, one can first construct Sylow $p$-subgroups of each of the $S_{V_{i}}$’s that get permuted by the $\iota(w)$’s and then lift to $\widetilde{S}_{V_{i}}$.
-By the construction of the $T_{w}$’s we also have $T_{w}\widetilde{S}_{V_{i}}T_{w}^{-1}=\widetilde{S}_{V_{w(i)}}$, for all $1\leq i\leq d$ and $w\in S_{d}$.
-Note that $R\widetilde{S}_{V_{1}}e_{z}\cong R\widetilde{S}_{p}e_{z}$ via $t_{r+m}e_{z}\mapsto t_{m}e_{z}$.
-We now fix isomorphisms between the $R\widetilde{S}_{V_{i}}e_{z}$’s.
-For each $1<i\leq d$, we identify $R\widetilde{S}_{V_{i}}e_{z}$ with $R\widetilde{S}_{V_{1}}e_{z}\cong R\widetilde{S}_{p}e_{z}$ via the isomorphism
-$R\widetilde{S}_{V_{1}}e_{z}\to R\widetilde{S}_{V_{i}}e_{z},\quad a\mapsto(-1)^{i|a|}T_{(1,i)}aT_{(1,i)}^{-1}.$ (5.5)
-(This does not depend on our choice of $T_{(1,i)}$, because $T_{(1,i)}$ is uniquely determined up to multiplication by $z$.) Through these isomorphisms and using (5.3), we can identify the superalgebra
-$R\widetilde{S}_{V_{1},\cdots,V_{d}}e_{z}\cong R\widetilde{S}_{V_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}_{V_{d}}e_{z}$ (5.6)
-with $(R\widetilde{S}_{p}e_{z})^{\otimes d}$. Clearly (5.5) and (5.6) still hold with $R$ replaced by $R_{p}$ - because all coefficients involved in these isomorphisms are contained in $R_{p}$.
-5.13.
+Indeed, one can first construct Sylow $p$-subgroups of each of the $S_{V_{i}}$’s that get permuted by the $\iota(w)$’s and then lift to $\widetilde{S}^{\eta}_{V_{i}}$.
+By the construction of the $T_{w}$’s we also have $T_{w}\widetilde{S}^{\eta}_{V_{i}}T_{w}^{-1}=\widetilde{S}^{\eta}_{V_{w(i)}}$, for all $1\leq i\leq d$ and $w\in S_{d}$.
+Note that $R\widetilde{S}^{\eta}_{V_{1}}e_{z}\cong R\widetilde{S}^{\eta}_{p}e_{z}$ via $t_{r+m}e_{z}\mapsto t_{m}e_{z}$.
+We now fix isomorphisms between the $R\widetilde{S}^{\eta}_{V_{i}}e_{z}$’s.
+For each $1<i\leq d$, we identify $R\widetilde{S}^{\eta}_{V_{i}}e_{z}$ with $R\widetilde{S}^{\eta}_{V_{1}}e_{z}\cong R\widetilde{S}^{\eta}_{p}e_{z}$ via the isomorphism
+$R\widetilde{S}^{\eta}_{V_{1}}e_{z}\to R\widetilde{S}^{\eta}_{V_{i}}e_{z},\quad a\mapsto(-1)^{i|a|}T_{(1,i)}aT_{(1,i)}^{-1}.$ (6.5)
+(This does not depend on the choice of $T_{(1,i)}$, because $T_{(1,i)}$ is uniquely determined up to multiplication by $z$.) Through these isomorphisms and using (6.3), we can identify the superalgebra
+$R\widetilde{S}^{\eta}_{V_{1},\cdots,V_{d}}e_{z}\cong R\widetilde{S}^{\eta}_{V_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}^{\eta}_{V_{d}}e_{z}$ (6.6)
+with $(R\widetilde{S}^{\eta}_{p}e_{z})^{\otimes d}$.
+Clearly (6.3), (6.5) and (6.6) still hold with $R$ replaced by $R_{p}$ - because all coefficients involved in these isomorphisms are contained in $R_{p}$.
+Lemma 6.12 (a slight refinement of [23, Lemma 4.7]).
+Keep the notation in 6.11. Up to modifying the choice of $T_{i}$ for $1\leq i\leq d-1$, the set $\{T_{i}\mid 1\leq i\leq d-1\}$ satisfies the following equalities:
+  1. 1.
+    $),(T_{i}e_{z})(a_{1}\otimes\cdots\otimes a_{d})(T_{i}^{-1}e_{z})=(-1)^{\sum_{j\neq i,i+1}|a_{j}|}({}^{s_{i}}(a_{1}\otimes\cdots\otimes a_{d}$)), for all $1\leq i\leq d-1$ and $a_{1}\otimes\dots\otimes a_{d}\in(R\widetilde{S}^{\eta}_{p}e_{z})^{\otimes d}$.
+  2. 2.
+    $(T_{i}e_{z})^{2}=\eta(-1)^{\frac{(p-1)}{2}}e_{z},$ for all $1\leq i\leq d-1$.
+  3. 3.
+    $(T_{i}e_{z})(T_{l}e_{z})=-(T_{l}e_{z})(T_{i}e_{z}),$ for all $1\leq i,l\leq d-1$ with |il|>1|i-l|>1.
+  4. 4.
+    $(T_{i}e_{z}T_{i+1}e_{z})^{3}=(T_{i+1}e_{z}T_{i}e_{z})^{3},$ for all $1\leq i\leq d-2$.
+The result also holds for $R_{p}$ instead of $R$.
+Proof. Equality (i) does not depend on the choice of $T_{i}$’s. Indeed, if $a_{1}\otimes\cdots\otimes a_{d}\in(R\widetilde{S}^{\eta}_{p}e_{z})^{\otimes d}$, we have
+$(zT_{i}e_{z})(a_{1}\otimes\cdots\otimes a_{d})(z^{-1}T_{i}^{-1}e_{z})=T_{i}(a_{1}\otimes\cdots\otimes a_{d})T_{i}^{-1}$
+$=$ $T_{(1,i)}T_{(1,i+1)}T_{(1,i)}(a_{1}\otimes\cdots\otimes a_{d})T_{(1,i)}^{-1}T_{(1,i+1)}^{-1}T_{(1,i)}^{-1}.$
+Equality (i) now follows from a direct calculation using (6.5).
+Equality (iii) also does not depend on the $T_{i}$’s. Noting that $ze_{z}=-e_{z}$, equality (iii) follows immediately from (6.3).
+Recall that for each $1\leq i\leq d-1$, $T_{i}$ is a lift of $\iota(s_{i})$, where
+$\iota(s_{i})=(r+p(i-1)+1,r+pi+1)(r+p(i-1)+2,r+pi+2)\cdots(r+pi,r+pi+p).$
+So $T_{i}^{2}=z^{p(p-1)/2}$ if $\widetilde{S}^{\eta}_{n}=\widetilde{S}_{n}^{+}$, and $T_{i}^{2}=z\cdot z^{p(p-1)/2}$ if $\widetilde{S}^{\eta}_{n}=\widetilde{S}_{n}^{-}$.
+Noting that $ze_{z}=-e_{z}$, we have $(T_{i}e_{z})^{2}=(-1)^{p(p-1)/2}e_{z}$ if $\widetilde{S}^{\eta}_{n}=\widetilde{S}_{n}^{+}$, and $(T_{i}e_{z})^{2}=(-1)\cdot(-1)^{p(p-1)/2}e_{z}$ if $\widetilde{S}^{\eta}_{n}=\widetilde{S}_{n}^{-}$.
+In other words, we have $(T_{i}e_{z})^{2}=\eta(-1)^{\frac{(p-1)}{2}}e_{z}$.
+This proves equality (ii).
+Again, equality (ii) does not depend on the choice of $T_{i}$.
+Finally, since the $\pi_{n}(T_{i})$’s satisfy the braid relations, property (iv) holds up to a $z$, for each $1\leq i\leq d-2$.
+We may therefore change the choice of some of $T_{i}$’s to ensure that equality (iv) holds.
+[…]
+6.13.
 Fix a $\bar{p}$-core $\rho$ and a non-negative integer $d$. We adopt the notation in (6.4), in particular, $r:=|\rho|$ and $n:=r+dp$. Let
-$G$ $:=\widetilde{S}_{n},$
-$L$ $:=\widetilde{S}_{r,p^{d}},$
-$N^{\rho,d}$ $:=N_{G}(\widetilde{S}_{V_{1},\cdots,V_{d}}).$
+$G$ $:=\widetilde{S}^{\eta}_{n},$
+$L$ $:=\widetilde{S}^{\eta}_{r,p^{d}},$
+$N^{\rho,d}$ $:=N_{G}(\widetilde{S}^{\eta}_{V_{1},\cdots,V_{d}}).$
 Note that $N^{\rho,d}$ must permute $V_{1},\cdots,V_{d}$.
-In particular, $N^{\rho,d}\leq\widetilde{S}_{n}$ is the subgroup generated by $L=\widetilde{S}_{U,V_{1},\cdots,V_{d}}$ and the set $\{T_{w}\mid w\in S_{n}\}$ where the notation $T_{w}$ is from 5.12.
+In particular, $N^{\rho,d}\leq\widetilde{S}^{\eta}_{n}$ is the subgroup generated by $L=\widetilde{S}^{\eta}_{U,V_{1},\cdots,V_{d}}$ and the set $\{T_{w}\mid w\in S_{d}\}$ where the notation $T_{w}$ is from 6.11.
 Recall the notation in (6.1) that we will also use the subgroups $G_{\bar{0}}=\widetilde{A}_{n}$, $N_{\bar{0}}^{\rho,d}=N^{\rho,d}\cap\widetilde{A}_{n}$, etc.
 We make the following identifications using (6.3):
-$\begin{split}RLe_{z}&\cong R\widetilde{S}_{U}e_{z}\otimes_{R}R\widetilde{S}_{V_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}_{V_{d}}e_{z}.\end{split}$ (5.7)
+$\begin{split}RLe_{z}&\cong R\widetilde{S}^{\eta}_{U}e_{z}\otimes_{R}R\widetilde{S}^{\eta}_{V_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}^{\eta}_{V_{d}}e_{z}.\end{split}$ (6.7)
 Set $b:=e_{\rho,d}\in RGe_{z}$ and
 $f=:e_{\rho,0}\otimes e_{\varnothing,1}^{(1)}\otimes\cdots\otimes e_{\varnothing,1}^{(d)}\in RLe_{z}\subseteq RN^{\rho,d}e_{z}.$
-By Proposition 5.9, we have $f\in R_{p}L\subseteq R_{p}N^{\rho,d}$. By [22, page 128], the isomorphism
-$RLe_{z}\cong R\widetilde{S}_{r}e_{z}\otimes_{R}(R\widetilde{S}_{p}e_{z})^{\otimes d}$
+By Proposition 6.8, we have $f\in R_{p}L\subseteq R_{p}N^{\rho,d}$. Since $N^{\rho,d}$ is generated by $L$ and $\{T_{w}\mid w\in S_{d}\}$, the isomorphism
+$RLe_{z}\cong R\widetilde{S}^{\eta}_{r}e_{z}\otimes_{R}(R\widetilde{S}^{\eta}_{p}e_{z})^{\otimes d}$
 from (6.7) extends to an isomorphism
-$RN^{\rho,d}e_{z}\cong R\widetilde{S}_{r}e_{z}\otimes_{R}(R\widetilde{S}_{p}e_{z}\wr_{s}\mathcal{T}_{d}^{R}),\ T_{i}e_{z}\mapsto\kappa_{i}^{-1}(1\otimes\tau_{i}),$
-where $\kappa_{i}$’s are from [22, Lemma 4.7].
-More specifically, according to the proof of [22, Lemma 4.7], if $\widetilde{S}_{n}=\widetilde{S}_{n}^{+}$, then $\kappa_{i}\in\{1,-1\}$ for all $1\leq i\leq d-1$, and if $\widetilde{S}_{n}=\widetilde{S}_{n}^{-}$, then $\kappa_{i}\in\{\sqrt{-1},-\sqrt{-1}\}$ for all $1\leq i\leq d-1$.
-In particular, we have the isomorphisms of superalgebras
-$\begin{split}RLf&\cong R\widetilde{S}_{r}e_{\rho,0}\otimes_{R}(R\widetilde{S}_{p}e_{\varnothing,1})^{\otimes d}\cong B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\otimes d}\\ RN^{\rho,d}f&\cong R\widetilde{S}_{r}e_{\rho,0}\otimes_{R}(R\widetilde{S}_{p}e_{\varnothing,1}\wr_{s}\mathcal{T}_{d}^{R})\cong B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R}).\end{split}$ (5.8)
+$RN^{\rho,d}e_{z}\cong R\widetilde{S}^{\eta}_{r}e_{z}\otimes_{R}(R\widetilde{S}^{\eta}_{p}e_{z}\wr_{s}\mathcal{T}^{\eta}_{d,R}),\ T_{i}e_{z}\mapsto(1\otimes\xi_{i}),$
+where $T_{i}$’s are chosen to satisfy Lemma 6.12. In particular, we have the isomorphisms of superalgebras
+$\begin{split}RLf&\cong R\widetilde{S}^{\eta}_{r}e_{\rho,0}\otimes_{R}(R\widetilde{S}^{\eta}_{p}e_{\varnothing,1})^{\otimes d}=B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R})^{\otimes d}\\ RN^{\rho,d}f&\cong R\widetilde{S}^{\eta}_{r}e_{\rho,0}\otimes_{R}(R\widetilde{S}^{\eta}_{p}e_{\varnothing,1}\wr_{s}\mathcal{T}^{\eta}_{d,R})=B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R}).\end{split}$ (6.8)
 Remark 6.14.
-For the group $\widetilde{S}_{n}^{+}$, since $\kappa_{i}=\pm 1\in R_{p}$ and since $R$ is a flat $R_{p}$-module (see [19, Lemma 4.1]), the isomorphisms (5.8) of $R$-superalgebras restricts to isomorphisms of $R_{p}$-superalgebras
-$\begin{split}R_{p}Lf&\cong R_{p}\widetilde{S}_{r}^{+}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{+}e_{\varnothing,1})^{\otimes d}\\ R_{p}N^{\rho,d}f&\cong R_{p}\widetilde{S}_{r}^{+}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{+}e_{\varnothing,1}\wr_{s}\mathcal{T}_{d}^{R_{p}}).\end{split}$
-For the group $\widetilde{S}_{n}^{-}$, $\kappa_{i}=\pm\sqrt{-1}$.
-Then $\kappa_{i}\in R_{p}$ if and only if $p\equiv 1~{\rm mod}~4$.
-Hence if $p\equiv 1~{\rm mod}~4$, the isomorphisms (5.8) of $R$-superalgebras restricts to isomorphisms of $R_{p}$-superalgebras
-$\begin{split}R_{p}Lf&\cong R_{p}\widetilde{S}_{r}^{-}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{-}e_{\varnothing,1})^{\otimes d}\\ R_{p}N^{\rho,d}f&\cong R_{p}\widetilde{S}_{r}^{-}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}_{p}^{-}e_{\varnothing,1}\wr_{s}\mathcal{T}_{d}^{R_{p}}).\end{split}$
-Lemma 5.15 ([22, Lemmas 4.9 and 4.10]).
-Let $0<d<p$. As in 5.12, for $i=1,\cdots,d$, let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$ such that $T_{w}D_{i}T_{w}^{-1}=D_{w(i)}$, for all $i=1,\cdots,d$ and $w\in S_{d}$, and let
-$D:=D_{1}\times\cdots\times D_{d}\leq\widetilde{S}_{V}.$
+Since all coefficients involved in the isomorphisms (6.8) are contained in the subring $R_{p}$ of $R$. The isomorphisms (6.8) of $R$-superalgebras restricts to isomorphisms of $R_{p}$-superalgebras
+$\begin{split}R_{p}Lf&\cong R_{p}\widetilde{S}^{\eta}_{r}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}^{\eta}_{p}e_{\varnothing,1})^{\otimes d}=B^{\eta,\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}})^{\otimes d}\\ R_{p}N^{\rho,d}f&\cong R_{p}\widetilde{S}^{\eta}_{r}e_{\rho,0}\otimes_{R_{p}}(R_{p}\widetilde{S}^{\eta}_{p}e_{\varnothing,1}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})=B^{\eta,\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}}).\end{split}$
+Lemma 6.15 ([23, Lemmas 4.9 and 4.10]).
+Let $0<d<p$.
+As in 6.11, for $i=1,\cdots,d$, let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$ such that $T_{w}D_{i}T_{w}^{-1}=D_{w(i)}$, for all $i=1,\cdots,d$ and $w\in S_{d}$, and let
+$D:=D_{1}\times\cdots\times D_{d}\leq\widetilde{S}^{\eta}_{V}.$
   1. 1.
@@
     $RN^{\rho,d}f$ (resp. $RN_{\bar{0}}^{\rho,d}f$) is the Brauer correspondent of $RGb$ (resp. $RG_{\bar{0}}b$) in $N^{\rho,d}$ (resp. $N_{\bar{0}}^{\rho,d}$).
-5.16[22, Definition 5.1].
+6.16[23, Definition 5.1].
 Rouquier cores and RoCK blocks. Let $d\in\mathbb{N}$. A $d$-Rouquier $\bar{p}$}-core $\rho$ is a $\bar{p}$-core such that ${\rm Ab}_{\rho}$ has the following properties:
@@
     The $(i+1)$-th runner has at least d1d-1 more beads than the $i$-th runner, for $1\leq i\leq(p-3)/2$.
-If $\rho$ is a $d$-Rouquier $\bar{p}$-core, we refer to $B^{\rho,d}_{R}$ as an RoCK block of weight $d$ over $R$.
-For more information on RoCK blocks we refer to [22, §5].
+If $\rho$ is a $d$-Rouquier $\bar{p}$-core, we refer to $B^{\eta,\rho,d}_{R}$ as an RoCK block of weight $d$ over $R$.
+For more information on RoCK blocks we refer to [23, §5].
 7 On Brauer correspondents, and proofs of Theorems 1.3 and 1.5
@@
 We also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$.
-Let $R\in\{\mathcal{O},k\}$ and $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$.
-In this section we continue to adopt the notation in 5.13; we always assume that $d>0$.
-Let $D$ be a Sylow $p$-subgroup of $\widetilde{S}_{V}$.
-Then by Theorem 5.6, $RN_{\widetilde{S}_{n}}(D)e_{\rho,0}$ is the Brauer correspondent of $RGb=R\widetilde{S}_{n}e_{\rho,d}$, and by [21, Theorem 5.2.10], $RN_{\widetilde{A}_{n}}(D)e_{\rho,0}$ is the Brauer correspondent of $RG_{\bar{0}}b=R\widetilde{A}_{n}e_{\rho,d}$.
+Let $(R,R_{p})\in\{(\mathcal{O},\mathbb{Z}_{p}),(k,\mathbb{F}_{p})\}$.
+In this section we continue to adopt the notation in 6.13; we always assume that $d>0$.
+Let $D$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V}$.
+Then by Theorem 6.5, $RN_{\widetilde{S}^{\eta}_{n}}(D)e_{\rho,0}$ is the Brauer correspondent of $RGb=R\widetilde{S}^{\eta}_{n}e_{\rho,d}$, and by [22, Theorem 5.2.10], $RN_{\widetilde{A}_{n}}(D)e_{\rho,0}$ is the Brauer correspondent of $RG_{\bar{0}}b=R\widetilde{A}_{n}e_{\rho,d}$.
 Note that we interpret $e_{\varnothing,0}$ as $e_{z}$.
 We set
-$\begin{split}b^{\rho,d}_{R}&:=RN_{\widetilde{S}_{n}}(D)e_{\rho,0},~~(b^{\rho,d}_{R})_{\bar{0}}:=RN_{\widetilde{A}_{n}}(D)e_{\rho,0},\\ b^{\rho,d}_{R_{p}}&:=R_{p}N_{\widetilde{S}_{n}}(D)e_{\rho,0},~~(b^{\rho,d}_{R_{p}})_{\bar{0}}:=R_{p}N_{\widetilde{A}_{n}}(D)e_{\rho,0}.\end{split}$
-Theorem 6.1 ([21, Proposition 5.2.13]).
+$\begin{split}b^{\eta,\rho,d}_{R}&:=RN_{\widetilde{S}^{\eta}_{n}}(D)e_{\rho,0},~~(b^{\eta,\rho,d}_{R})_{\bar{0}}:=RN_{\widetilde{A}_{n}}(D)e_{\rho,0},\\ b^{\eta,\rho,d}_{R_{p}}&:=R_{p}N_{\widetilde{S}^{\eta}_{n}}(D)e_{\rho,0},~~(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}:=R_{p}N_{\widetilde{A}_{n}}(D)e_{\rho,0}.\end{split}$
+Note that $(b^{\eta,\rho,d}_{R})_{\bar{0}}$ (resp. $(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$) is indeed the even part of the superalgebra $B^{\eta,\rho,d}_{R}$ (resp. $B^{\eta,\rho,d}_{R_{p}}$).
+Theorem 7.1 ([22, Proposition 5.2.13]).
 Suppose that $0<d<p$.
-Let $P$ be a Sylow $p$-subgroup of $\widetilde{S}_{p}$, let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$ chosen as in 5.12, and let $D=D_{1}\times\cdots\times D_{d}$.
+Let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$ chosen as in 6.11, and let $D=D_{1}\times\cdots\times D_{d}$.
 Then we have an isomorphism of $k$-superalgebras
-$b^{\rho,d}_{k}\cong k\widetilde{S}_{U}e_{\rho,0}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k})\cong B^{\rho,0}_{k}\otimes_{k}(b^{\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d}^{k}).$ (6.1)
+$b^{\eta,\rho,d}_{k}\cong k\widetilde{S}^{\eta}_{U}e_{\rho,0}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k})\cong B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,k}).$ (7.1)
 Theorem 7.2.
-Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$.
 Suppose that $0<d<p$.
-Let $P$ be a Sylow $p$-subgroup of $\widetilde{S}_{p}$, let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$ chosen as in 5.12, and let $D=D_{1}\times\cdots\times D_{d}$.
+Let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$ chosen as in 6.11, and let $D=D_{1}\times\cdots\times D_{d}$.
 Then we have an isomorphism of $\mathbb{F}_{p}$-superalgebras
-$b^{\rho,d}_{\mathbb{F}_{p}}\cong\mathbb{F}_{p}\widetilde{S}_{U}e_{\rho,0}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}).$ (6.2)
+$b^{\eta,\rho,d}_{\mathbb{F}_{p}}\cong\mathbb{F}_{p}\widetilde{S}^{\eta}_{U}e_{\rho,0}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})\cong B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}).$ (7.2)
+Proof. By the isomorphism (7.1) and by Proposition 3.4 (i),
+$b^{\eta,\rho,d}_{k}\cong B^{\eta,\rho,0}_{k}\otimes_{k}(b^{\eta,\varnothing,1}_{k}\wr_{s}\mathcal{T}_{d,\eta}^{k})$ (7.3)
 […]
-6.3.
+7.3.
 Keep the notation of Theorem 7.2.
 We have the following isomorphisms all of which are obtained by restricting either the isomorphism (7.2) or the isomorphisms in Remark 6.14 to appropriate subalgebras:
-$\begin{split}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}&\cong b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}\\ \mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}&\cong B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}\end{split}$ (6.3)
-We realise [21, Lemma 5.2.19] over $\mathbb{F}_{p}$:
+$\begin{split}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,0}^{\otimes d}&\cong b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}\\ \mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d}&\cong B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}\end{split}$ (7.4)
+We realise [22, Lemma 5.2.19] over $\mathbb{F}_{p}$:
 Lemma 7.4.
 Let $P$ be a Sylow $p$-subgroup of $\widetilde{A}_{p}$.
-There is a splendid Rickard complex of $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodules that extends to a complex of $(\mathbb{F}_{p}\widetilde{S}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{S}_{p}}(P)e_{\varnothing,0})_{C_{2}}$-modules.
+There is a splendid Rickard complex of $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodules that extends to a complex of $(\mathbb{F}_{p}\widetilde{S}^{\eta}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{S}^{\eta}_{p}}(P)e_{\varnothing,0})_{C_{2}}$-modules.
 Proof.
-In [19, Theorem 1.10], Kessar and Linckelmann construct a splendid Rickard complex for blocks with a cyclic defect group over a non-splitting field, hence a splendid Rickard complex $X$ of $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodules.
-We review the construction of $X$ from the proof of [19, Theorem 1.10].
+In [20, Theorem 1.10], Kessar and Linckelmann construct a splendid Rickard complex for blocks with a cyclic defect group over a non-splitting field, hence a splendid Rickard complex $X$ of $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodules.
+We review the construction of $X$ from the proof of [20, Theorem 1.10].
 Note that $|P|=p$ and $O_{p}(\widetilde{A}_{p})$, the largest $p$-normal subgroup of $\widetilde{A}_{n}$, is the trivial subgroup.
@@
 where $M$ is in degree zero.
-Moreover $M$ is the unique (up to isomorphism) summand of the $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodule $e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}$.
+Moreover $M$ is up to isomorphism the unique nonprojective indecomposable summand of the $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodule $e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}$.
 Let $Z\to M$ be a projective cover of the $(\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})$-bomodule $M$.
 Then $N$ is a direct summand of $Z$ and $\varphi$ is the restriction of the map $Z\to M$ to $N$; moreover, if we write $Z=N\oplus N^{\prime}$, then $N$ and $N^{\prime}$ do not have nonzero isomorphic direct summand.
-As in the proof of [21, Lemma 5.2.19], we transport our setup to the group algebra setting. Set
-$\mathcal{A}=\langle\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P),(s,s)\rangle\leq\widetilde{S}_{p}\times N_{\widetilde{S}_{p}}(P),$
-for some fixed $s\in N_{\widetilde{S}_{p}}(P)\backslash N_{\widetilde{S}_{p}}(P)$. We have an isomorphism of $\mathbb{F}_{p}$-superalgebras
-$\mathbb{F}_{p}\mathcal{A}(e_{\varnothing,1}\otimes e_{\varnothing,0})$ $\cong(\mathbb{F}_{p}\widetilde{S}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{S}_{p}}(P)e_{\varnothing,0})_{C_{2}},$
+As in the proof of [22, Lemma 5.2.19], we transport our setup to the group algebra setting. Set
+$\mathcal{A}=\langle\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P),(s,s)\rangle\leq\widetilde{S}^{\eta}_{p}\times N_{\widetilde{S}^{\eta}_{p}}(P),$
+for some fixed $s\in N_{\widetilde{S}^{\eta}_{p}}(P)\backslash N_{\widetilde{A}_{p}}(P)$. We have an isomorphism of $\mathbb{F}_{p}$-superalgebras
+$\mathbb{F}_{p}\mathcal{A}(e_{\varnothing,1}\otimes e_{\varnothing,0})$ $\cong(\mathbb{F}_{p}\widetilde{S}^{\eta}_{p}e_{\varnothing,1},\mathbb{F}_{p}N_{\widetilde{S}^{\eta}_{p}}(P)e_{\varnothing,0})_{C_{2}},$
 $(a_{1},a_{2})(e_{\varnothing,1}\otimes e_{\varnothing,0})$ $\mapsto a_{1}e_{\varnothing,1}\otimes a_{2}^{-1}e_{\varnothing,0},$
@@
 We identify $\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1}\otimes_{\mathbb{F}_{p}}(\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0})^{\rm sop}$ with the subalgebra $\mathbb{F}_{p}(\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P))(e_{\varnothing,1}\otimes e_{\varnothing,0})$ of $\mathbb{F}_{p}\mathcal{A}(e_{\varnothing,1}\otimes e_{\varnothing,0})$.
+In this setting, we need to show that the complex $X$ from (7.5) extends to a complex of $\mathbb{F}_{p}\mathcal{A}$-modules.
+Clearly the $\mathbb{F}_{p}(\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P))$-module $e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}$ extends to an $\mathbb{F}_{p}\mathcal{A}$-module via $(s,s)\cdot x=sxs^{-1}$ for all $x\in e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}$.
+If we write $e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}=M\oplus M^{\prime}$, then (by the construction of $X$) $M$ is the nonprojective part and $M^{\prime}$ is the projective part of $e_{\varnothing,1}\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,0}$.
+Hence both $M$ and $M^{\prime}$ are $C_{2}$-invariant; that is, $(s,s)M\cong M$ and $(s,s)M^{\prime}\cong M^{\prime}$ as $\mathbb{F}_{p}(\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P))$-modules.
+By [8, Theorem 3.11], $M$ and $M^{\prime}$ extend to $\mathbb{F}_{p}\mathcal{A}(e_{\varnothing,1}\otimes e_{\varnothing,0})$-modules.
+Let $\tilde{M}$ be such an extension of $M$ to $\mathbb{F}_{p}\mathcal{A}(e_{\varnothing,1}\otimes e_{\varnothing,0})$.
+Let $\tilde{Z}$ be a projective cover of the $\mathbb{F}_{p}\mathcal{A}$-module $\tilde{M}$.
+Since $2$ is invertible in $\mathbb{F}_{p}$, we have
+$\tilde{Z}/J(\mathbb{F}_{p}\mathcal{A})\tilde{Z}\cong\tilde{M}/J(\mathbb{F}_{p}\mathcal{A})\tilde{M}\cong\tilde{M}/J(\mathbb{F}_{p}(\widetilde{A}_{p}\times N_{\widetilde{A}_{p}}(P)))\tilde{M}\cong M/{\rm rad}(M)$
 […]
@@
 […]
-In [21, Proposition 5.2.21], Kleshchev and Livesey proved the following result for Brauer correspondents of blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ with abelian defect groups.
-Theorem 6.5 ([21, Proposition 5.2.33]).
-$kN_{\widetilde{S}_{n}}(D)e_{\rho,0}$, i.e.
-$b^{\rho,d}_{k}$, is splendid Rickard equivalent to $kNf$, and $kN_{\widetilde{A}_{n}}(D)e_{\rho,0}$, i.e.
-$(b^{\rho,d}_{k})_{\bar{0}}$, is splendid Rickard equivalent to $kN_{\bar{0}}f$.
+In [22, Proposition 5.2.21], Kleshchev and Livesey proved the following result for Brauer correspondents of blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$ with abelian defect groups.
+Theorem 7.5 ([22, Proposition 5.2.33]).
+Suppose that $0<d<p$.
+Let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$ chosen as in 6.11, and let $D=D_{1}\times\cdots\times D_{d}$.
+Then $kN_{\widetilde{S}^{\eta}_{n}}(D)e_{\rho,0}$, i.e.
+$b^{\eta,\rho,d}_{k}$, is splendid Rickard equivalent to $kNf$, and $kN_{\widetilde{A}_{n}}(D)e_{\rho,0}$, i.e.
+$(b^{\eta,\rho,d}_{k})_{\bar{0}}$, is splendid Rickard equivalent to $kN_{\bar{0}}f$.
 Now we realise Theorem 7.5 over $\mathbb{F}_{p}$:
 Theorem 7.6.
-Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$.
+Suppose that $0<d<p$.
+Let $D_{i}$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$ chosen as in 6.11, and let $D=D_{1}\times\cdots\times D_{d}$.
 Then $\mathbb{F}_{p}N_{G}(D)e_{\rho,0}$, i.e.
-$b^{\rho,d}_{\mathbb{F}_{p}}$, is splendid Rickard equivalent to $\mathbb{F}_{p}N^{\rho,d}f$, and $\mathbb{F}_{p}N_{G}(D)e_{\rho,0}$, i.e.
-$(b^{\rho,d}_{\mathbb{F}_{p}})_{\bar{0}}$, is splendid Rickard equivalent to $\mathbb{F}_{p}N_{\bar{0}}^{\rho,d}f$.
+$b^{\eta,\rho,d}_{\mathbb{F}_{p}}$, is splendid Rickard equivalent to $\mathbb{F}_{p}N^{\rho,d}f$, and $\mathbb{F}_{p}N_{G_{\bar{0}}}(D)e_{\rho,0}$, i.e.
+$(b^{\eta,\rho,d}_{\mathbb{F}_{p}})_{\bar{0}}$, is splendid Rickard equivalent to $\mathbb{F}_{p}N_{\bar{0}}^{\rho,d}f$.
 Proof. By Remark 6.14, we have
-$\mathbb{F}_{p}N^{\rho,d}f\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}).$
-Let $P$ be a Sylow $p$-subgroup of $\widetilde{S}_{p}$.
+$\mathbb{F}_{p}N^{\rho,d}f\cong B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}).$
+Let $P$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{p}$.
 The block algebra $\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1}$ has a cyclic defect group $P$ and Brauer correspondent $\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0}$.
-By Lemma 6.4, there is a splendid Rickard complex $X$ of $((B^{\varnothing,1}_{\mathbb{F}_{p}})_{\bar{0}},(b^{\varnothing,1}_{\mathbb{F}_{p}})_{\bar{0}})$-bimodule that extends to a complex $\tilde{X}$ of $(B^{\varnothing,1}_{\mathbb{F}_{p}},b^{\varnothing,1}_{\mathbb{F}_{p}})_{C_{2}}$-modules.
-Let $H$ be the subgroup of $G$ generated by $\widetilde{S}_{U}$ and all $\widetilde{A}_{V_{i}}$’s.
-Now (by Propositions 4.2, 4.12 and Remark 2.3) $X^{\otimes d}$ induces a splendid Rickard equivalence between
-$\mathbb{F}_{p}(H\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}\cong\mathbb{F}_{p}\widetilde{A}_{V_{1}}e_{\varnothing,1}\otimes_{\mathbb{F}_{p}}\cdots\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}\widetilde{A}_{V_{d}}e_{\varnothing,1}\cong(B^{\varnothing,1}_{\bar{0}})^{\otimes d}$
+By Lemma 7.4, there is a splendid Rickard complex $X$ of $((B^{\eta,\varnothing,1}_{\mathbb{F}_{p}})_{\bar{0}},(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}})_{\bar{0}})$-bimodule that extends to a complex $\tilde{X}$ of $(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}},b^{\eta,\varnothing,1}_{\mathbb{F}_{p}})_{C_{2}}$-modules.
+Let $H$ be the subgroup of $G$ generated by $\widetilde{S}^{\eta}_{U}$ and all $\widetilde{A}_{V_{i}}$’s.
+Now (by Propositions 5.2, 5.14 and Remark 2.3) $X^{\otimes d}$ induces a splendid Rickard equivalence between
+$\mathbb{F}_{p}(H\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d}\cong\mathbb{F}_{p}\widetilde{A}_{V_{1}}e_{\varnothing,1}\otimes_{\mathbb{F}_{p}}\cdots\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}\widetilde{A}_{V_{d}}e_{\varnothing,1}\cong(B^{\eta,\varnothing,1}_{\bar{0}})^{\otimes d}$
 and
-$\mathbb{F}_{p}(N_{G}(D)\cap H\cap\widetilde{S}_{V})e_{\varnothing,0}\cong\mathbb{F}_{p}N_{\widetilde{A}_{V_{1}}}(D_{1})e_{\varnothing,0}\otimes_{\mathbb{F}_{p}}\cdots\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}N_{\widetilde{A}_{V_{d}}}(D_{d})e_{\varnothing,0}\cong(b^{\varnothing,1}_{\bar{0}})^{\otimes d}.$
-We now adopt the identifications (6.3). By Proposition 2.12, $\tilde{X}^{\otimes d}$ extends to a complex $Y_{0}$ of modules for the group algebra of
-$\langle(H\cap\widetilde{S}_{V})\times(N_{G}(D)\cap H\cap\widetilde{S}_{V}),(u_{i},u_{i}),(T_{w},T_{w})\rangle\leq(N^{\rho,d}\cap\widetilde{S}_{V})\times(N_{G}(D)\cap\widetilde{S}_{V}),$
-where $i$ ranges over $\{1,\cdots,d\}$ with $u_{i}\in N_{\widetilde{S}_{V_{i}}}(D_{i})\backslash N_{\widetilde{A}_{V_{i}}}(D_{i})$, and $w$ ranges over $S_{d}$ with $T_{w}$ defined as in 5.12.
-By Proposition 4.7 (iii), $Y_{0}$ is a splendid complex.
+$\mathbb{F}_{p}(N_{G}(D)\cap H\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,0}\cong\mathbb{F}_{p}N_{\widetilde{A}_{V_{1}}}(D_{1})e_{\varnothing,0}\otimes_{\mathbb{F}_{p}}\cdots\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}N_{\widetilde{A}_{V_{d}}}(D_{d})e_{\varnothing,0}\cong(b^{\eta,\varnothing,1}_{\bar{0}})^{\otimes d}.$
+We now adopt the identifications (7.4). By Proposition 3.8, $X^{\otimes d}$ extends to a complex $Y_{0}$ of modules for the group algebra of
+$\langle(H\cap\widetilde{S}^{\eta}_{V})\times(N_{G}(D)\cap H\cap\widetilde{S}^{\eta}_{V}),(u_{i},u_{i}),(T_{w},T_{w})\rangle\leq(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})\times(N_{G}(D)\cap\widetilde{S}^{\eta}_{V}),$
+where $i$ ranges over $\{1,\cdots,d\}$ with $u_{i}\in N_{\widetilde{S}^{\eta}_{V_{i}}}(D_{i})\backslash N_{\widetilde{A}_{V_{i}}}(D_{i})$, and $w$ ranges over $S_{d}$ with $T_{w}$ chosen as in Lemma 6.12.
 We then induce $Y_{0}$ up to
-$(N^{\rho,d}\cap\widetilde{S}_{V})\times(N_{G}(D)\cap\widetilde{S}_{V})$
+$(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})\times(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})$
 to obtain a complex $Y$. Since $d<p$, we have $p\nmid|C_{2}\wr S_{d}|$. It follows from Theorem 5.4 that $Y$ induces a $(C_{2}\wr S_{d})$-graded Rickard equivalence between
-$\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}\cong B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$
+$\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d}\cong B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}$
 and
-$\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}\cong b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}.$
-By Proposition 4.7 (i), $Y$ is a splendid complex.
+$\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,0}^{\otimes d}\cong b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}.$
+By Proposition 5.6 (iii), $Y$ is a splendid complex.
 Next consider the group homomorphism
@@
 Let $\mathcal{K}$ be the kernel of this homomorphism.
-Since $B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$ is a $C_{2}\wr S_{d}$-graded algebra, we use the notation $(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})_{g}$ to denote the $g$-component of $B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$ for any $g\in C_{2}\wr S_{d}$.
+Since $B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}$ is a $C_{2}\wr S_{d}$-graded algebra, we use the notation $(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})_{g}$ to denote the $g$-component of $B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}$ for any $g\in C_{2}\wr S_{d}$.
 Then we have
-$(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})_{\bar{0}}=\bigoplus_{g\in\mathcal{K}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})_{g}.$
+$(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})_{\bar{0}}=\bigoplus_{g\in\mathcal{K}}(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})_{g}.$
 Since $Y$ is a $C_{2}\wr S_{d}$-graded bimodule, $Y$ can be regarded as a bisupermodule ($\mathbb{Z}/2\mathbb{Z}$-graded bimodule) via
@@
 By the definition of a $(C_{2}\wr S_{d})$-graded Rickard equivalence, we see that $Y$ (considered as a bisupermodule) induces a $\mathbb{Z}/2\mathbb{Z}$-graded Rickard equivalence between
-$\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}\cong B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$
+$\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d}\cong B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}$
 and
-$\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}\cong b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}.$
-It is straightforward to check that the $(B^{\rho,0}_{\mathbb{F}_{p}},B^{\rho,0}_{\mathbb{F}_{p}})$-bisupermodule $B^{\rho,0}_{\mathbb{F}_{p}}$ induces a Morita self-superequivalence of $B^{\rho,0}_{\mathbb{F}_{p}}$; this bimodule is splendid because it is isomorphic to a direct summand of $\mathbb{F}_{p}\widetilde{S}_{r}\cong{\rm Ind}_{\Delta\widetilde{S}_{r}}^{\widetilde{S}_{r}\times\widetilde{S}_{r}}(\mathbb{F}_{p})$.
-Hence by Proposition 4.2, the complex $B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y$ of $\mathbb{Z}/2\mathbb{Z}$-graded $(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\rho,1}^{\otimes d})$-$(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\rho,0}^{\otimes d})$-bimodule induces a $\mathbb{Z}/2\mathbb{Z}$-graded Rickard equivalence between
-$\mathbb{F}_{p}N^{\rho,d}f\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}$
+$\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,0}^{\otimes d}\cong b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}.$
+It is straightforward to check that the $(B^{\eta,\rho,0}_{\mathbb{F}_{p}},B^{\eta,\rho,0}_{\mathbb{F}_{p}})$-bisupermodule $B^{\eta,\rho,0}_{\mathbb{F}_{p}}$ induces a Morita self-superequivalence of $B^{\eta,\rho,0}_{\mathbb{F}_{p}}$; this bimodule is splendid because it is isomorphic to a direct summand of $\mathbb{F}_{p}\widetilde{S}^{\eta}_{r}\cong{\rm Ind}_{\Delta\widetilde{S}^{\eta}_{r}}^{\widetilde{S}^{\eta}_{r}\times\widetilde{S}^{\eta}_{r}}(\mathbb{F}_{p})$.
+Hence by Proposition 5.2, the complex $B^{\eta,\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y$ of $\mathbb{Z}/2\mathbb{Z}$-graded $(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d})$-$(B^{\eta,\varnothing,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})e_{\rho,0}^{\otimes d})$-bimodule induces a $\mathbb{Z}/2\mathbb{Z}$-graded Rickard equivalence between
+$\mathbb{F}_{p}N^{\rho,d}f\cong B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d}$
 and
-$\mathbb{F}_{p}N_{G}(D)e_{\rho,d}\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}.$
-By Proposition 4.12, $B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y$ is a splendid complex of $(\mathbb{F}_{p}N^{\rho,d}f,\mathbb{F}_{p}N_{G}(D)e_{\rho,d})$-bimodules.
-Now by Theorem 4.4, $(B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y)_{\bar{0}}$ induces a Rickard equivalence between $\mathbb{F}_{p}N^{\rho,d}_{\bar{0}}f$ and $\mathbb{F}_{p}N_{G}(D)_{\bar{0}}e_{\rho,d}$.
-Since $(B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y)_{\bar{0}}$ is isomorphic to a direct summand of the complex
-${\rm Res}_{N^{\rho,d}_{\bar{0}}\times N_{G}(D)_{\bar{0}}}^{N^{\rho,d}\times N_{G}(D)}(B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y),$
+$\mathbb{F}_{p}N_{G}(D)e_{\rho,d}\cong B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,0}^{\otimes d}.$
+By Proposition 5.14, $B^{\eta,\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y$ is a splendid complex of $(\mathbb{F}_{p}N^{\rho,d}f,\mathbb{F}_{p}N_{G}(D)e_{\rho,d})$-bimodules.
+Now by Theorem 5.3, $(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y)_{\bar{0}}$ induces a Rickard equivalence between $\mathbb{F}_{p}N^{\rho,d}_{\bar{0}}f$ and $\mathbb{F}_{p}N_{G_{\bar{0}}}(D)e_{\rho,d}$.
+Since $(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y)_{\bar{0}}$ is isomorphic to a direct summand of the complex
+${\rm Res}_{N^{\rho,d}_{\bar{0}}\times N_{G_{\bar{0}}}(D)}^{N^{\rho,d}\times N_{G}(D)}(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y),$
 […]
 8 The Morita superequivalence bimodule, and proofs of Theorems 1.8 and 1.9
-7.1.
+8.1.
 Throughout this section we set $d$ to be an integer with $1\leq d<p$ and $\rho$ a $d$-Rouquier $\bar{p}$-core.
@@
 We also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$.
-Let $R\in\{\mathcal{O},k\}$ and $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$.
-We will adopt the notation in 5.13, in particular, $r=|\rho|$, $n=r+dp$,
-$G=\widetilde{S}_{n},\quad L=\widetilde{S}_{U,V_{1},\dots,V_{d}},\quad N^{\rho,d}=N_{G}(\widetilde{S}_{V_{1},\cdots,V_{d}}),\quad$
+Let $(R,R_{p})\in\{(\mathcal{O},\mathbb{Z}_{p}),(k,\mathbb{F}_{p})\}$.
+We may assume that $R$ is finitely generated as $R_{p}$-module.
+We will adopt the notation in 6.13, in particular, $r=|\rho|$, $n=r+dp$,
+$G=\widetilde{S}^{\eta}_{n},\quad L=\widetilde{S}^{\eta}_{U,V_{1},\dots,V_{d}},\quad N^{\rho,d}=N_{G}(\widetilde{S}^{\eta}_{V_{1},\cdots,V_{d}}),\quad$
 $b=e_{\rho,d}\in RG,\quad f=e_{\rho,0}\otimes e_{\varnothing,1}^{(1)}\otimes\dots\otimes e_{\varnothing,1}^{(d)}\in RLe_{z},$
@@
 is chosen as in Lemma 6.15.
-We identify $RLf$ with $B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\otimes d}$ and $RN^{\rho,d}f$ with $B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})$ via the isomorphisms (5.8).
+We identify $RLf$ with $B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R})^{\otimes d}$ and $RN^{\rho,d}f$ with $B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})$ via the isomorphisms (6.8).
 Let $\ell:=(p-1)/2$.
-7.2.
-The bisupermodule which induces a Morita superequivalence. We set $M^{R}$ to be the $(B^{\varnothing,1}_{R},B^{\varnothing,1}_{R})$-bisupermodule
-$M^{R}:=\Omega_{B^{\varnothing,1}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\rm sop}}^{\ell}(B^{\varnothing,1}_{R}).$
-Then we have the $((B^{\varnothing,1}_{R})^{\otimes d},(B^{\varnothing,1}_{R})^{\otimes d})$-bisupermodule $(M^{R})^{\boxtimes d}$.
-For $1\leq i\leq d$, we identify $(B^{\varnothing,1}_{R})^{(i)}:=R\widetilde{S}_{V_{i}}e_{\varnothing,1}^{(i)}$ with $B^{\varnothing,1}_{R}=R\widetilde{S}_{p}e_{\varnothing,1}$ via (5.5), as in 5.12.
+8.2.
+The bisupermodule which induces a Morita superequivalence. We set $M^{R}$ to be the $(B^{\eta,\varnothing,1}_{R},B^{\eta,\varnothing,1}_{R})$-bisupermodule
+$M^{R}:=\Omega_{B^{\eta,\varnothing,1}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R})^{\rm sop}}^{\ell}(B^{\eta,\varnothing,1}_{R}).$
+Then we have the $((B^{\eta,\varnothing,1}_{R})^{\otimes d},(B^{\eta,\varnothing,1}_{R})^{\otimes d})$-bisupermodule $(M^{R})^{\boxtimes d}$.
+For $1\leq i\leq d$, we identify $(B^{\eta,\varnothing,1}_{R})^{(i)}:=R\widetilde{S}^{\eta}_{V_{i}}e_{\varnothing,1}^{(i)}$ with $B^{\eta,\varnothing,1}_{R}=R\widetilde{S}^{\eta}_{p}e_{\varnothing,1}$ via (6.5), as in 6.11.
 Denote by $(M^{R})^{(i)}$ the $i$-th factor in $(M^{R})^{\boxtimes d}$.
-That is, $(M^{R})^{(i)}$ is a $((B^{\varnothing,1}_{R})^{(i)},(B^{\varnothing,1}_{R})^{(i)})$-bisupermodule that we identify with $M^{R}$.
-Denote by $(M^{R})_{S_{d}}^{\boxtimes d}$ the $(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R},B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})_{S_{d}}$-supermodule defined as in (2.5).
-Recalling the identification $RLf\cong B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\otimes d}$, define the $(RLf,RLf)$-bisupermodule
-$M_{L}^{R}:=B^{\rho,0}_{R}\boxtimes_{R}(M^{R})^{\boxtimes d}.$
-We have also identified $RN^{\rho,d}f$ with $B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})$. Recalling (2.6), define the $(RN^{\rho,d}f,RN^{\rho,d}f)$-bisupermodule
-$M_{N^{\rho,d}}^{R}:=B^{\rho,0}_{R}\boxtimes_{R}(M^{R}\wr_{s}\mathcal{T}_{d}^{R})=B^{\rho,0}_{R}\boxtimes_{R}\Big({\rm Ind}_{(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R},B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})_{S_{d}}}^{(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})^{\rm sop}}(M^{R})^{\boxtimes d}_{S_{d}}\Big).$
-By [22, Lemma 7.2 (ii)], the $(RN^{\rho,d}f,RN^{\rho,d}f)$-bimodule $M_{N^{\rho,d}}^{R}$ is module-indecomposable with vertex $\Delta D:=\{(u,u)\mid u\in D\}$, and by [22, Lemma 4.8 (ii)], we have $N_{G\times N^{\rho,d}}(\Delta D)\leq N^{\rho,d}\times N^{\rho,d}$.
+That is, $(M^{R})^{(i)}$ is a $((B^{\eta,\varnothing,1}_{R})^{(i)},(B^{\eta,\varnothing,1}_{R})^{(i)})$-bisupermodule that we identify with $M^{R}$.
+Denote by $(M^{R})_{S_{d}}^{\boxtimes d}$ the $(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R},B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})_{S_{d}}$-supermodule defined as in (3.2).
+Recalling the identification $RLf\cong B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R})^{\otimes d}$, define the $(RLf,RLf)$-bisupermodule
+$M_{L}^{R}:=B^{\eta,\rho,0}_{R}\boxtimes_{R}(M^{R})^{\boxtimes d}.$
+We have also identified $RN^{\rho,d}f$ with $B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})$.
+Recalling (3.3), define the $(RN^{\rho,d}f,RN^{\rho,d}f)$-bisupermodule
+$M_{N^{\rho,d}}^{R}:=B^{\eta,\rho,0}_{R}\boxtimes_{R}(M^{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})=B^{\eta,\rho,0}_{R}\boxtimes_{R}\Big({\rm Ind}_{(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R},B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})_{S_{d}}}^{(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})\otimes_{R}(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})^{\rm sop}}(M^{R})^{\boxtimes d}_{S_{d}}\Big).$
+By [23, Lemma 7.2 (ii)] and Proposition 3.6 (i), the $(RN^{\rho,d}f,RN^{\rho,d}f)$-bimodule $M_{N^{\rho,d}}^{R}$ is module-indecomposable with vertex $\Delta D:=\{(u,u)\mid u\in D\}$, and by [23, Lemma 4.8 (ii)], we have $N_{G\times N^{\rho,d}}(\Delta D)\leq N^{\rho,d}\times N^{\rho,d}$.
 Hence we can now define $X^{R}$ to be the super Green correspondent of $M_{N^{\rho,d}}^{R}$ in $G\times N^{\rho,d}$; see Theorem 4.11 for the definition of the super Green correspondence.
-By [22, Lemma 7.5], $X^{R}$ is an $(RGb,RN^{\rho,d}f)$-bisupermodule.
-Theorem 7.3 ([22, Theorem 8.13]).
+By [23, Lemma 7.5], $X^{R}$ is an $(RGb,RN^{\rho,d}f)$-bisupermodule.
+Theorem 8.3 ([23, Theorem 8.13 and Lemma 9.2]).
 The bisupermodules $X^{R}$ and $(X^{R})^{*}$ induce a Morita superequivalence between $RGb$ and $RN^{\rho,d}f$.
-$X^{R}_{\bar{0}}$ and $(X^{R}_{\bar{0}})^{*}$ induce a Morita superequivalence between $RG_{\bar{0}}b$ and $RN_{\bar{0}}^{\rho,d}f$.
+The bimodules $X^{R}_{\bar{0}}$ and $(X^{R}_{\bar{0}})^{*}$ induce a Morita equivalence between $RG_{\bar{0}}b$ and $RN_{\bar{0}}^{\rho,d}f$.
 Both the bimodules $X^{R}$ and $X^{R}_{\bar{0}}$ have endopermutation sources.
 Theorem 8.4.
-Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$.
-Then $R_{p}Gf$ and $R_{p}Nf$ are Morita superequivalent.
+The blocks $R_{p}Gf$ and $R_{p}Nf$ are Morita superequivalent.
 More precisely, there is an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$, together with its dual $(X^{R_{p}})^{*}$, inducing a Morita superequivalence between $R_{p}Gb$ and $R_{p}N^{\rho,d}f$ and fulfilling $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$.
-The bimodule $X^{\mathbb{F}_{p}}$ has endopermutation source.
+The bimodule $X^{R_{p}}$ has the endopermutation $R_{p}\Delta D$-source $U^{R_{p}}:=\Omega_{R_{p}\Delta D_{1}}^{\ell}(R_{p})\otimes_{R_{p}}\cdots\otimes_{R_{p}}\Omega_{R_{p}\Delta D_{d}}^{\ell}(R_{p})$.
 Proof.
-By Propositions 3.5 and 3.4, it suffices to show that there exists an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$ such that $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$.
-By Remark (5.14) we have an isomorphism of $R_{p}$-superalgebras
-$R_{p}N^{\rho,d}f\cong B^{\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}}).$
-We let $M^{R_{p}}$ be the $(B^{\varnothing,1}_{R_{p}},B^{\varnothing,1}_{R_{p}})$-bimodule
-$M^{R_{p}}:=\Omega_{B^{\varnothing,1}_{R_{p}}\otimes_{R_{p}}(B^{\varnothing,1}_{R_{p}})^{\rm sop}}^{\ell}(B^{\varnothing,1}_{R_{p}}).$
+For the first statement, by Proposition 4.5, it suffices to show that there exists an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$ such that $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$.
+By Remark 6.14 we have an isomorphism of $R_{p}$-superalgebras
+$R_{p}N^{\rho,d}f\cong B^{\eta,\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}}).$
+We let $M^{R_{p}}$ be the $(B^{\eta,\varnothing,1}_{R_{p}},B^{\eta,\varnothing,1}_{R_{p}})$-bimodule
+$M^{R_{p}}:=\Omega_{B^{\eta,\varnothing,1}_{R_{p}}\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}})^{\rm sop}}^{\ell}(B^{\eta,\varnothing,1}_{R_{p}}).$
 Then by Proposition 4.8 we have $M^{R}\cong R\otimes_{R_{p}}M^{R_{p}}$. Define the $(R_{p}N^{\rho,d}f,R_{p}N^{\rho,d}f)$-bisupermodule
-$\begin{split}M_{N^{\rho,d}}^{R_{p}}&:=B^{\rho,0}_{R_{p}}\boxtimes_{R_{p}}(M^{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})\\ &=B^{\rho,0}_{R_{p}}\boxtimes_{R_{p}}\Big({\rm Ind}_{(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}},B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})_{S_{d}}}^{(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})\otimes_{R_{p}}(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})^{\rm sop}}(M^{R_{p}})^{\boxtimes d}_{S_{d}}\Big).\end{split}$
-Then we have $M_{N^{\rho,d}}^{R}\cong R\otimes_{R_{p}}M_{N^{\rho,d}}^{R_{p}}$ as $(RN^{\rho,d}f,RN^{\rho,d}f)$-bisupermodules.
+$\begin{split}M_{N^{\rho,d}}^{R_{p}}&:=B^{\eta,\rho,0}_{R_{p}}\boxtimes_{R_{p}}(M^{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})\\ &=B^{\eta,\rho,0}_{R_{p}}\boxtimes_{R_{p}}\Big({\rm Ind}_{(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}},B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})_{S_{d}}}^{(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})^{\rm sop}}(M^{R_{p}})^{\boxtimes d}_{S_{d}}\Big).\end{split}$
+Then by Proposition 3.6, we have $M_{N^{\rho,d}}^{R}\cong R\otimes_{R_{p}}M_{N^{\rho,d}}^{R_{p}}$ as $(RN^{\rho,d}f,RN^{\rho,d}f)$-bisupermodules.
 Now by Proposition 4.12, there is an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$ such that $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$, as desired.
+Denote by $U^{R_{p}}$ the $R_{p}\Delta D$-module $\Omega_{R_{p}\Delta D_{1}}^{\ell}(R_{p})\otimes_{R_{p}}\cdots\otimes_{R_{p}}\Omega_{R_{p}\Delta D_{d}}^{\ell}(R_{p})$. By Proposition 4.8, we have
+$R\otimes_{R_{p}}U^{R_{p}}\cong\Omega_{R\Delta D_{1}}^{\ell}(R)\otimes_{R}\cdots\otimes_{R}\Omega_{R\Delta D_{d}}^{\ell}(R)$
 […]
-By [22, Lemma 3.10], Theorem 7.4 has the following corollary:
 Corollary 8.5.
-With the notation of Theorem 7.4, the $(R_{p}G_{\bar{0}}b,R_{p}N_{\bar{0}}^{\rho,d}f)$-bimodule $X_{\bar{0}}^{R_{p}}$ and its dual induce a Morita equivalence between $R_{p}G_{\bar{0}}b$ and $R_{p}N_{\bar{0}}^{\rho,d}f$.
-Corollary 7.5 answers a question of Kleshchev and Livesey proposed in [22, Remark 8.15] in the highest generality of their question.
+With the notation of Theorem 8.4, the $(R_{p}G_{\bar{0}}b,R_{p}N_{\bar{0}}^{\rho,d}f)$-bimodule $X_{\bar{0}}^{R_{p}}$ induces a Morita equivalence between $R_{p}G_{\bar{0}}b$ and $R_{p}N_{\bar{0}}^{\rho,d}f$ and has an endopermutation source.
+Proof.
+By Theorem 8.4 and [23, Lemma 3.10], $X_{\bar{0}}^{R_{p}}$ induces a Morita equivalence between $R_{p}G_{\bar{0}}b$ and $R_{p}N_{\bar{0}}^{\rho,d}f$.
+Since ${\rm Res}_{G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}}^{G\times N^{\rho,d}}(X^{R_{p}})=X_{\bar{0}}^{R_{p}}\oplus X^{R_{p}}_{\bar{1}}$, $X^{R_{p}}_{\bar{0}}$ is isomorphic to a direct summand of
+${\rm Res}_{G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}}^{G\times N^{\rho,d}}{\rm Ind}_{\Delta D}^{G\times N^{\rho,d}}(U^{R_{p}}).$
+[…]
+Corollary 8.5 answers a question of Kleshchev and Livesey proposed in [23, Remark 8.15] in the highest generality of their question.
 Theorem 8.6.
@@
 Proof.
-Since there is a finite extension of $\mathbb{Q}_{p}$ containing a primitive $(2n!)$-th root of unity and containing $\sqrt{-1}$, we may assume that $k$ is finite.
-By [22, Lemma 9.1], the $(\mathcal{O}Gb,\mathcal{O}N^{\rho,d}f)$-bimodule $X^{\mathcal{O}}$ is isomorphic to a direct summand of
+Since there is a finite extension of $\mathbb{Q}_{p}$ containing a primitive $(2n!)$-th root of unity and containing $\sqrt{-1}$ and $\sqrt{2}$, we may assume that $k$ is finite.
+By [23, Lemma 9.2], the $(\mathcal{O}Gb,\mathcal{O}N^{\rho,d}f)$-bimodule $X^{\mathcal{O}}$ is isomorphic to a direct summand of
 $\mathcal{O}Gi\otimes_{\mathcal{O}D}{\rm Ind}_{\Delta D}^{D\times D}(U)\otimes j\mathcal{O}N^{\rho,d}$
-for some source idempotents $i\in(\mathcal{O}Gb)^{D}$ and $j\in(\mathcal{O}N^{\rho,d}f)^{D}$, where $U$ is an endopermutation $\mathcal{O}\Delta D$-module.
-We can regard $U$ as an endopermutation $\mathcal{O}D$-module via the isomorphism $D\cong\Delta D$.
-By [22, Lemma 9.1], $U$ has an $N_{G}(D)$-stable endosplit $p$-permutation resolution $\mathcal{Y}$; see [30, Section 7] or [24, Definition 7.11.1] for this notion.
-It follows that $\mathcal{Y}$ is $\mathcal{F}$-stable, where $\mathcal{F}$ is the fusion system on $D$ determined by the source idempotent $i$ of $\mathcal{O}Gb$.
-By [24, Theorem 9.11.5 (i)], $X^{\mathcal{O}}$ has an endosplit $p$-permutation resolution $Y^{\mathcal{O}}$, and $Y^{\mathcal{O}}$ induces a splendid Rickard equivalence between $\mathcal{O}Gb$ and $\mathcal{O}N^{\rho,d}f$.
-Let $X^{k}=k\otimes_{\mathcal{O}}X^{\mathcal{O}}$, $Y^{k}=k\otimes_{\mathcal{O}}Y^{\mathcal{O}}$ and $X^{\mathbb{F}_{p}}=\mathbb{F}_{p}\otimes_{\mathbb{Z}_{p}}X^{\mathbb{Z}_{p}}$.
-By definition, one checks that $Y^{k}$ is an endosplit $p$-permutation resolution of $X^{k}$.
-By Theorem 7.4, we have $X^{k}\cong k\otimes_{\mathbb{F}_{p}}X^{\mathbb{F}_{p}}$.
-Then by [28, Theorem 1.3 (b)], there is a splendid complex of $(\mathbb{F}_{p}G\bar{b},\mathbb{F}_{p}N^{\rho,d}\bar{f})$-bimodules $Y^{\mathbb{F}_{p}}$ such that $Y^{k}\cong k\otimes_{\mathbb{F}_{p}}Y^{\mathbb{F}_{p}}$.
-By [19, Proposition 4.5 (a)], $Y^{\mathbb{F}_{p}}$ induce a splendid Rickard equivalence between $\mathbb{F}_{p}G\bar{b}$ and $\mathbb{F}_{p}N^{\rho,d}\bar{f}$.
-By lifting property of splendid Rickard equivalences ([30, Theorem 5.2]), $Y^{\mathbb{F}_{p}}$ lifts to a splendid Rickard complex $Y^{\mathbb{Z}_{p}}$ of $(\mathbb{Z}_{p}Gb,\mathbb{Z}_{p}N^{\rho,d}f)$-bimodules.
 […]
 Theorem 8.7.
-Let $G=\widetilde{S}_{n}^{+}$ (and hence $G_{\bar{0}}:=\widetilde{A}_{n}$). Then $\mathbb{Z}_{p}G_{\bar{0}}b$ is splendidly Rickard equivalent to $\mathbb{Z}_{p}N^{\rho,d}_{\bar{0}}f$.
+The block $\mathbb{Z}_{p}G_{\bar{0}}b$ is splendidly Rickard equivalent to $\mathbb{Z}_{p}N^{\rho,d}_{\bar{0}}f$.
 Proof.
@@
 where the isomorphism is by using the Mackey formula.
-Again since $p\nmid|G\times N^{\rho,d}:G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}|$, any vertex of the bimodule $X_{\bar{0}}^{\mathcal{O}}$ has the same order as a vertex of the bimodule $X^{\mathcal{O}}$.
+Again since $p\nmid|G\times N^{\rho,d}:G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}|$, any vertex of the bimodule $X_{\bar{0}}^{\mathcal{O}}$ has the same order with any vertex of the bimodule $X^{\mathcal{O}}$.
 It follows that $X_{\bar{0}}^{\mathcal{O}}$ has an endopermutation $)\mathcal{O}({}^{t}(\Delta D$))-source $^{t}U$ for some $t\in G\times N^{\rho,d}/G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}$.
@@
 Write $P:={}^{t_{1}}D$, $Q:={}^{t_{2}}D$ and $\varphi:P\cong Q$ the group isomorphism induced by conjugation action of $t_{2}t_{1}^{-1}$.
-By [24, Proposition 9.7.1], $P$ is a defect group of $\mathcal{O}G_{\bar{0}}b$ and $Q$ is a defect group of $\mathcal{O}N^{\rho,d}_{\bar{0}}f$.
+By [25, Proposition 9.7.1], $P$ is a defect group of $\mathcal{O}G_{\bar{0}}b$ and $Q$ is a defect group of $\mathcal{O}N^{\rho,d}_{\bar{0}}f$.
 Note that $X_{\bar{0}}^{\mathcal{O}}$ is isomorphic to a direct summand of
 $\mathcal{O}G_{\bar{0}}b\otimes_{\mathcal{O}P}{\rm Ind}_{\Delta\varphi}^{P\times Q}({}^{t_{1}}U)\otimes_{\mathcal{O}Q}\mathcal{O}N^{\rho,d}_{\bar{0}}f.$
-Then using the same argument as in [24, page 320], $X_{\bar{0}}^{\mathcal{O}}$ is isomorphic to a direct summand of
+Then using the same argument as in [25, page 320], $X_{\bar{0}}^{\mathcal{O}}$ is isomorphic to a direct summand of
 $\mathcal{O}G_{\bar{0}}i_{0}\otimes_{\mathcal{O}P}{\rm Ind}_{\Delta\varphi}^{P\times Q}({}^{t_{1}}U)\otimes_{\mathcal{O}Q}j_{0}\mathcal{O}N^{\rho,d}_{\bar{0}}$
@@
 Remark 8.8.
-By [22, Theorem 9.7], when $p\equiv 1~{\rm mod}~4$, the $\mathcal{O}D$-source $U$ of $X^{\mathcal{O}}$ is the trivial module.
+By [23, Theorem 9.7], when $p\equiv 1~{\rm mod}~4$, the $\mathcal{O}D$-source $U$ of $X^{\mathcal{O}}$ is the trivial module.
 Hence $X^{\mathcal{O}}$ induces a splendid Morita equivalence between $\mathcal{O}Gb$ and $\mathcal{O}N^{\rho,d}d$, and $X^{\mathcal{O}}_{\bar{0}}$ induces a splendid Morita equivalence between $\mathcal{O}G_{\bar{0}}b$ and $\mathcal{O}N^{\rho,d}_{\bar{0}}f$.
@@
 If $p\equiv 1~{\rm mod}~4$, then $X^{\mathbb{Z}_{p}}$ induces a splendid Morita equivalence between $\mathbb{Z}_{p}Gb$ and $\mathbb{Z}_{p}N^{\rho,d}d$, and $X^{\mathbb{Z}_{p}}_{\bar{0}}$ induces a splendid Morita equivalence between $\mathbb{Z}_{p}G_{\bar{0}}b$ and $\mathbb{Z}_{p}N^{\rho,d}_{\bar{0}}f$.
-We can now construct splendid Rickard equivalences between $\mathbb{Z}_{p}Gb$ (resp. $\mathbb{Z}_{p}G_{\bar{0}}b$) and its Brauer correspondent in the normaliser of a defect group.
+We can now construct splendid Rickard equivalences between $\mathbb{Z}_{p}Gb$ (resp.
+$\mathbb{Z}_{p}G_{\bar{0}}b$) and its Brauer correspondent in $\mathbb{Z}_{p}N_{G}(D)$ (resp.
+$\mathbb{Z}_{p}N_{G_{\bar{0}}}(D)$).
 Theorem 8.10.
-Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$.
-Let $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$.
-Then $R_{p}Gb$ and $R_{p}G_{\bar{0}}b$ are splendidly Rickard equivalent to their Brauer correspondent in $R_{p}N_{G}(D)$ and $R_{p}N_{G_{0}}(D)$, respectively.
+The block $R_{p}Gb$ (resp. $R_{p}G_{\bar{0}}b$) is splendidly Rickard equivalent to its Brauer correspondent in $R_{p}N_{G}(D)$ (resp. $R_{p}N_{G_{0}}(D)$).
 Proof.
-By the lifting theorem of splendid Rickard equivalences [30, Theorem 5.2], it suffices to consider $R_{p}=\mathbb{F}_{p}$.
+By the lifting theorem of splendid Rickard equivalences (Theorem 5.9), it suffices to consider $R_{p}=\mathbb{F}_{p}$.
 By Theorem 7.6, there is a splendid Rickard equivalence between the Brauer correspondent of $\mathbb{F}_{p}Gb$ (resp.
@@
 $\mathbb{F}_{p}N^{\rho,d}_{\bar{0}}f$).
-Now the statements follows from Theorems 7.6, 7.7 and Proposition 4.8.
+Now the statements follows from Theorems 8.6, 8.7 and Proposition 5.7.
 […]
@@
 Alexander Kleshchev for helpful explanation.
-The first author is supported by National Key R & D Profram of China (No.
+The first author is supported by National Key R & D Program of China (No.
 2020YFE020 4200) and NSFC (Nos.
 12431001, 12350710787).
-The second author is supported by Fundamental Research Funds for the Central Universities (CCNU24XJ028/25JC025/25JCPT031), and NSFC (Nos.
+The second author is supported by Fundamental Research Funds for the Central Universities (CCNU24XJ028/25JC025/25JCPT031), China Postdoctoral Science Foundation (GZC20252006, 2025T001HB) and NSFC (Nos.
 12501024, 12471016).